### v1

# Estimating the linear relation between variables that are never jointly observed: an application in in vivo experiments

Abstract
This work is motivated by in vivo experiments in which measurement are destructive so that the variables of interest can never be observed simultaneously when the aim is to estimate the regression coefficients of a linear regression. Assuming that the global experiment can be decomposed into sub experiments (corresponding for example to different doses) with distinct first moments, we propose different estimators of the linear regression which take account of that additional information. We consider estimators based on moments as well as estimators based optimal transport theory. These estimators are proved to be consistent as well as asymptotically Gaussian under weak hypotheses. The asymptotic variance has no explicit expression, except in some particular cases, and specific bootstrap approaches are developed to build confidence intervals for the estimated parameter. A Monte Carlo study is conducted to assess and compare the finite sample performances of the different approaches.
1 Motivation and context
In order to study the effects of a certain treatment on a living organism, in vivo experiments are often conducted. In the context of a complex organism response, scientists may be interested in studying multiple variables describing the effect from different perspectives. In particular, such variables of interest often include a macroscopic biomarker only available through in vivo data on the one hand, and a microscopic biomarker that can also be observed on a cellular level. The interest in this case lies in predicting the former with the latter. For instance, in the context of studying the adverse effects induced by radiotherapy on healthy tissues, the potential outcomes of interest may manifest in the form of lesions, that are quantified in a form a certain macroscopic biomarker, and the levels of a some predictor such as gene expression. These measures often require sacrificing the animal. As a result, the quantities of interest cannot be observed on the same animals. Since the goal is to establish relationships between these variables, a problem of statistical data fusion arises.
In this work we propose an approach to estimate the relationship between the variables that are not simultaneously observed under the experimental setting described above, based on a conditional model assuming linear relationship within every component (experimental condition). An estimator derived with the method of moments as well as optimal transport solution using Wasserstein distance are considered for the real data problem in question.
1.1 Example of an in vivo experiment
Figure 1: Schematic representation of the design of an in vivo experiment studying the effect of irradiated volume.
An illustrative example of an in vivo experiment where the variables of interest are never observed simultaneously is presented in Figure 1. In this experiment, mice are irradiated on the lungs with different volume, with a goal of studying the role of irradiated volume in the appearing of radio-induced adverse effects. The latter are assessed by measuring septal thickening, a macroscopic biomarker of radio-induced adverse effects in the lungs. The other variable that is measured with the purpose of predicting the adverse effect related variable is the expression of multiple pro-inflammatory genes. As shown in Figure 1, there are two independent cohorts in the study, one is used to measure gene expression, whereas the other for measuring septal thickness. This is a results of both measures being of destructive nature, which does not allow them to be taken on the same animals.
Comparing distributions of the measurements arising from the two cohorts, such as those presented in Figure 2, one may suspect a correlation or even a linear relationship between the variables. In order to assert whether such a relationship exists, one has to connect two variables that are never observed simultaneously, which translates into solving a data fusion problem. In this example, there are four groups indicating time points (1, 3, 6 and 12 months after irradiation) when the corresponding animals are sacrificed and the measurements are taken. Thus, the categorical variable indication the time point, which is known for every observation, can be used as an additional variable in order to link the predictor and the predicted variables between each other.
Figure 2: Distribution of the data, collected from the irradiated patch under SBRT with 3 mm beam size: the expression of the gene IL6 on the left, and septal thickness on the right. The measurements were made 1, 3, 6 and 12 months after irradiation.
1.2 Existing research
The task of linking variables that are never jointly observed cannot be approached as a typical missing values problem, since most methods for inference on incomplete data require a sufficient overlap, which is completely absent in the case treated here. As a result, any approaches applying related frameworks such as multiple imputation in the context of data fusion are unsuitable for our application. For example, Carrig et al., (2015) use multiple imputation to integrate disparate datasets allowing for the absence of the overlap, but requiring a calibration dataset, where all the variables of interest must be jointly observed.
Other approaches to data fusion available in literature include factor analysis (Cudeck,, 2000), statistical matching (Mitsuhiro and Hoshino,, 2020), Bayesian network inference (Triantafillou et al.,, 2010; Tsamardinos et al.,, 2012) and Gaussian Markov combinations (Massa and Riccomagno,, 2017). These methods are designed for linking variables that are not observed simultaneously through covariates, present for both variables of interest. This corresponds to the properties of the in vivo data described in Section 1.1. However, the covariates in these approaches are random variables, typically continuous, and often assumed to be Gaussian, which is the case in Cudeck, (2000) and Massa and Riccomagno, (2017). The groups variable available through in vivo experiments may present in a continuous form, but the presence of such categories as control and sham makes it impossible to assume continuity and normality. The Bayesian network approaches introduced by Triantafillou et al., (2010) and Tsamardinos et al., (2012), aimed at inferring binary causal relationships between variables, are more suitable for large datasets with a high number of covariates. Finally, currently available research in statistical matching addresses such aspects as not-at-random missingness (Mitsuhiro and Hoshino,, 2021) and high dimensionality (Mitsuhiro and Hoshino,, 2020). This approach is based on the idea of comparing distances between the covariates from the datasets of interest, which cannot be done by taking the group variable as the covariate. It can be noted that the goal of the aforementioned examples in statistical matching is to group individuals before imputation, which is not necessary in our case since the groups are already known.
2 Estimating the linear relation between variables that are never jointly observed
2.1 Problem and presentation of the different identification approaches
We consider a real random variable $Y$ and a vector of $d$ real valued random regressors $\mathbf{X}=(X_{1},\ldots,X_{d})$ and suppose that the following linear regression hold
$Y$ $=\beta_{0}+\sum_{j=1}^{d}\beta_{j}X_{j}+\epsilon.$ (1)
[…]
In the absence of additional information and without any strong additional hypothesis, the parameters $(\beta_{0},\beta_{1},\ldots,\beta_{d})$ and the variance of the noise $\sigma^{2}_{\epsilon}$ cannot be identified. Indeed, if for example $X_{1}$ is centered with symmetric distribution, the coefficient $\beta_{1}$ can only be determined up to sign change since $\beta_{1}X_{1}$ and $\beta_{1}(-X_{1})$ have the same distribution.
To deal with this identification issue, we consider that we can perform different experiments in which the mean value of $X$ is allowed to vary. For that, we suppose that there are $K$ groups (corresponding to $K$ different experiments), defined by a discrete variable $G$ taking values in $\{1,\cdots,K\}$ observed simultaneously with $Y$ and with $X$. This means that we have now access to $(\mathbf{X},G)$ and $(Y,G)$ but not to $(\mathbf{X},Y,G)$. We also suppose that $\epsilon$ is independent of $G$.
Given $G=k$, for $k=1,\ldots,K$, we denote by $\mu_{Y}^{k}={\mathbb{E}}(Y|G=k)$ and $\mu_{X_{j}}^{k}={\mathbb{E}}(X_{j}|G=k)$, $j=1,\ldots,d$, the expected values within each group.
We now present two different approaches developed to identify the vector $\boldsymbol{\beta}=(\beta_{0},\ldots,\beta_{d})$ of unknown regression coefficients and the noise variance $\sigma_{\epsilon}^{2}$, taking account of the additional information given by the discrete variable $G$.
2.1.1 Moment approach
A first and simple approach is based on first moments identification. Taking the conditional expectation, given $G=k$, in (1), we have for $k=1,\ldots,K$,
$\mu_{Y}^{k}$ $=\beta_{0}+\sum_{j=1}^{d}\beta_{j}\mu_{X_{j}}^{k},$ (2)
since the residual term $\epsilon$ is supposed to satisfy $\mathbb{E}(\epsilon|G=k)=0$ for $k=1,\ldots,K$.
We denote by $\boldsymbol{\mu}_{1,X}$ the $K\times(d+1)$ design matrix, whose $k$th row is equal to $(1,\boldsymbol{\mu}_{X}^{k\top})$ with $\boldsymbol{\mu}_{X}^{k}=(\mu_{X_{1}}^{k},\cdots,\mu_{X_{d}}^{k})^{\top}$, by $\boldsymbol{\mu}_{Y}$ the $K$ dimensional vector with elements $(\mu_{Y}^{1},\ldots,\mu_{Y}^{K})$. The $K$ linear equations in (2) can we written in a matrix way, $\boldsymbol{\mu}_{Y}=\boldsymbol{\mu}_{1,X}\boldsymbol{\beta}$.
We introduce the following assumption, guaranteeing the identifiability of the model parameters:
$\mathbf{H}_{1}$ $\quad\mbox{rank}(\boldsymbol{\mu}_{1,X})=d+1,$
meaning that there are at least $K\geq d+1$ groups and that the $d+1$ column vectors of $\boldsymbol{\mu}_{1,X}$ span a vector space of dimension $d+1$ in $\mathbb{R}^{K}$.
Lemma 2.1.
If model (1) holds and assumption $\mathbf{H}_{1}$ is fulfilled, $\boldsymbol{\beta}$ is uniquely identified in terms of the conditional first order moments of $\mathbf{X}$ and $Y$ given $G$,
$\boldsymbol{\beta}$ $=\left(\boldsymbol{\mu}_{1,X}^{\top}\boldsymbol{\mu}_{1,X}\right)^{-1}\boldsymbol{\mu}_{1,X}^{\top}\boldsymbol{\mu}_{Y}.$
Additionally, the noise variance $\sigma_{\epsilon}^{2}$ satisfies
$\sigma_{\epsilon}^{2}$ $=\sigma_{Y}^{2}-\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}\boldsymbol{\beta}_{-0},$
where $\sigma_{Y}^{2}$ is the variance of $Y$, $\boldsymbol{\Gamma}_{X}$ is the covariance matrix of $\mathbf{X}$, and $\boldsymbol{\beta}_{-0}=(\beta_{1},\ldots,\beta_{d})$.
The proof of Lemma 2.1 is direct and thus omitted.
2.1.2 Optimal transport and minimum Wasserstein distance approach
The second approach is based on optimal transport, in particular on the idea of estimating the linear transformation of the distribution of $\mathbf{X}$ that is the closest to that of $Y$ with respect to the Wasserstein distance (see Panaretos and Zemel, (2019) for a general introduction for statisticians).
The optimal transport map T between Gaussian measures on $\mathbb{R}^{d}$ is linear, and the Wasserstein distance of order 2 between two Gaussian distributions $D_{1}$ and $D_{2}$, with $D_{1}\sim\mathcal{N}(\boldsymbol{\mu}_{1},\boldsymbol{\Gamma}_{1})$ and $D_{2}\sim\mathcal{N}(\boldsymbol{\mu}_{2},\boldsymbol{\Gamma}_{2})$, is equal to
$W_{2}^{2}(D_{1},D_{2})$ $=\|\boldsymbol{\mu}_{1}-\boldsymbol{\mu}_{2}\|^{2}+\mbox{tr}\left(\boldsymbol{\Gamma}_{1}+\boldsymbol{\Gamma}_{2}-2\left(\boldsymbol{\Gamma}_{2}^{1/2}\boldsymbol{\Gamma}_{1}\boldsymbol{\Gamma}_{2}^{1/2}\right)^{1/2}\right),$
[…]
If we assume that, given $G=k$, $\mathbf{X}$ is a Gaussian random vector and $\epsilon$ is Gaussian, we have, if the linear model (1) holds, that $Y$ is also Gaussian given $G=k$, with expectation $\mu_{Y}^{k}=\beta_{0}+\sum_{j=1}^{d}\beta_{j}\mu_{X_{j}}^{k}$ and variance $\sigma_{\epsilon}^{2}+\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}$, where $\boldsymbol{\Gamma}_{X}^{k}$ is the variance matrix of $\mathbf{X}$ given $G=k$. Thus, the Wasserstein distance between $D_{\gamma}$, the distribution of $\gamma_{0}+\boldsymbol{\gamma}_{-0}^{\top}\mathbf{X}+\epsilon$, and $D_{Y}$, the distribution of $Y$, is equal to
$W_{2}^{2}(D_{\gamma},D_{Y})$ $=\mathbb{E}\left[W_{2}^{2}(D_{\gamma},D_{Y})|G\right]$
$=\sum_{k=1}^{K}\pi_{k}\left[(\mu_{Y}^{k}-\alpha_{0}-\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\mu}_{X}^{k})^{2}+\left(\sigma_{Y,k}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma_{\epsilon}^{2}}\right)^{2}\right],$
where $\pi_{k}=\mathbb{P}[G=k]$ and $\sigma_{Y,k}^{2}=\mbox{Var}(Y|G=k)$.
We introduce the following loss criterion
$\varphi(\boldsymbol{\gamma},\sigma^{2})$ $=\sum_{k=1}^{K}\pi_{k}\left[(\mu_{Y}^{k}-\gamma_{0}-\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\mu}_{X}^{k})^{2}+\left(\sigma_{Y,k}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma^{2}}\right)^{2}\right].$ (3)
which evaluates the weighted Wasserstein distance between $Y$ and a linear combination of the $\mathbf{X}$ variables, contaminated by a Gaussian noise, with variance $\sigma^{2}$. We state, without proof, the following Lemma ensuring that the parameters $\boldsymbol{\beta}$ and $\sigma_{\epsilon}^{2}$ can be identified under general conditions.
Lemma 2.2.
If model (1) holds and if assumption $\mathbf{H}_{1}$ is fulfilled, $\varphi(\boldsymbol{\gamma},\sigma^{2})$ has its unique minimum at $\boldsymbol{\gamma}=\boldsymbol{\beta}$ and $\sigma^{2}=\sigma_{\epsilon}^{2}$.
2.2 Sampled data and estimators
We suppose that experiments are made for $K\geq 2$ different groups and that for each group $k$, for $k=1,\ldots,K$, we have two independent samples $(Y^{k}_{1},\ldots,Y^{k}_{n_{y}^{k}})$ and $(X^{k}_{j,1},\ldots,X^{k}_{j,n_{x}^{k}})_{j=1,\ldots,d}$, with sizes $n_{y}^{k}$ and $n_{x}^{k}$. For each unit $i=1,\ldots,n_{x}^{k}$ from group $k$, the vector of covariates is denoted by $\mathbf{X}_{i}^{k}=(X_{1,i},\ldots,X_{d,i})$. We also define $N_{x}=\sum_{k=1}^{K}n_{x}^{k}$ and $N_{y}=\sum_{k=1}^{K}n_{y}^{k}$, the total number of observations of the response $Y$ and the covariates $X_{1},\ldots,X_{d}$.
As shown in Lemma 2.1 and Lemma 2.2, identification of parameter $\boldsymbol{\beta}$ relies on the knowledge of the first two conditional moments, given $G$, of $Y$ and $\mathbf{X}$. For $k=1,\ldots,K$, we denote by $\widehat{\mu}_{Y}^{k}=\frac{1}{n_{y}^{k}}\sum_{i=1}^{n_{y}^{k}}Y^{k}_{i}$ and $\widehat{\mu}_{X_{j}}^{k}=\frac{1}{n_{x}^{k}}\sum_{i=1}^{n_{x}^{k}}X^{k}_{j,i}$ the empirical means within each group $k$ and by $\widehat{\sigma}^{2}_{Y,k}=\frac{1}{n_{y}^{k}}\sum_{i=1}^{n_{y}^{k}}\left(Y_{i}^{k}\right)^{2}-\left(\widehat{\mu}_{Y}^{k}\right)^{2}$, and $\widehat{\boldsymbol{\Gamma}}^{k}_{X}=\frac{1}{n_{x}^{k}}\sum_{i=1}^{n_{x}^{k}}\mathbf{X}_{i}^{k}(\mathbf{X}_{i}^{k})^{\top}-\widehat{\boldsymbol{\mu}}_{X}^{k}(\widehat{\boldsymbol{\mu}}_{X}^{k})^{\top}$, the empirical variances, with $\widehat{\boldsymbol{\mu}}_{X}^{k}=(\widehat{\mu}_{X_{1}}^{k},\ldots,\widehat{\mu}_{X_{d}}^{k})$. We also define the overall empirical mean and variance $\widehat{\mu}_{Y}=\frac{1}{N_{y}}\sum_{k=1}^{K}{n_{y}^{k}}\widehat{\mu}_{Y}^{k}$, $\widehat{\boldsymbol{\mu}}_{X}=\frac{1}{N_{x}}\sum_{k=1}^{K}{n_{x}^{k}}\widehat{\boldsymbol{\mu}}_{X}^{k}$, $\widehat{\sigma}^{2}_{Y}=\frac{1}{N_{y}}\sum_{k=1}^{K}\sum_{i=1}^{n_{y}^{k}}\left(Y_{i}^{k}\right)^{2}-\left(\widehat{\mu}_{Y}\right)^{2}$, $\widehat{\boldsymbol{\Gamma}}_{X}=\frac{1}{N_{x}}\sum_{k=1}^{K}\sum_{i=1}^{n_{x}^{k}}\mathbf{X}_{i}^{k}(\mathbf{X}_{i}^{k})^{\top}-\widehat{\boldsymbol{\mu}}_{X}\widehat{\boldsymbol{\mu}}_{X}^{\top}$.
We denote by $\widehat{\boldsymbol{\mu}}_{1,X}$ the $K\times(d+1)$ matrix, with the first column consisting of ones, and the rest equal to $\widehat{\boldsymbol{\mu}}_{X}$.
2.2.1 Moment estimators
If $\widehat{\boldsymbol{\mu}}_{1,X}$ is full rank, moment estimators of $\boldsymbol{\beta}=(\beta_{0},\beta_{1},\ldots,\beta_{d})$ can be built by considering the empirical counterpart of identification equations given in Lemma 2.1,
$\widehat{\boldsymbol{\beta}}^{M}$ $=\left(\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\widehat{\boldsymbol{\mu}}_{1,X}\right)^{-1}\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\widehat{\boldsymbol{\mu}}_{Y}$
where $\widehat{\boldsymbol{\mu}}_{Y}=(\widehat{\mu}_{Y}^{1},\ldots,\widehat{\mu}_{Y}^{K})$. We consider in the following a slightly more general moment estimator of $\boldsymbol{\beta}$, by introducing a weight $w_{k}$ given to each group $k$ of observations, with $w_{k}>0$ and $\sum_{k=1}^{K}w_{k}=1$. The weighted moment estimator $\boldsymbol{\beta}$ is defined as the minimizer of
$\psi(\boldsymbol{\gamma})$ $=\sum_{k=1}^{K}w_{k}\left[\widehat{\mu}_{Y}^{k}-\left(\gamma_{0}+\sum_{j=1}^{d}\gamma_{j}\widehat{\mu}_{X_{j}}^{k}\right)\right]^{2},$
which is unique if $\widehat{\boldsymbol{\mu}}_{1,X}$ is full rank and defined by
$\widehat{\boldsymbol{\beta}}^{M}$ $=\left(\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\mathbf{w}\widehat{\boldsymbol{\mu}}_{1,X}\right)^{-1}\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\mathbf{w}\widehat{\boldsymbol{\mu}}_{Y}$ (4)
with $\mathbf{w}=(w_{1},\ldots,w_{K})$. The first moment estimator considered just before corresponds to equal weights $w_{k}=K^{-1}$.
We can then define the following estimator of the noise variance,
$\widehat{\sigma}_{\epsilon}^{2,M}$ $=\widehat{\sigma}_{Y}^{2}-(\widehat{\boldsymbol{\beta}}_{-0}^{M})^{\top}\widehat{\boldsymbol{\Gamma}}_{X}\widehat{\boldsymbol{\beta}}_{-0}^{M}.$ (5)
2.2.2 Minimum Wasserstein distance estimators
Estimators of $\boldsymbol{\beta}$ and $\sigma^{2}_{\epsilon}$ based on an optimal transport criterion are derived by minimizing the empirical version $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2})$ of functional $\varphi(\boldsymbol{\gamma},\sigma^{2})$ defined by
$\varphi_{n}(\boldsymbol{\gamma},\sigma^{2})$ $=\sum_{k=1}^{K}\pi_{k}\left[(\widehat{\mu}_{Y}^{k}-\gamma_{0}-\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k})^{2}+\left(\widehat{\sigma}_{Y,k}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma^{2}}\right)^{2}\right].$ (6)
Note that in absence of an a priori information on the probability $\pi_{k}$ of observing group $k$, we can set $\pi_{k}=K^{-1}$. We denote by $(\widehat{\boldsymbol{\beta}}^{W},\widehat{\sigma}^{2,W})$ minimizers of $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2})$ which are obtained with iterative optimization algorithms based on gradient descent. The algorithm can be initialized with $(\widehat{\boldsymbol{\beta}}^{M},\widehat{\sigma}^{2,M})$
2.3 Consistency and asymptotic distribution
To study the asymptotic behavior of the estimators of $\boldsymbol{\beta}$ defined in previous Section, we suppose that the number $K$ of groups is kept fixed and that for all groups and all variables $\mathbf{X}$ and $Y$, the number of observations tends to infinity. This means that $n_{\min}=\min(n_{y}^{1},\cdots,n_{y}^{K},n_{x}^{1},\cdots,n_{x}^{K})$ the smallest sample size among all experiments is also supposed to tend to infinity.
Under general moment conditions, without any particular assumption on the probability law of $Y$ and $\mathbf{X}$ we get in the following two Lemmas that the moment estimators and the Wasserstein distance based estimators are consistent. The proofs are postponed to the Appendix.
Lemma 2.3.
If ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{2})<+\infty$, and assumption $\mathbf{H}_{1}$ is fulfilled, the sequence of estimators $(\widehat{\boldsymbol{\beta}}^{M},\widehat{\sigma}_{\epsilon}^{2,M})$ defined by (4) and (5) converges in probability to $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$ when $n_{\min}$ tends to infinity.
For the Wasserstein minimum distance estimator, since there is no explicit expression for the estimators, a compactness assumption is also made to obtain the consistency.
Lemma 2.4.
If ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{2})<+\infty$, $(\boldsymbol{\beta},\sigma^{2}_{\epsilon})\in\Theta$ and $\Theta$ is a compact set that does not contain 0, suppose that model (1) holds and hypothesis $\mathbf{H}_{1}$ is fulfilled, then the sequence of estimators $(\widehat{\boldsymbol{\beta}}^{W},\widehat{\sigma}^{2,W}_{\epsilon})$ that minimize (6) converges in probability to $(\beta,\sigma^{2}_{\epsilon})$ when $n_{\min}$ tends to infinity.
As far as the asymptotic distribution of the estimators is concerned, and for sake of simplicity and lighter notations, we suppose from now on that the number of experiments is the same for all groups and all variables, that is to say $n=n_{y}^{1}=\cdots n_{y}^{K}=n_{x}^{1}\cdots=n_{x}^{K}$.
Proposition 2.1.
Assume that the assumptions of Lemma 2.3 are fulfilled. Then as $n$ tends to infinity,
$\sqrt{n}\left(\widehat{\boldsymbol{\beta}}^{M}-\boldsymbol{\beta}\right)\rightsquigarrow\mathcal{N}\left(0,\boldsymbol{\Gamma}_{\beta_{M}}\right)$
where the expression of the asymptotic covariance matrix $\boldsymbol{\Gamma}_{\beta_{M}}$ is given in the proof.
This result is based on the CLT for empirical means and the application of the delta method, which involves computing the Jacobian of the inverse of matrices, making obtaining the explicit expression for $\boldsymbol{\Gamma}_{\beta_{M}}$ a difficult task when $d>1$.
Remark 2.1.
The weak convergence toward a Gaussian distribution presented in Proposition 2.1 would remain true, at the expense of heavier notations and a different asymptotic covariance matrix $\boldsymbol{\Gamma}_{\beta_{M}}$, provided that there exist two constants, $0<c\leq C$ such that
$0<c\leq\frac{\max(n_{y}^{1},\cdots,n_{y}^{K},n_{x}^{1},\cdots,n_{x}^{K})}{\min(n_{y}^{1},\cdots,n_{y}^{K},n_{x}^{1},\cdots,n_{x}^{K})}\leq C<+\infty,$ (7)
and $\min(n_{y}^{1},\cdots,n_{y}^{K},n_{x}^{1},\cdots,n_{x}^{K})\to\infty$.
The asymptotic normality of $\widehat{\boldsymbol{\beta}}^{W}$ relies on classical results for M-estimators recalled in the Appendix (see Theorem A.4). Note that since we estimate simultaneously $\beta$ and $\sigma_{\epsilon}^{2}$, an additional condition on the existence of the moments of order four for the covariates is required.
Proposition 2.2.
If model (1) holds and hypothesis $\mathbf{H}_{1}$ is fulfilled, ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{4})<+\infty$, $(\boldsymbol{\beta},\sigma^{2}_{\epsilon})\in\Theta$ and $\Theta$ is a compact set that does not contain $(0,0)$, then, as $n$ tends to infinity,
$\sqrt{n}\left(\begin{pmatrix}\widehat{\boldsymbol{\beta}}^{W}\\ \widehat{\sigma}^{2,W}_{\epsilon}\end{pmatrix}-\begin{pmatrix}\boldsymbol{\beta}\\ \sigma^{2}_{\epsilon}\end{pmatrix}\right)$ $\rightsquigarrow\mathcal{N}\left(0,\boldsymbol{\Gamma}_{W}\right),$
for some covariance matrix $\boldsymbol{\Gamma}_{W}$.
Similarly to $\boldsymbol{\Gamma}_{\beta_{M}}$, the expression of the asymptotic covariance matrix $\boldsymbol{\Gamma}_{W}$ is almost impossible to explicitly derive manually, with the exception of some particularly simple cases.
3 Bootstrapping for confidence intervals
Since, as noted in the previous section, it is complicated to compute explicitly the asymptotic variance matrix of $\widehat{\boldsymbol{\beta}}^{M}$ and $\widehat{\boldsymbol{\beta}}^{W}$, we consider stratified bootstrap approaches in order to build confidence sets for $\boldsymbol{\beta}$.
Our bootstrap procedure takes account of the independence between the different groups $k=1,\ldots,K$, as well as the independence of inputs $(X_{1}^{k},\ldots,X_{d}^{k})$ and output $Y^{k}$ within each group, meaning more formally that, given $G=k$, the joint probability measure $\mathbb{P}^{k}$ of $Y$ and $\mathbf{X}$ is a product measure of the marginal measures $\mathbb{P}^{k}=\mathbb{P}_{Y}^{k}\otimes\mathbb{P}_{\mathbf{X}}^{k}$.
Within each group $k$, we draw, with equal probability and with replacement, $n_{y}^{k}$ observations among $Y_{1}^{k},\ldots,Y_{n_{y}^{k}}^{k}$, We also draw independently, with equal probability and with replacement, $n_{x}^{k}$ observations among $\mathbf{X}_{1}^{k},\ldots,\mathbf{X}_{n_{x}^{k}}^{k}$. We denote by $\mu_{Y}^{k*}$ and by $\boldsymbol{\mu}_{X}^{k*}$ the empirical means and by $\sigma_{Y,k}^{2,*}$ and by $\boldsymbol{\Gamma}_{X}^{k*}$ the empirical variances of $Y$ and $\mathbf{X}$ in these these bootstrap samples. Bootstrapped estimators $\boldsymbol{\beta}^{M,*}$ and $\boldsymbol{\beta}^{W,*}$ of $\boldsymbol{\beta}$ can now be computed by replacing the empirical moments by the bootstrap moments in (4) and (6).
To build confidence sets for the components of $\boldsymbol{\beta}$ based on previous bootstrap procedure, the bootstrap percentile technique, described in Chapter 4 of Shao and Tu, (1995), is simple to use.
It can be noted that our estimators are smooth functions of sample means so that classical bootstrap theory applies (see for example Shao and Tu, (1995), Chapter 3). For simplicity, as in Proposition 2.1, we suppose that $n=n_{y}^{1}=\ldots=n_{y}^{K}=n_{x}^{1}=\ldots=n_{x}^{K}$. Because of the considered experimental design, our global "empirical distribution" is made of products of marginal empirical distributions, the bootstrap for means is almost surely consistent for the Kolmogorov metric, and with Theorem 3.1 in Shao and Tu, (1995) the same result holds for the estimators of $\boldsymbol{\beta}$ considered in this work. Then, the application of Theorem 4.1 in Shao and Tu, (1995) allows to conclude that bootstrap percentile method gives consistent confidence sets for each component of $\boldsymbol{\beta}$.
Proposition 3.1.
Suppose that $\mathbb{E}(Y^{2})<\infty$ and $\mathbb{E}\|\mathbf{X}\|^{2}<\infty$ and hypothesis $\mathbf{H}_{1}$ is fulfilled. Then as $n\to+\infty$, the bootstrap estimator $\boldsymbol{\beta}^{M,*}$ is strongly consistent for $\boldsymbol{\beta}$ in the Kolmogorov metric. Consider a risk $\alpha\in(0,1)$, for each component of $\boldsymbol{\beta}$, the bootstrap percentile approach provides, for a given nominal level $1-\alpha$, a consistent confidence set.
We can also state a similar result for the estimators minimizing the Wasserstein distance, under slightly more restrictive moment conditions and compactness assumption.
Proposition 3.2.
Suppose that the assumptions of Proposition 2.2 as fulfilled. As $n\to+\infty$, the bootstrap estimator $(\boldsymbol{\beta}^{W,*},\sigma_{\epsilon}^{2,W,*})$ is strongly consistent for $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$ in the Kolmogorov metric. Consider a risk $\alpha\in(0,1)$, for each component of $\boldsymbol{\beta}$, the bootstrap percentile approach provides, for a given nominal level $1-\alpha$, a consistent confidence set.
Propositions 3.1 and 3.2 are very important for practical applications since they ensure that even if we are not able to compute the explicit expression of the asymptotic variance of our estimators, we can still construct asymptotic confidence intervals thanks to simple bootstrap procedures.
4 A simulation study
Some simulations are performed in order to evaluate the finite sample performances of the proposed approaches on data that resemble in vivo data originating from real experiments on mice.
4.1 Simulation design
4.1.1 General setting
Let $K\in\mathbb{N}$ be the number of subpopulations, $g_{k}=\{0,1,\dots,k-1\}$ be the level of the grouping variable for subpopulation $k\in\{1,\dots,K\}$, and the number of animals observed per group was choose for simplicity $n=n_{y}^{1}=\cdots=n_{y}^{K}=n_{x}^{1}\cdots=n_{x}^{K}$. For every animal $i\in\{1,\dots,n\}$ and every subpopulation $k\in\{1,\dots,K\}$ we choose the predictor variable $X^{k}_{i}$ in the Gaussian univariate setting:
$X^{k}_{i}\sim\mathcal{N}(\mu_{X}^{k},\sigma_{X,k}^{2})\text{, where }\mu_{X}^{k}=\Delta_{\mu}g_{k}+C_{\mu},$
where $C_{\mu}\in\mathbb{R}$ is the global mean scale parameter, and $\Delta_{\mu}\in\mathbb{R}$ is the parameter that determines the difference between groups in terms of the mean. We simulate the predicted variable independently as follows, with regression parameters $\beta_{0}$ and $\beta_{1}$:
$Y^{k}_{i}=\beta_{0}+\beta_{1}X^{\prime k}_{i}+\epsilon^{k}_{i}\text{, where }X^{\prime k}_{i}\sim X^{k}_{i}\text{ and }\epsilon^{k}_{i}\sim\mathcal{N}(0,\sigma_{\epsilon,k}^{2}).$
It should be noted that $X^{k}_{i}$ and $X^{\prime k}_{i}$ are independent, we thus recreate the situation of the predictor and predicted variables not being simultaneously observed.
The variable sets $X=(X^{k}_{i})_{1\leq i\leq n,1\leq k\leq K}$ and $Y=(Y^{k}_{i})_{1\leq i\leq n,1\leq k\leq K}$ are simulated $N_{sim}\in\mathbb{N}$ times. For each simulation, $N_{boot}\in\mathbb{N}$ bootstrap samples of the size $n$ are generated from $X^{k}=(X^{k}_{i})_{1\leq k\leq K}$ and $Y^{k}=(Y^{k}_{i})_{1\leq k\leq K}$ for each subpopulation $k\in\{1,\dots,K\}$ independently, then moment estimators $\mu_{X}^{k*}$ and $\mu_{Y}^{k*}$ are calculated. Finally, the bootstrap sample-based estimators $\boldsymbol{\beta}^{M,*}=(\beta_{0}^{M,*},\beta_{1}^{M,*})$ and $\boldsymbol{\beta}^{W,*}=(\beta_{0}^{W,*},\beta_{1}^{W,*})$ are calculated. Based on $N_{boot}$ estimates, we calculate 95% confidence intervals using the function quantile of the Python library NumPy. As a result, we obtain $N_{sim}$ confidence intervals for each regression parameter, that we use to calculate the following quantities of interest: coverage rate of the intervals, their average amplitude, and power, i.e. the proportion of intervals not including 0. Along with the bootstrap procedure, we estimate the confidence intervals for the parameter estimators of the regression on the means by group with a naive method, assuming that the deviation of the parameter estimator from the true value divided by the estimator’s standard error follows a Student’s t-distribution:
$\frac{\widehat{\beta}_{j}-\beta_{j}}{SE(\widehat{\beta_{j}})}\rightsquigarrow t_{K-(d+1)}\text{ for }j\in\{0,\dots,d\}.$
4.1.2 Parameter levels
We fix the following parameters throughout all simulations: the regression parameters $\beta_{0}=1$ and $\beta_{1}=2$, the number of simulations $N_{sim}=500$, the number of bootstrap samples $N_{boot}=500$, the location parameters of the predictor variable $\Delta_{\mu}=1$ and $C_{\mu}=10$. Multiple parameters are varied to study their effect. We take the number of animals $n\in\{10,30\}$, in particular to test whether the inference is significantly impaired in case of a small number of animals, which is often the case for real experimental data. We consider the number of groups $K\in\{4,10\}$, 4 being the number of groups that is often observed in real data, and 10 being a higher number that may produce sufficiently good results with the naive approach to approximating confidence intervals with Student distribution. Additionally, the parameter $\sigma^{2}_{X,k}$ can be adjusted to control the extent to which the observations per group can be easily distinguish one from another. We set $\sigma_{X,1}^{2}=\dots=\sigma_{X,K}^{2}$, and $\sigma^{2}_{X,k}\in\{0.75,2\}$, the first value corresponding to lower overlap between groups, and the second to higher overlap. Finally, we introduce an additional parameter $\rho\in\mathbb{R}^{+}$ controlling the variance of the response to the variance of noise ratio, i.e. $\rho=\cfrac{\sigma_{Y,k}}{\sigma_{\epsilon,k}}$. The choice of adjusting the noise to signal ratio instead of the quantity of noise itself through $\sigma_{\epsilon,k}^{2}$ is motivated by the fact that $\sigma_{Y,k}$ depends on $\sigma_{X,k}$, hence the same level of $\sigma_{\epsilon,k}$ cannot be interpreted the same way for different values of $\sigma_{X,k}$. The variance of the noise can be expressed as follows: $\sigma_{\epsilon,k}^{2}=\cfrac{\beta_{1}^{2}\sigma_{X,k}^{2}}{\rho^{2}-1}$. The values of $\rho$ are chosen to correspond to the realistic situation, namely the very noisy case and a slightly less noisy one: $\rho\in\{1.01,1.1\}$. The effect of different values of $\rho$ on the simulated response variable is illustrated on Figure 3.
Figure 3: The effect of different values of $\rho$ on the data, with $K=4$ and $\sigma_{X,1}^{2}=\dots=\sigma_{X,4}^{2}=0.75$. a) Boxplots constructed from the simulated values of $X^{k}_{i}$. b) Boxplots constructed from the simulated values of $Y^{k}_{i}$ with lower relative noise level, i.e. $\rho=1.1$. c) Boxplots constructed from the simulated values of $Y^{k}_{i}$ with higher relative noise level, i.e. $\rho=1.01$.
4.2 Results
The results of the simulation study are presented in Figure 4, namely the obtained coverage rates for the linear regression slope confidence intervals on Figure 3(a), the amplitudes on Figure 3(b), and the associated test powers on Figure 3(c). Every figure contains four tables, each corresponding to a combination of parameters with respect to the number of animals $N_{anim}$ and the number of groups $K$. The columns of the tables represent combinations of parameters with respect to the group dispersion parameter $\sigma^{2}_{X,k}$ and the noise to signal ratio $\rho$. Finally, the lines of each table indicate the method used to estimate the confidence intervals: the proposed method of moments and optimal transport based bootstrap estimators, and the classical linear regression procedure based on estimated means within each group.
Figure 4: a) Coverage rates, b) amplitudes, and c) powers of the confidence intervals for the estimators of $\beta_{1}$ obtained from $N_{sim}$ simulations for each parameter combination. The method labels on the left: "mm (boot)" stands for the method of moments with bootstrap, "ot (boot)" for the optimal transport method with bootstrap, and "mm (student)" for the naive linear regression on means approach based on Student’s distribution.
In general, it can be observed that the bootstrap estimators produce confidence intervals with smaller amplitudes and with higher power, but with lower coverage rate. Whereas the coverage rates in all cases remain within 91-96%, the extent to which the amplitudes are smaller and the powers are bigger is significantly more important for the bootstrap estimators in almost all cases. This implies that the naive approach based on the Student’s distribution is more likely to produce false negatives in terms of significance.This trend of further amplified by the number of groups parameter: whereas the results are overall worsened with the decrease in either the number of animals or groups, it is the case with the small number of groups that demonstrates the biggest difference in the approaches. Indeed, in almost all cases within the tables with $K=4$ we observe the amplitudes approximately twice as important for the naive approach, and a similar trend in terms of lower powers. The latter result is important since lower power implies bigger probability of not detecting a significant relationship between the predictor and the predicted variables, when it is actually present. Overall, these results mean that the proposed bootstrap estimators are more effective if the experimental design entails a small number of groups.
Concerning the remaining two parameters, as expected, in general the best results are obtained with lower $\sigma^{2}_{X,k}$ and higher $\rho$. In most case, the results for the naive and the bootstrap estimators are either both good or both bad in terms of power, with the latter being slightly better. A particularly complicated case can be distinguished, with high dispersion and high noise level, with few groups and few animals, where all estimators drastically fail: we observe almost equally bad powers (0.1 for both bootstrap estimators and 0.09 for the naive estimator), despite the significant difference in amplitudes. On the other hand, we can also distinguish two case where the powers of the bootstrap estimators are more than 90%, whereas those of the naive estimator are under 50%: in both cases there are 4 groups and high noise, in the first case there are only 10 animals but lower dispersion, in the second case high dispersion level is compensated but a higher number of animals. This implies that if the underlying distributions per group are characterized by a reasonable amount of overlap, or a significant overlap is compensated by having more observations, the bootstrap estimators manage to detect the significant relationship in most cases, unlike the naive estimator.
Lastly, it can be observed that the estimator based on optimal transport produces confidence intervals with slightly smaller amplitudes compared to the method of moments estimator. The difference appears to be relatively more important in the cases with higher group overlap $\sigma^{2}_{X,k}=2$. However, the powers are not affected by this difference. This may be explained by a more important bias associated with the optimal transport estimator. The estimator will likely produce better results in terms of the power than the method of moments estimator if the bias is corrected.
5 Application to real data
In order to illustrate the proposed estimators on real data, we studied the data from experiments conducted on mice in order to assess the adverse effects induced in the context of different irradiated volume. In the course of these experiments, mice were exposed to either stereotactic body radiation therapy (SBRT) with different beam sizes at 90 Gy on the left lung, or whole-thorax irradiation (WTI) at 19 Gy. For one cohort, the expressions of three pro-inflammatory genes (IL1 $\alpha$ and $\beta$, IL6 and TNF) were measured, for the other cohort the measurements of the thickness of the alveolar septas were made due to its role as a macroscopic biomarker of radio-induced pulmonary lesions. In the case of the SBRT, the measurements were taken in multiple locations: the irradiated patch (inside the irradiation field), the remaining part of the left lung referred to as ipsilateral lung, and the right lung (contralateral lung). The goal of this statistical analysis is to determine whether there is a statistical association between the gene expression as predictors and the septal thickening as as outcome. Our approach is applied since the variables are measured on different animals, but within each irradiation condition there are shared groups in terms of the measurement time points.
Table 1: Results of estimation of the linear regression slope predicting septal thickening with the pro-inflammatory genes expression, with three methods, for WTI and SBRT with different beam sizes, with measurements taken in different parts of lungs.
Method of moments Optimal transport Lin. Reg. on means
(bootstrap) (bootstrap) (Student)
Loc. Vol. Gene ${\widehat{\beta}_{1}}$ 95% C.I. Signif. ${\widehat{\beta}_{1}}$ 95% C.I. Signif. ${\widehat{\beta}_{1}}$ 95% C.I. Signif.
IL1b -0.35 (-1.03, 0.4) -0.19 (-0.44, 0.23) -0.35 (-1.99, 1.28)
1 mm IL6 0.43 (-0.23, 1.2) 0.2 (-0.2, 0.84) 0.43 (-1.32, 2.17)
Ipsilateral TNF 0.2 (-0.23, 0.84) 0.16 (-0.18, 0.84) 0.2 (-1.26, 1.66)
lung IL1b 0.9 (0.0, 2.11) 1.05 (0.01, 1.44) 0.9 (-0.43, 2.24)
3 mm IL6 0.05 (-0.34, 0.46) 0.06 (-0.33, 0.49) 0.05 (-1.65, 1.76)
TNF 0.65 (-0.12, 1.6) 0.63 (-0.12, 1.46) 0.65 (-1.57, 2.87)
IL1b 2.31 (-1.73, 3.87) 0.82 (-0.48, 0.94) 2.31 (-0.73, 5.35)
1 mm IL6 0.85 (-0.7, 2.38) 0.61 (-0.56, 1.6) 0.85 (-3.75, 5.45)
Irradiated TNF 0.85 (-0.6, 2.3) 0.69 (-0.54, 1.53) 0.85 (-3.96, 5.66)
patch IL1b 2.85 (0.84, 5.14) 2.59 (0.73, 3.58) 2.85 (-1.07, 6.78)
3 mm IL6 1.35 (0.22, 2.47) 1.3 (0.21, 2.26) 1.35 (-1.8, 4.5)
TNF 3.81 (1.01, 6.37) 3.37 (0.99, 5.33) 3.81 (-1.86, 9.48)
IL1b -0.92 (-2.2, 1.06) -0.42 (-0.89, 0.51) -0.92 (-3.97, 2.14)
0 mm IL6 2.23 (-1.99, 2.28) 0.83 (-0.83, 0.97) 2.23 (-2.65, 7.12)
TNF 2 (-2.0, 2.82) 0.88 (-1.0, 1.11) 2 (-1.72, 5.72)
IL1b 0.63 (-0.82, 2.32) 0.23 (-0.33, 0.56) 0.63 (-2.95, 4.21)
1 mm IL6 1.03 (-0.57, 2.19) 0.46 (-0.3, 1.0) 1.03 (-0.88, 2.94)
Right TNF 1.05 (-0.47, 2.43) 0.66 (-0.31, 1.26) 1.05 (-1.01, 3.11)
lung IL1b 1.07 (0.17, 1.97) 0.92 (0.17, 1.13) 1.07 (-1.96, 4.1)
3 mm IL6 0.3 (-1.45, 1.12) 0.37 (-0.74, 0.86) 0.3 (-4.94, 5.53)
TNF 2.02 (0.27, 4.1) 1.05 (0.04, 1.44) 2.02 (0.03, 4.02)
IL1b 6.41 (-11.6, 15.4) 1.62 (-2.28, 3.1) 6.41 (-16.7, 29.52)
7 mm IL6 -0.58 (-1.26, 0.09) -0.6 (-1.23, 0.1) -0.58 (-6.23, 5.08)
TNF 0.51 (-1.7, 3.86) 0.43 (-1.45, 1.89) 0.51 (-23.21, 24.22)
Whole IL1a 2.93 (-1.18, 6.63) 2.47 (-1.49, 3.02) 2.93 (-11.08, 16.94)
thorax IL6 3.7 (1.53, 5.99) 2.51 (1.39, 3.43) 3.7 (1.11, 6.29)
irradiation TNF 2.35 (0.57, 4.73) 1.67 (0.53, 1.88) 2.35 (-3.86, 8.57)
The linear regression parameters were estimated with three estimators the same way it was done in the simulation study in Section 4.2. The results are presented in Table 1. The focus is placed on estimating the slope parameter $\beta_{1}$ in particular. The table contains the estimations of $\beta_{1}$ as well as the estimated confidence intervals for the slope estimator, and the corresponding test result on the significance of the estimated relationship.
Figure 5: Distribution of the data, collected from the irradiated patch under SBRT with 7 mm beam size: the expression of the gene TNF on the left, and septal thickness on the right. The measurements were made 30, 45 and 60 days after irradiation.
Major differences between the results for the naive and the bootstrap estimators can be observed in terms of the detected significance. On the one hand, the relationship between the pro-inflammatory genes and septal thickening has been identified by all methods in the case of whole-thorax irradiation (two out of three genes for the bootstrap estimators, and only one for the naive estimator), which is an expected result. On the other hand, we also expect to identify a strong correlation in the case of the measurements taken directly from the irradiated patch. This is only the case for the bootstrap estimators, but not for the naive one. This results is in accordance with the results we obtain with simulated data: the confidence intervals are often overestimated with the naive approach, which may result in false negatives in terms of significance.
For the irradiated patch, as well as in all other cases with identified significance, it is only the case for the beam size of 3 mm, the results that is consistent with literature, indicating it as the beam size starting with which the long-term lesions start appearing (Bertho et al.,, 2020). Multiple significant associations have been identified with the bootstrap estimators in the ipsilateral lung and in the right lung for the beam size of 3 mm. However, none of the genes has been identified as significantly linked to lesions in the case of 7 mm. This may be caused by the relationship being of non-linear nature, as appears to be the case for the gene TNF, presented in Figure 5.
Finally, among the cases where a significant relationship has been detected, the estimated values of the slope are always positive, which indicates a general radio-induced up-regulation trend. These values are in general bigger in case of the whole-thorax irradiation and within the patch than for the ipsilateral or right lung for SBRT, which is also in line with biological knowledge, since the genes are expected to be more strongly up-regulated in the areas of injury than further away. This effect is illustrated in Figure 6 on the example of the linear model prediction made for the gene IL6 .
Figure 6: Linear model prediction of septal thickness based on IL6 expression, plotted for different locations and beam sizes, with the results from two bootstrap estimators.
6 Discussion
This work focuses on a statistical framework designed for extracting dependencies form in vivo experiments, specifically introducing linear regression estimators in the context where the predictor and the predicted variables are never jointly observed. In this work we chose the basic linear multivariate setting, prioritizing simplicity and computational feasibility. Particularly, the estimator based on the method of moments makes no hypotheses on data distribution and can be calculated explicitly. The estimator based on optimal transport includes a simple optimization problem, and is based on the Gaussian form of the Wasserstein distance but does not technically require the data to be Gaussian, seeking to approximate them with Gaussian variables in whatever case.
However, these approaches are inapplicable in the cases where linear relationship hypothesis cannot be satisfied. For instance, it is the case with predicting survival data with some continuous biomarker, which is of particular interest in research into radio-induced adverse effects. To be able to consider such scenarios, our model can be extended to a more general case, namely with generalized linear model. The optimal transport estimator appears promising in this context given the fact that Wasserstein distance allows to compare probability distributions of different nature (for example, continuous and discrete).
Another further research direction lies in investigating alternative methods based on integrated likelihood and Bayesian approaches, which are likely to produce better results in many cases but require putting priors on distributions. Concerning a likelihood-based approach, the joint likelihood to be maximized can be written in a factorized form due to the independence of $X$ and $Y$: $L(x,y;\beta_{0},\beta_{1},\mu_{x}^{k},\sigma_{\epsilon},\sigma_{x})=L_{X}(x;\mu_{x},\sigma_{x})L_{Y}(y;\beta_{0},\beta_{1},\mu_{x},\sigma_{\epsilon},\sigma_{x})$ for a group $k$ in the univariate setting. The marginal likelihoods can be expressed explicitly, assuming that the model (1) is true, and that $X|(G=k)\sim\mathcal{N}(\mu_{x}^{k},\sigma_{x,k}^{2})$, and thus $Y|(G=k)\sim\mathcal{N}(\beta_{0}+\beta_{1}\mu_{x}^{k},\sigma_{\epsilon}^{2}+\beta_{1}^{2}\sigma_{x,k}^{2})$, with the noise $\epsilon\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$ independent of $X$.
Finally, it would be of interest to work on improving theoretical properties of our estimators, namely correcting the negative bias that appears for both estimators. The latter is particularly important in the case of the optimal transport estimator, which may be arising naturally with the Wasserstein distance (Wasserman,, 2023). Correcting this bias would considerably improve the estimator, making it competitive with the approaches mentioned earlier that make numerous assumptions on the data.
Funding
This work is supported by the European Union through the PO FEDER-FSE Bourgogne 2014/2020 programs as part of the project ModBioCan2020, and by Institut de Radioprotection et de Sûreté nucléaire as part of the project ROSIRIS.
Conflict of Interest: none declared.
Acknowledgements: The authors warmly thank Dr. Patrick Tardivel for his numerous remarks that helped to improve the presentation of the manuscript.
Appendix A Some classical theorems in asymptotic statistics
A proof of the classical continuous mapping theorem can be found in van der Vaart, (1998) (Theorem 2.3).
Theorem A.1.
(Continuous mapping theorem).
Let $g:\mathbb{R}^{d}\to\mathbb{R}^{m}$ be continuous at every point of $C$ such that $\mathbb{P}[X\in C]=1$.
If the sequence of random variables $(X_{n})_{n\geq 1}$ converges in distribution (resp. probability, resp. almost surely) to $X$ then $(g(X_{n}))_{n\geq 1}$ converges in distribution (resp. probability, resp. almost surely) to $g(X)$.
We also recall some well known results that are useful to show the consistency of estimators $\widehat{\theta}_{n}$ defined as the minimizers of functionals $Q_{n}(\theta)$ which have some regularity properties at the limit.
Theorem A.2.
(Lemma 2.9 in Newey and McFadden, (1994))
Suppose that $\theta\in\Theta$ and $\Theta$ is compact, $Q_{0}(\theta)$ is continuous and $\forall\theta\in\Theta$, $Q_{n}(\theta)\to Q_{0}(\theta)$ in probability as $n$ tends to infinity. If there is $\alpha>0$ and $B_{n}=O_{p}(1)$ such that
$\forall(\widetilde{\theta},\theta)\in\Theta\times\Theta,\ |Q_{n}(\widetilde{\theta})-Q_{n}(\theta)|\leq B_{n}\|\widetilde{\theta}-\theta\|^{\alpha}$
then
$\sup_{\theta\in\Theta}\left|Q_{n}(\theta)-Q_{0}(\theta)\right|\to 0\ \mbox{ inprobability.}$.
Theorem A.3.
(Theorem 2.1 in Newey and McFadden, (1994))
Suppose that $\theta\in\Theta$ and $\Theta$ is compact, $Q_{0}(\theta)$ is continuous $\forall\theta\in\Theta$. If $Q_{0}(\theta)$ is uniquely maximized at $\theta_{0}$ and, as $n$ tends to infinity, $\sup_{\theta\in\Theta}\left|Q_{n}(\theta)-Q_{0}(\theta)\right|\to 0$ in probability, then $\widehat{\theta}_{n}\to\theta_{0}$ in probability.
Under additional hypotheses, we also get the asymptotic normality of the sequence of estimators $\widehat{\theta}_{n}$ of $\theta_{0}$. we denote by $\nabla_{00}Q_{n}(\theta)$ the Hessian matrix of functional $Q_{n}$ evaluated at $\theta$.
Theorem A.4.
(Theorem 3.1 in Newey and McFadden, (1994))
Suppose that $\widehat{\theta}_{n}\to\theta_{0}$ in probability, (i) $\theta_{0}$ is an interior point of $\Theta$, (ii) $Q_{n}(\theta)$ is twice differentiable in a neighborhood $\mathcal{N}$ of $\theta_{0}$, (iii) $\sqrt{n}\nabla_{0}Q_{n}(\theta_{0})\rightsquigarrow\mathcal{N}\left(0,\boldsymbol{\Sigma}\right)$, (iv) there is $\mathbf{H}(\theta)$ continuous at $\theta_{0}$ and $\sup_{\theta\in\mathcal{N}}\|\nabla_{00}Q_{n}(\theta)-\mathbf{H}(\theta)\|\to 0$ in probability (v) $\mathbf{H}=\mathbf{H}(\theta_{0})$ is non singular. Then
$\sqrt{n}\left(\widehat{\theta}_{n}-\theta_{0}\right)\rightsquigarrow\mathcal{N}\left(0,\mathbf{H}^{-1}\boldsymbol{\Sigma}\mathbf{H}^{-1}\right)$
We also recall the central limit theorem for bootstrap means (see Theorem 23.4 in van der Vaart, (1998) for a proof).
Theorem A.5.
(CLT for bootstrap means)
Let $X_{1},X_{2},\ldots$ be i.i.d. random vectors with mean $\boldsymbol{\mu}$ and covariance matrix $\boldsymbol{\Gamma}$. Then conditionally on $X_{1},X_{2},\ldots$, for almost every sequence $X_{1},X_{2},\ldots$
$\sqrt{n}\left(\overline{X}^{*}_{n}-\overline{X}_{n}\right)\rightsquigarrow\mathcal{N}\left(0,\boldsymbol{\Gamma}\right)$
where $\overline{X}_{n}$ is the empirical mean and $\overline{X}^{*}_{n}$ is the empirical mean of $n$ independent observations drawn from the empirical distribution.
Appendix B Proofs
Proof.
of Lemma 2.3
First note that the assumptions ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{2})<+\infty$ ensure the existence of $\sigma_{Y}^{2}$ and $\boldsymbol{\Gamma}_{X}$. From the law of large numbers, we have that for all $k\in\{1,\ldots,K\},\ \widehat{\boldsymbol{\mu}}_{1,X}^{k}\to\boldsymbol{\mu}_{X}^{k}$ and $\widehat{\mu}_{Y}^{k}\to\mu_{Y}^{k}$ in probability when $n_{\min}$ tends to infinity. We deduce from the continuous mapping theorem that $\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\widehat{\boldsymbol{\mu}}_{1,X}\to\boldsymbol{\mu}_{1,X}^{\top}\mathbf{w}\boldsymbol{\mu}_{1,X}$ and $\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\mathbf{w}\widehat{\boldsymbol{\mu}}_{Y}\to\boldsymbol{\mu}_{1,X}^{\top}\mathbf{w}\boldsymbol{\mu}_{Y}$ in probability. Under hypothesis $\mathbf{H}_{1}$, the inverse being continuous in a neighborhood of $\boldsymbol{\mu}_{1,X}^{\top}\mathbf{w}\boldsymbol{\mu}_{1,X}$ another application of the continuous mapping theorem gives that $\left(\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\mathbf{w}\widehat{\boldsymbol{\mu}}_{1,X}\right)^{-1}\to\left(\boldsymbol{\mu}_{1,X}^{\top}\mathbf{w}\boldsymbol{\mu}_{1,X}\right)^{-1}$ and
$\widehat{\boldsymbol{\beta}}^{M}=\left(\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\mathbf{w}\widehat{\boldsymbol{\mu}}_{1,X}\right)^{-1}\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\mathbf{w}\widehat{\boldsymbol{\mu}}_{Y}\to\left(\boldsymbol{\mu}_{1,X}^{\top}\mathbf{w}\boldsymbol{\mu}_{1,X}\right)^{-1}\boldsymbol{\mu}_{1,X}^{\top}\mathbf{w}\boldsymbol{\mu}_{Y}=\boldsymbol{\beta}$
in probability as $n_{\min}$ tends to infinity. The law of large numbers gives that $\widehat{\boldsymbol{\Gamma}}_{X}^{2}\to\boldsymbol{\Gamma}_{X}^{2}$ and $\widehat{\sigma}_{Y}^{2}\to\sigma_{Y}^{2}$ in probability and we deduce, with another application of the continuous mapping theorem, that $\widehat{\sigma}_{Y}^{2}-\widehat{\boldsymbol{\beta}}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{2}\widehat{\boldsymbol{\beta}}\to\sigma_{Y}^{2}-\boldsymbol{\beta}^{\top}\boldsymbol{\Gamma}_{X}^{2}\boldsymbol{\beta}=\sigma_{\epsilon}^{2}$ in probability as $n_{\min}$ tends to infinity. ∎
Proof.
of Lemma 2.4
The proof is based on Lemma 2.9 and Theorem 2.1 in Newey and McFadden, (1994), which are recalled in Appendix A. The law of large numbers and the continuous mapping theorem give us that for all $(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\in\Theta$, $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\to\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$ in probability, when $n_{\min}$ tends to infinity. Consider now $(\boldsymbol{\alpha},\sigma_{\alpha}^{2})\in\Theta$. We have,
$\left|(\widehat{\mu}_{Y}^{k}-\gamma_{0}-\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k})^{2}-(\widehat{\mu}_{Y}^{k}-\alpha_{0}-\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k})^{2}\right|$ $=\left|(\boldsymbol{\alpha}-\boldsymbol{\gamma})^{T}\begin{pmatrix}1\\ \widehat{\boldsymbol{\mu}}_{X}^{k}\end{pmatrix}\left(2\widehat{\mu}_{Y}^{k}-(\boldsymbol{\alpha}+\boldsymbol{\gamma})^{T}\begin{pmatrix}1\\ \widehat{\boldsymbol{\mu}}_{X}^{k}\end{pmatrix}\right)\right|$
$\leq\left\lVert\boldsymbol{\alpha}-\boldsymbol{\gamma}\right\rVert A_{n}^{k},$
with Cauchy-Schwarz inequality and $A_{n,k}=\mathcal{O}_{p}(1)$ because $\|\widehat{\boldsymbol{\mu}}_{X}^{k}\|=O_{p}(1)$, $\widehat{\mu}_{Y}^{k}=O_{p}(1)$ and for some constant $C_{1}$ that does not depend on $\boldsymbol{\alpha}$ and $\boldsymbol{\gamma}$, $\|\boldsymbol{\alpha}+\boldsymbol{\gamma}\|\leq C_{1}<\infty$ because $\Theta$ is supposed to be compact. On the other hand, we have
$\left|\left(\widehat{\sigma}_{Y,k}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma_{\gamma}^{2}}\right)^{2}-\left(\widehat{\sigma}_{Y,k}-\sqrt{\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\alpha}_{-0}+\sigma_{\alpha}^{2}}\right)^{2}\right|$
$=\left|\sqrt{\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\alpha}_{-0}+\sigma_{\alpha}^{2}}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma_{\gamma}^{2}}\right|\left(2\widehat{\sigma}_{Y,k}+\sqrt{\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\alpha}_{-0}+\sigma_{\alpha}^{2}}+\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma_{\gamma}^{2}}\right)$
$=\left|\sqrt{\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\alpha}_{-0}+\sigma_{\alpha}^{2}}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma_{\gamma}^{2}}\right|O_{p}(1)$
since $\Theta$ is compact and $\|\widehat{\boldsymbol{\Gamma}}_{X}^{k}\|_{sp}=O_{p}(1)$, where .$\|.\|_{sp}$ denotes the spectral norm. Because $\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\alpha}_{-0}-\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}=\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\left(\boldsymbol{\alpha}_{-0}-\boldsymbol{\gamma}_{-0}\right)+\left(\boldsymbol{\alpha}_{-0}-\boldsymbol{\gamma}_{-0}\right)^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}$ we have, for some constant $C_{2,k}>0$,
$\left|\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\alpha}_{-0}-\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}\right|$ $\leq C_{2,k}\left\|\widehat{\boldsymbol{\Gamma}}_{X}^{k}\right\|_{sp}\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|.$ (8)
Using now the fact that function $x\mapsto\sqrt{x}$ is concave and differentiable, we have for $x>0$ and $y>0$ that $\sqrt{y}\leq\sqrt{x}+\frac{y-x}{2\sqrt{x}}$. Thus, if $y>x>0$ then $0<\sqrt{y}-\sqrt{x}\leq\frac{y-x}{2\sqrt{x}}$ and if $x>y>0$, then $0<\sqrt{x}-\sqrt{y}\leq\frac{x-y}{2\sqrt{y}}$. Consequently, we have $|\sqrt{y}-\sqrt{x}|\leq\frac{|x-y|}{2\min(\sqrt{x},\sqrt{y})}$ and we deduce that,
$\left|\sqrt{\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\alpha}_{-0}+\sigma_{\alpha}^{2}}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma_{\gamma}^{2}}\right|$ $\leq B_{n}^{k}\left(\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|+|\sigma_{\alpha}^{2}-\sigma_{\gamma}^{2}|\right)$ (9)
where $B_{n}^{k}=O_{p}(1)$. Combining previous inequalities, we get
$\left|\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi_{n}(\boldsymbol{\alpha},\sigma_{\alpha}^{2})\right|$ $\leq\left(\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|+|\sigma_{\alpha}^{2}-\sigma_{\gamma}^{2}|\right)\sum_{k=1}^{K}\pi_{k}\left(B_{n}^{k}+A_{n}^{k}\right),$ (10)
with $\sum_{k=1}^{K}\pi_{k}\left(B_{n}^{k}+A_{n}^{k}\right)=O_{p}(1)$. As a result, it can be deduced from Lemma 2.9 in Newey and McFadden, (1994) that
$\sup_{(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\in\Theta}\left|\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\right|\to 0\quad\mbox{ in probability.}$.
We conclude the proof by recalling that $\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$ attains its unique minimum at $(\boldsymbol{\beta},\sigma^{2}_{\epsilon})\in\Theta$ if assumption $\mathbf{H}_{1}$ is fulfilled, so that $(\widehat{\boldsymbol{\beta}}^{W},\widehat{\sigma}^{2,W})\to(\boldsymbol{\beta},\sigma^{2}_{\epsilon})$ in probability in view of Theorem 2.1 in Newey and McFadden, (1994). ∎
Proof.
of Proposition 2.1
The central limit theorem applies directly to the independent sequences of independent random variables $(\mathbf{X}_{1}^{1},\cdots,\mathbf{X}_{n}^{1}),\ldots,(\mathbf{X}_{1}^{K},\cdots,\mathbf{X}_{n}^{K})$ and $(Y_{1}^{1},\cdots,Y_{n}^{1}),,\ldots,(Y_{1}^{K},\cdots,Y_{n}^{K})$ so that, as $n$ tends to infinity
$\sqrt{n}\begin{pmatrix}\widehat{\boldsymbol{\mu}}_{X}^{1}-\boldsymbol{\mu}_{X}^{1}\\ \vdots\\ \widehat{\boldsymbol{\mu}}_{X}^{K}-\boldsymbol{\mu}_{X}^{K}\\ \widehat{\mu}_{Y}^{1}-\mu_{Y}^{1}\\ \vdots\\ \widehat{\mu}_{Y}^{K}-\mu_{Y}^{K}\end{pmatrix}$ $\rightsquigarrow\mathcal{N}\left(0,\boldsymbol{\Gamma}_{\mu}\right)$ (17)
where $\boldsymbol{\Gamma}_{\mu}$ is a block diagonal matrix, with diagonal elements $(\boldsymbol{\Gamma}_{X}^{1},\ldots,\boldsymbol{\Gamma}_{X}^{K},\sigma_{Y,1}^{2},\ldots,\sigma_{Y,K}^{2})$, with $\boldsymbol{\Gamma}_{X}^{k}=\mbox{Var}(\mathbf{X}|G=k)={\mathbb{E}}\left(\mathbf{X}^{k}(\mathbf{X}^{k})^{\top}\right)-\boldsymbol{\mu}_{X}^{k}(\boldsymbol{\mu}_{X}^{k})^{\top}$ and $\sigma_{Y,k}^{2}=\mbox{Var}(Y|G=k)$. Consider the application $g:\mathbb{R}^{dK+K}\to\mathbb{R}^{d+1}$ defined by
$g(\boldsymbol{\mu}_{X}^{1},\ldots,\boldsymbol{\mu}_{X}^{K},\mu_{Y}^{1},\ldots,\mu_{Y}^{K})=\left(\boldsymbol{\mu}_{1,X}^{\top}\mathbf{w}\boldsymbol{\mu}_{1,X}\right)^{-1}\boldsymbol{\mu}_{1,X}^{\top}\mathbf{w}\boldsymbol{\mu}_{Y}.$
Application $g$ is differentiable at $\boldsymbol{\theta}=(\boldsymbol{\mu}_{X}^{1},\ldots,\boldsymbol{\mu}_{X}^{K},\mu_{Y}^{1},\ldots,\mu_{Y}^{K})$, with non null Jacobian matrix denoted by $\mathbf{J}_{\theta}$ (see Chapter 8 and more particularly Theorem 8.3 in Magnus and Neudecker, (2019)). The application of the Delta method (see Theorem 3.1 in van der Vaart, (1998)) permits to get the asymptotic normality convergence result,
$\sqrt{n}\left(\widehat{\boldsymbol{\beta}}^{M}-\boldsymbol{\beta}\right)$ $\rightsquigarrow\mathcal{N}\left(0,\boldsymbol{\Gamma}_{\beta_{M}}\right),$
where $\boldsymbol{\Gamma}_{\beta_{M}}=\mathbf{J}_{\theta}\boldsymbol{\Gamma}_{\mu}\mathbf{J}_{\theta}^{\top}$. ∎
Proof.
of Proposition 2.2
The proof consists in checking the different points of Theorem A.4. Point (i) is satisfied by the hypotheses, and the point (ii) follows directly from the fact $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2})$ is twice-differentiable in a neighborhood of $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$. To show that (iii) is fulfilled, we consider the following expansion, based on the empirical version of the gradient of $\varphi$,
$\nabla\varphi_{n}$ $=\begin{pmatrix}-2\sum_{k=1}^{K}\pi_{k}\left(\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\\ -2\sum_{k=1}^{K}\pi_{k}\left[\left(\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\widehat{\boldsymbol{\mu}}_{X}^{k}+\left(\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-1\right)\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}\right]\\ \sum_{k=1}^{K}\pi_{k}\left(1-\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}\right)\end{pmatrix}$ (21)
Since model (1) holds, $\nabla\varphi=0$ and $\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}=(\widehat{\mu}_{Y}^{k}-{\mu}_{Y}^{k})-\boldsymbol{\beta}_{-0}^{\top}\left(\widehat{\boldsymbol{\mu}}_{X}^{k}-\boldsymbol{\mu}_{X}^{k}\right)$, we thus deduce with (17) the asymptotic normality of the first component of the gradient $\nabla\varphi_{n}$, that is to say $\sqrt{n}\left(-2\sum_{k=1}^{K}\pi_{k}\left(\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\right)$ converges in distribution to a centered Gaussian distribution. As far as the second component is concerned, it can be noted that $\widehat{\boldsymbol{\Gamma}}_{X}^{k}$ converges in probability to $\boldsymbol{\Gamma}_{X}^{k}$ and by the continuous mapping theorem, $\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}\to\sigma_{Y,k}$ in probability. It can also be noted that, under the moment condition ${\mathbb{E}}\left[\|\mathbf{X}\|^{4}|G=k\right]<\infty$, the central limit theorem gives that $\sqrt{n}\left(\widehat{\boldsymbol{\Gamma}}_{X}^{k}-\boldsymbol{\Gamma}_{X}^{k}\right)$ converges in distribution to a centered Gaussian multivariate distribution, and we deduce with the Cramer-Wold device, the continuous mapping theorem and Slutsky’s theorem that the second component of $\nabla\varphi_{n}$ multiplied by $\sqrt{n}$ also in distribution to a centered Gaussian random vector. It is immediate to deduce that the same convergence result holds for the third component, which is to say that $\sqrt{n}\left(\sum_{k=1}^{K}\pi_{k}\left(1-\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}\right)\right)$ converges in distribution to a centered Gaussian random variable. We finally deduce, with the Cramer-Wold device, that (iii) is fulfilled. To prove that (iv) also holds, consider the Hessian matrix of functional $\varphi_{n}$, evaluated at $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$:
$\nabla_{00}\varphi_{n}$ $=\begin{pmatrix}2&2\left(\sum_{k=1}^{K}\pi_{k}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)^{\top}&0\\ 2\sum_{k=1}^{K}\pi_{k}\widehat{\boldsymbol{\mu}}_{X}^{k}&\widehat{\mathbf{H}}(\boldsymbol{\beta}_{-0})&\sum_{k=1}^{K}\pi_{k}\widehat{\sigma}_{Y,k}\left(\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}\right)^{-3/2}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}\\ 0&\sum_{k=1}^{K}\pi_{k}\widehat{\sigma}_{Y,k}\left(\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}\right)^{-3/2}\left(\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}&\frac{1}{2}\sum_{k=1}^{K}\pi_{k}\widehat{\sigma}_{Y,k}\left(\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}\right)^{-3/2}\end{pmatrix},$
where
$\widehat{\mathbf{H}}(\boldsymbol{\beta}_{-0})=$ $2\sum_{k=1}^{K}\pi_{k}\bigg{[}\widehat{\sigma}_{Y,k}\left(\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}\right)^{-3/2}\bigg{[}\left(\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}\right)\left(\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}-\left(\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}\right)\widehat{\boldsymbol{\Gamma}}_{X}^{k}\bigg{]}$
$+\widehat{\boldsymbol{\mu}}_{X}^{k}\left(\widehat{\boldsymbol{\mu}}_{X}^{k}\right)^{\top}+\widehat{\boldsymbol{\Gamma}}_{X}^{k}\bigg{]}.$
By similar arguments as those used to show that $\varphi_{n}(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$ converges in probability to $\varphi(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$, we deduce that $\nabla_{00}\varphi_{n}$ converges in probability to some matrix $\mathbf{H}(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$, defined as follows
$\mathbf{H}(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$ $=\begin{pmatrix}2&2\left(\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}\right)^{\top}&0\\ 2\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}&\mathbf{H}(\boldsymbol{\beta}_{-0})&\sum_{k=1}^{K}\pi_{k}\sigma_{Y,k}\left(\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}\right)^{-3/2}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\\ 0&\sum_{k=1}^{K}\pi_{k}\sigma_{Y,k}\left(\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}\right)^{-3/2}\left(\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}&\frac{1}{2}\sum_{k=1}^{K}\pi_{k}\sigma_{Y,k}\left(\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}\right)^{-3/2}\end{pmatrix}$
where
$\mathbf{H}(\boldsymbol{\beta}_{-0})=$ $2\sum_{k=1}^{K}\pi_{k}\bigg{(}\sigma_{Y,k}\left(\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}\right)^{-3/2}\bigg{[}\left(\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)\left(\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}-\left(\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}\right)\boldsymbol{\Gamma}_{X}^{k}\bigg{]}$
$+\boldsymbol{\mu}_{X}^{k}\left(\boldsymbol{\mu}_{X}^{k}\right)^{\top}+\boldsymbol{\Gamma}_{X}^{k}\bigg{)}.$
We now must check that $\mathbf{H}(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$ is a positive definite matrix. For that we show that at the minimizer value $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$ its determinant is strictly positive. We first note that $\sigma_{Y,k}=\left(\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}\right)^{1/2}$ so that $\sigma_{Y,k}\left(\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}\right)^{-3/2}=\frac{1}{\sigma_{Y,k}^{2}}$ and $\mathbf{H}(\boldsymbol{\beta}_{-0})$ can be written in a simpler form,
$\mathbf{H}(\boldsymbol{\beta}_{-0})=$ $2\sum_{k=1}^{K}\pi_{k}\left[\boldsymbol{\mu}_{X}^{k}\left(\boldsymbol{\mu}_{X}^{k}\right)^{\top}+\frac{1}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\left(\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}\right],$ (22)
which is a definite positive matrix under hypothesis $\mathbf{H}_{1}$. Using a block matrix determinant formula, we have
$\left|\mathbf{H}(\boldsymbol{\beta},\sigma_{\epsilon}^{2})\right|$ $=\begin{vmatrix}2&0\\ 0&\frac{1}{2}\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\end{vmatrix}\ \begin{vmatrix}\mathbf{H}(\boldsymbol{\beta}_{-0})-\mathbf{C}\begin{pmatrix}\frac{1}{2}&0\\ 0&\frac{2}{\sum_{k}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}}\end{pmatrix}\mathbf{C}^{\top}\end{vmatrix}$ (28)
where $\mathbf{C}=\begin{pmatrix}2\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}&\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\end{pmatrix}$ and it only has to be verified that the second determinant at the righthand side of (28) is strictly positive. We now have to show that
$\mathbf{H}(\boldsymbol{\beta}_{-0})-\mathbf{C}\begin{pmatrix}\frac{1}{2}&0\\ 0&\frac{2}{\sum_{k}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}}\end{pmatrix}\mathbf{C}^{\top}=2\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}\left(\boldsymbol{\mu}_{X}^{k}\right)^{\top}-2\left(\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}\right)\left(\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}\right)^{\top}$ (31)
$+2\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\left(\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}-\frac{2}{\sum_{k}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}}\left(\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)\left(\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}$ (32)
is a positive matrix. We can remark that by Cauchy Schwarz inequality, for $\mathbf{u}\in\mathbb{R}^{d}$,
$\mathbf{u}^{\top}\left(\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}\right)\left(\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}\right)^{\top}\mathbf{u}=\left(\sum_{k=1}^{K}\pi_{k}\mathbf{u}^{\top}\boldsymbol{\mu}_{X}^{k}\right)^{2}$ $\leq\sum_{k=1}^{K}\pi_{k}\left(\mathbf{u}^{\top}\boldsymbol{\mu}_{X}^{k}\right)^{2}$
$=\mathbf{u}^{\top}\left(\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}\left(\boldsymbol{\mu}_{X}^{k}\right)^{\top}\right)\mathbf{u}$
using the fact that $\sum_{k}(\sqrt{\pi_{k}})^{2}=1$. It can be noted that if $\mathbf{u}\neq 0$, previous inequality is strict unless $\mathbf{u}^{\top}\boldsymbol{\mu}_{X}^{1}=\cdots=\mathbf{u}^{\top}\boldsymbol{\mu}_{X}^{K}$, which can not happen under hypothesis $\mathbf{H}_{1}$. The second part at the righthand side of (32) is handled the same way. We have
$\mathbf{u}^{\top}\left(\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)\left(\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}\mathbf{u}$ $=\left(\mathbf{u}^{\top}\left(\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)\right)^{2}$
$\leq\sum_{k=1}^{K}\sqrt{\frac{\pi_{k}}{\sigma_{Y,k}^{2}}}^{2}\sum_{k=1}^{K}\left(\sqrt{\frac{\pi_{k}}{\sigma_{Y,k}^{2}}}^{2}\mathbf{u}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{2}$
$=\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\mathbf{u}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\left(\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}\mathbf{u},$
and consequently the determinant of $\mathbf{H}(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$ is strictly positive. To finish the proof, it remains to check that in a neighborhood $\mathcal{N}$ of $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$, we have
$\sup_{(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\in\mathcal{N}}\|\nabla_{00}\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\mathbf{H}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\|\to 0\ \mbox{ in probability.}$.
This is a direct consequence of the continuous mapping theorem, which gives us that for all $(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\in\mathcal{N}$, $\|\nabla_{00}\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\mathbf{H}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\|\to 0$ in probability, and the fact that third order partial derivatives of $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$ are bounded in probability for $(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$ so that Theorem A.2 can apply. ∎
Proof.
of Proposition 3.1
The fact that the bootstrap estimator $\boldsymbol{\beta}^{M,*}$ is strongly consistent for $\boldsymbol{\beta}$ is a direct consequence of Theorem 3.1 in Shao and Tu, (1995), noting that
$\widehat{\boldsymbol{\beta}}^{M}=g(\widehat{\boldsymbol{\mu}}_{X}^{1},\ldots,\widehat{\boldsymbol{\mu}}_{X}^{K},\widehat{\mu}_{Y}^{1},\ldots,\widehat{\mu}_{Y}^{K})$
is a continuously differentiable function of means at $(\boldsymbol{\mu}_{X}^{1},\ldots,\boldsymbol{\mu}_{X}^{K},\mu_{Y}^{1},\ldots,\mu_{Y}^{K})$. The fact that confidence sets based on the percentile approach are consistent is proved by checking the assumptions in Theorem 4.1 (iii) Shao and Tu, (1995), namely the bootstrap estimator $\boldsymbol{\beta}^{M,*}$ is consistent, $\widehat{\boldsymbol{\beta}}^{M}$ is consistent (Lemma 2.3), with asymptotic Gaussian distribution (Proposition 2.1). ∎
Proof.
of Proposition 3.2
We denote by $\boldsymbol{\theta}_{0}=(\boldsymbol{\beta},\sigma_{\epsilon}^{2,W})$ the vector of true parameters, by $\widehat{\boldsymbol{\theta}}=(\boldsymbol{\beta}^{W},\sigma_{\epsilon}^{2})$ the sequence of minimum Wasserstein distance estimators and by $\boldsymbol{\theta}^{*}=(\boldsymbol{\beta}^{W,*},\sigma_{\epsilon}^{2,W,*})$ bootstrap estimators of $\boldsymbol{\theta}_{0}$. The vector of parameters $\boldsymbol{\theta}^{*}$ is the minimizer of functional $\varphi_{n}^{*}$ defined as follows,
$\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2})$ $=\sum_{k=1}^{K}\pi_{k}\left[(\mu_{Y}^{k,*}-\gamma_{0}-\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\mu}_{X}^{k,*})^{2}+\left(\sigma_{Y,k}^{*}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k,*}\boldsymbol{\gamma}_{-0}+\sigma^{2}}\right)^{2}\right].$ (33)
We first show with arguments similar to those employed in the proof of Lemma 2.4, that $\boldsymbol{\theta}^{*}$ is a consistent estimator for $\boldsymbol{\theta}_{0}$, based on the fact that $\varphi_{n}^{*}$ is a smooth function converging to $\varphi$ and the sample mean theorem for bootstrap (see for example Theorem 23.4 in van der Vaart, (1998)). Indeed, we first recall that for all $(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\in\Theta$, $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\to\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$ in probability, when $n_{\min}$ tends to infinity and
$\left|\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\right|$ $\leq\left|\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\right|+\left|\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\right|.$ (34)
Since the bootstrap means converge to the empirical ones we deduce with the continuous mapping theorem that $\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\to\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$ in probability, when $n_{\min}$ tends to infinity, so that $\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\to\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$. We also have, as in (10), where empirical means are replaced by bootstrap means,
$\left|\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi_{n}^{*}(\boldsymbol{\alpha},\sigma_{\alpha}^{2})\right|$ $\leq\left(\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|+|\sigma_{\alpha}^{2}-\sigma_{\gamma}^{2}|\right)\sum_{k=1}^{K}\pi_{k}\left(B_{n}^{k,*}+A_{n}^{k,*}\right),$ (35)
for any $(\boldsymbol{\alpha},\sigma_{\alpha}^{2})\in\Theta$, with $\sum_{k=1}^{K}\pi_{k}\left(B_{n}^{k,*}+A_{n}^{k,*}\right)=O_{p}(1)$. As a result, we deduce from Lemma 2.4, inequality (34) and Lemma 2.9 in Newey and McFadden, (1994) that
$\sup_{(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\in\Theta}\left|\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\right|\to 0\quad\mbox{ in probability.}$.
We conclude that $\boldsymbol{\theta}^{*}\to\boldsymbol{\theta}_{0}$ in probability in view of Theorem 2.1 in Newey and McFadden, (1994). We now prove that $\sqrt{n}\left(\boldsymbol{\theta}^{*}-\widehat{\boldsymbol{\theta}}\right)$ and $\sqrt{n}\left(\widehat{\boldsymbol{\theta}}-\boldsymbol{\theta}_{0}\right)$ have the same asymptotic distribution. By definition of $\widehat{\boldsymbol{\theta}}$ and Taylor expansion we have
$\nabla\varphi_{n}(\widehat{\boldsymbol{\theta}})=\nabla\varphi_{n}(\boldsymbol{\theta}_{0})+\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\left(\widehat{\boldsymbol{\theta}}-\boldsymbol{\theta}_{0}\right)=0$ (36)
where $\overline{\boldsymbol{\theta}}$ belongs, componentwise, to the segment between $\boldsymbol{\theta}_{0}$ and $\widehat{\boldsymbol{\theta}}$. We have a similar expansion for boostrap estimators, as well as
$\nabla\varphi_{n}^{*}(\boldsymbol{\theta}^{*})=\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})+\nabla_{00}^{*}\varphi_{n}(\overline{\boldsymbol{\theta}}^{*})\left(\boldsymbol{\theta}^{*}-\boldsymbol{\theta}_{0}\right)=0$ (37)
where $\overline{\boldsymbol{\theta}}^{*}$ belongs, componentwise, to the segment between $\boldsymbol{\theta}_{0}$ and $\boldsymbol{\theta}^{*}$. Combining (36) and (37), we deduce
$\boldsymbol{\theta}^{*}-\widehat{\boldsymbol{\theta}}$ $=\left(\nabla_{00}^{*}\varphi_{n}(\overline{\boldsymbol{\theta}}^{*})\right)^{-1}\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})-\left(\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\right)^{-1}\nabla\varphi_{n}(\boldsymbol{\theta}_{0})$
$=\left(\left(\nabla_{00}^{*}\varphi_{n}(\overline{\boldsymbol{\theta}}^{*})\right)^{-1}-\left(\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\right)^{-1}\right)\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})+\left(\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\right)^{-1}\left(\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})-\nabla\varphi_{n}(\boldsymbol{\theta}_{0})\right)$ (38)
Noticing that $\nabla_{00}^{*}\varphi_{n}(\overline{\boldsymbol{\theta}}^{*})$ and $\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})$ both tend in probability to the same limit $\mathbf{H}(\boldsymbol{\beta},\sigma^{2}_{\epsilon})$ and we have, with similar arguments as those used in the proof of Proposition 2.2, that $\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})$ is $O_{p}(n^{-1/2})$, it can be deduce that
$\boldsymbol{\theta}^{*}-\widehat{\boldsymbol{\theta}}$ $=\left(\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\right)^{-1}\left(\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})-\nabla\varphi_{n}(\boldsymbol{\theta}_{0})\right)+o_{P}(n^{-1/2}).$ (39)
Using arguments similar to those employed in the expansion of $\nabla\varphi_{n}$ in the proof of Proposition 2.2, we make appear the difference between bootstrap means and empirical means or a differentiable functional of these quantities:
$\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})-\nabla\varphi_{n}(\boldsymbol{\theta}_{0})=\footnotesize\begin{pmatrix}2\sum_{k=1}^{K}\pi_{k}\left((\widehat{\mu}_{Y}^{k}-{\mu}_{Y}^{k,*})-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\left(\widehat{\boldsymbol{\mu}}_{X}-\boldsymbol{\mu}_{X}^{k,*}\right)\right)\\ \begin{pmatrix}2\sum_{k=1}^{K}\pi_{k}\left[\left(\widehat{\mu}_{Y}^{k}-\mu_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\widehat{\boldsymbol{\mu}}_{X}^{k}+\left(\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-1\right)\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}\right]\\ -2\sum_{k=1}^{K}\pi_{k}\left[\left(\mu_{Y}^{k,*}-\mu_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\mu}_{X}^{k,*}\right)\boldsymbol{\mu}_{X}^{k,*}+\left(\frac{\sigma_{Y,k}^{*}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k,*}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-1\right)\boldsymbol{\Gamma}_{X}^{k,*}\boldsymbol{\beta}_{-0}\right]\end{pmatrix}\\ \sum_{k=1}^{K}\left(\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-\frac{\sigma_{Y,k}^{*}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k,*}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}\right)\end{pmatrix},$ (45)
which satisfies the central limit theorem for bootstrap means, or the Delta method for bootstrap estimators (see the Appendix as well as Theorem 23.4 and Theorem 23.5 in van der Vaart, (1998)). Consequently, $\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})-\nabla\varphi_{n}(\boldsymbol{\theta}_{0})$ and $\nabla\varphi_{n}(\boldsymbol{\theta}_{0})-\nabla\varphi(\boldsymbol{\theta}_{0})$ have the same asymptotic distribution. By Slustky’s theorem, the asymptotic distribution of $\sqrt{n}\left(\boldsymbol{\theta}^{*}-\widehat{\boldsymbol{\theta}}\right)$ is thus the same as the asymptotic distribution of $\mathbf{H}(\boldsymbol{\beta},\sigma^{2}_{\epsilon})\sqrt{n}\nabla\varphi_{n}(\boldsymbol{\theta}_{0})$, and we can conclude that $\sqrt{n}\left(\boldsymbol{\theta}^{*}-\widehat{\boldsymbol{\theta}}\right)$ and $\sqrt{n}\left(\widehat{\boldsymbol{\theta}}-\boldsymbol{\theta}_{0}\right)$ have also the same asymptotic Gaussian distribution. ∎

### v3

@@
-Estimating the linear relation between variables that are never jointly observed: an application in in vivo experiments
+Estimating the linear relation between variables that are never jointly observed
 Abstract
-This work is motivated by in vivo experiments in which measurement are destructive so that the variables of interest can never be observed simultaneously when the aim is to estimate the regression coefficients of a linear regression.
-Assuming that the global experiment can be decomposed into sub experiments (corresponding for example to different doses) with distinct first moments, we propose different estimators of the linear regression which take account of that additional information.
-We consider estimators based on moments as well as estimators based optimal transport theory.
-These estimators are proved to be consistent as well as asymptotically Gaussian under weak hypotheses.
-The asymptotic variance has no explicit expression, except in some particular cases, and specific bootstrap approaches are developed to build confidence intervals for the estimated parameter.
-A Monte Carlo study is conducted to assess and compare the finite sample performances of the different approaches.
-1 Motivation and context
+In modern experimental science there is a commonly encountered problem of estimating the coefficients of a linear regression in the context where the variables of interest can never be observed simultaneously.
+Assuming that the global experiment can be decomposed into sub-experiments with distinct first moments, we propose two estimators of the linear regression that take this additional information into account.
+We consider an estimator based on moments, and an estimator based on optimal transport theory.
+These estimators are proven to be consistent as well as asymptotically Gaussian under weak hypotheses.
+The asymptotic variance has no explicit expression, except in some particular cases, for which reason a stratified bootstrap approach is developed to build confidence intervals for the estimated parameters, whose consistency is also shown.
+A simulation study, assessing and comparing the finite sample performances of these estimators, demonstrated the advantages of the bootstrap approach in multiple realistic scenarios.
+An application to in vivo experiments, conducted in the context of studying radio-induced adverse effects on mice, revealed important relationships between the biomarkers of interest that could not be identified with the considered naive approach.
+1 Introduction
 In order to study the effects of a certain treatment on a living organism, in vivo experiments are often conducted.
-In the context of a complex organism response, scientists may be interested in studying multiple variables describing the effect from different perspectives.
-In particular, such variables of interest often include a macroscopic biomarker only available through in vivo data on the one hand, and a microscopic biomarker that can also be observed on a cellular level.
+In the context of a complex organism response, scientists may be interested in studying multiple variables describing the effect on different scales.
+In particular, such variables of interest often include a macroscopic biomarker, only available through in vivo data, and a microscopic biomarker that can also be observed on a cellular level (i.e.
+in vitro).
 The interest in this case lies in predicting the former with the latter.
-For instance, in the context of studying the adverse effects induced by radiotherapy on healthy tissues, the potential outcomes of interest may manifest in the form of lesions, that are quantified in a form a certain macroscopic biomarker, and the levels of a some predictor such as gene expression.
-These measures often require sacrificing the animal.
+For instance, in the context of studying the adverse effects induced by radiotherapy on healthy tissues, the potential outcomes of interest manifest in the form of the severity of macroscopic lesions, and predictors such as gene expression.
+These measures often require sacrificing the animals that they are taken from.
 As a result, the quantities of interest cannot be observed on the same animals.
 Since the goal is to establish relationships between these variables, a problem of statistical data fusion arises.
-In this work we propose an approach to estimate the relationship between the variables that are not simultaneously observed under the experimental setting described above, based on a conditional model assuming linear relationship within every component (experimental condition).
-An estimator derived with the method of moments as well as optimal transport solution using Wasserstein distance are considered for the real data problem in question.
-1.1 Example of an in vivo experiment
-Figure 1: Schematic representation of the design of an in vivo experiment studying the effect of irradiated volume.
-An illustrative example of an in vivo experiment where the variables of interest are never observed simultaneously is presented in Figure 1.
-In this experiment, mice are irradiated on the lungs with different volume, with a goal of studying the role of irradiated volume in the appearing of radio-induced adverse effects.
-The latter are assessed by measuring septal thickening, a macroscopic biomarker of radio-induced adverse effects in the lungs.
-The other variable that is measured with the purpose of predicting the adverse effect related variable is the expression of multiple pro-inflammatory genes.
-As shown in Figure 1, there are two independent cohorts in the study, one is used to measure gene expression, whereas the other for measuring septal thickness.
+Figure 1.1: Schematic representation of the design of an in vivo experiment studying the effect of irradiated volume.
+An illustrative example of an in vivo experiment where the variables of interest are never observed simultaneously is presented in Figure 1.1.
+In this experiment, mice are irradiated on the lungs with different volume, with a goal of studying the role of irradiated volume in the appearance of radio-induced adverse effects.
+The latter are assessed by measuring septal thickening, a histological marker of lung injuries.
+The other variable that is measured with the purpose of predicting the adverse effect-related variable is the expression of multiple pro-inflammatory genes.
+As shown in Figure 1.1, there are two independent cohorts in the study, one is used to measure gene expression, whereas the other for measuring septal thickness.
 This is a results of both measures being of destructive nature, which does not allow them to be taken on the same animals.
-Comparing distributions of the measurements arising from the two cohorts, such as those presented in Figure 2, one may suspect a correlation or even a linear relationship between the variables.
+Comparing distributions of the measurements arising from the two cohorts, such as those presented in Figure 1.2, one may suspect a correlation or even a linear relationship between the variables.
 In order to assert whether such a relationship exists, one has to connect two variables that are never observed simultaneously, which translates into solving a data fusion problem.
 In this example, there are four groups indicating time points (1, 3, 6 and 12 months after irradiation) when the corresponding animals are sacrificed and the measurements are taken.
-Thus, the categorical variable indication the time point, which is known for every observation, can be used as an additional variable in order to link the predictor and the predicted variables between each other.
-Figure 2: Distribution of the data, collected from the irradiated patch under SBRT with 3 mm beam size: the expression of the gene IL6 on the left, and septal thickness on the right.
+Thus, the categorical variable indicating the time point, which is known for every observation, can be used as an additional variable in order to link the predictor and the predicted variables between each other.
+Figure 1.2: Distribution of the data, collected from the irradiated patch under SBRT with 3 mm beam size: the expression of the gene IL6 on the left, and septal thickness on the right.
 The measurements were made 1, 3, 6 and 12 months after irradiation.
-1.2 Existing research
 The task of linking variables that are never jointly observed cannot be approached as a typical missing values problem, since most methods for inference on incomplete data require a sufficient overlap, which is completely absent in the case treated here.
-As a result, any approaches applying related frameworks such as multiple imputation in the context of data fusion are unsuitable for our application.
+As a result, any approaches applying frameworks such as multiple imputation in the context of data fusion are unsuitable for our application.
 For example, Carrig et al., (2015) use multiple imputation to integrate disparate datasets allowing for the absence of the overlap, but requiring a calibration dataset, where all the variables of interest must be jointly observed.
@@
 These methods are designed for linking variables that are not observed simultaneously through covariates, present for both variables of interest.
-This corresponds to the properties of the in vivo data described in Section 1.1.
+This corresponds to the properties of the in vivo data described above.
 However, the covariates in these approaches are random variables, typically continuous, and often assumed to be Gaussian, which is the case in Cudeck, (2000) and Massa and Riccomagno, (2017).
-The groups variable available through in vivo experiments may present in a continuous form, but the presence of such categories as control and sham makes it impossible to assume continuity and normality.
+The grouping variable available through in vivo experiments cannot be represented in a continuous form since categories such as control and sham make it impossible to assume continuity and normality.
 The Bayesian network approaches introduced by Triantafillou et al., (2010) and Tsamardinos et al., (2012), aimed at inferring binary causal relationships between variables, are more suitable for large datasets with a high number of covariates.
@@
 It can be noted that the goal of the aforementioned examples in statistical matching is to group individuals before imputation, which is not necessary in our case since the groups are already known.
-2 Estimating the linear relation between variables that are never jointly observed
-2.1 Problem and presentation of the different identification approaches
+In this work we propose an approach to estimate the relationship between the variables that are not simultaneously observed under the experimental setting described above, based on a conditional model assuming linear relationship within every component, defined by the grouping variable that is assumed to be present.
+An estimator derived with the method of moments as well as optimal transport solution using Wasserstein distance are considered for the real data problem in question.
+Both approaches do not require any overlap between the two cohorts of the experiment, and are based on weak assumptions ensuring model identifiability and moments existence.
+To construct confidence intervals, we propose both asymptotic and bootstrap-based results, the former being often unfeasible in practice, with the latter demonstrating practical advantages in numerous realistic settings.
+2 Identification approaches
 We consider a real random variable $Y$ and a vector of $d$ real valued random regressors $\mathbf{X}=(X_{1},\ldots,X_{d})$ and suppose that the following linear regression hold
-$Y$ $=\beta_{0}+\sum_{j=1}^{d}\beta_{j}X_{j}+\epsilon.$ (1)
+$Y$ $=\beta_{0}+\sum_{j=1}^{d}\beta_{j}X_{j}+\epsilon.$ (2.1)
 […]
@@
 We now present two different approaches developed to identify the vector $\boldsymbol{\beta}=(\beta_{0},\ldots,\beta_{d})$ of unknown regression coefficients and the noise variance $\sigma_{\epsilon}^{2}$, taking account of the additional information given by the discrete variable $G$.
-2.1.1 Moment approach
-A first and simple approach is based on first moments identification. Taking the conditional expectation, given $G=k$, in (1), we have for $k=1,\ldots,K$,
-$\mu_{Y}^{k}$ $=\beta_{0}+\sum_{j=1}^{d}\beta_{j}\mu_{X_{j}}^{k},$ (2)
+2.1 Moment approach
+A first and simple approach is based on first moments identification. Taking the conditional expectation, given $G=k$, in (2.1), we have for $k=1,\ldots,K$,
+$\mu_{Y}^{k}$ $=\beta_{0}+\sum_{j=1}^{d}\beta_{j}\mu_{X_{j}}^{k},$ (2.2)
 since the residual term $\epsilon$ is supposed to satisfy $\mathbb{E}(\epsilon|G=k)=0$ for $k=1,\ldots,K$.
 We denote by $\boldsymbol{\mu}_{1,X}$ the $K\times(d+1)$ design matrix, whose $k$th row is equal to $(1,\boldsymbol{\mu}_{X}^{k\top})$ with $\boldsymbol{\mu}_{X}^{k}=(\mu_{X_{1}}^{k},\cdots,\mu_{X_{d}}^{k})^{\top}$, by $\boldsymbol{\mu}_{Y}$ the $K$ dimensional vector with elements $(\mu_{Y}^{1},\ldots,\mu_{Y}^{K})$.
-The $K$ linear equations in (2) can we written in a matrix way, $\boldsymbol{\mu}_{Y}=\boldsymbol{\mu}_{1,X}\boldsymbol{\beta}$.
+The $K$ linear equations in (2.2) can we written in a matrix way, $\boldsymbol{\mu}_{Y}=\boldsymbol{\mu}_{1,X}\boldsymbol{\beta}$.
 We introduce the following assumption, guaranteeing the identifiability of the model parameters:
@@
 Lemma 2.1.
-If model (1) holds and assumption $\mathbf{H}_{1}$ is fulfilled, $\boldsymbol{\beta}$ is uniquely identified in terms of the conditional first order moments of $\mathbf{X}$ and $Y$ given $G$,
+If model (2.1) holds and assumption $\mathbf{H}_{1}$ is fulfilled, $\boldsymbol{\beta}$ is uniquely identified in terms of the conditional first order moments of $\mathbf{X}$ and $Y$ given $G$,
 $\boldsymbol{\beta}$ $=\left(\boldsymbol{\mu}_{1,X}^{\top}\boldsymbol{\mu}_{1,X}\right)^{-1}\boldsymbol{\mu}_{1,X}^{\top}\boldsymbol{\mu}_{Y}.$
@@
 $\sigma_{\epsilon}^{2}$ $=\sigma_{Y}^{2}-\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}\boldsymbol{\beta}_{-0},$
-where $\sigma_{Y}^{2}$ is the variance of $Y$, $\boldsymbol{\Gamma}_{X}$ is the covariance matrix of $\mathbf{X}$, and $\boldsymbol{\beta}_{-0}=(\beta_{1},\ldots,\beta_{d})$.
+where $\sigma_{Y}^{2}$ is the variance of $Y$, $\boldsymbol{\Gamma}_{X}$ is the covariance matrix of $\mathbf{X}$ with elements $\mbox{Cov}(X_{i},X_{j})=\sum_{k=1}^{K}\mbox{Cov}(X_{i},X_{j}|G=k)\mathbb{P}[G=k]$ for $i$ and $j$ in $\{1,\cdots,d\}$, and $\boldsymbol{\beta}_{-0}=(\beta_{1},\ldots,\beta_{d})$.
 The proof of Lemma 2.1 is direct and thus omitted.
-2.1.2 Optimal transport and minimum Wasserstein distance approach
+2.2 Optimal transport approach
 The second approach is based on optimal transport, in particular on the idea of estimating the linear transformation of the distribution of $\mathbf{X}$ that is the closest to that of $Y$ with respect to the Wasserstein distance (see Panaretos and Zemel, (2019) for a general introduction for statisticians).
@@
 […]
-If we assume that, given $G=k$, $\mathbf{X}$ is a Gaussian random vector and $\epsilon$ is Gaussian, we have, if the linear model (1) holds, that $Y$ is also Gaussian given $G=k$, with expectation $\mu_{Y}^{k}=\beta_{0}+\sum_{j=1}^{d}\beta_{j}\mu_{X_{j}}^{k}$ and variance $\sigma_{\epsilon}^{2}+\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}$, where $\boldsymbol{\Gamma}_{X}^{k}$ is the variance matrix of $\mathbf{X}$ given $G=k$.
+If we assume that, given $G=k$, $\mathbf{X}$ is a Gaussian random vector and $\epsilon$ is Gaussian, we have, if the linear model (2.1) holds, that $Y$ is also Gaussian given $G=k$, with expectation $\mu_{Y}^{k}=\beta_{0}+\sum_{j=1}^{d}\beta_{j}\mu_{X_{j}}^{k}$ and variance $\sigma_{\epsilon}^{2}+\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}$, where $\boldsymbol{\Gamma}_{X}^{k}$ is the variance matrix of $\mathbf{X}$ given $G=k$.
 Thus, the Wasserstein distance between $D_{\gamma}$, the distribution of $\gamma_{0}+\boldsymbol{\gamma}_{-0}^{\top}\mathbf{X}+\epsilon$, and $D_{Y}$, the distribution of $Y$, is equal to
@@
 We introduce the following loss criterion
-$\varphi(\boldsymbol{\gamma},\sigma^{2})$ $=\sum_{k=1}^{K}\pi_{k}\left[(\mu_{Y}^{k}-\gamma_{0}-\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\mu}_{X}^{k})^{2}+\left(\sigma_{Y,k}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma^{2}}\right)^{2}\right].$ (3)
+$\varphi(\boldsymbol{\gamma},\sigma^{2})$ $=\sum_{k=1}^{K}\pi_{k}\left[(\mu_{Y}^{k}-\gamma_{0}-\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\mu}_{X}^{k})^{2}+\left(\sigma_{Y,k}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma^{2}}\right)^{2}\right].$ (2.3)
 which evaluates the weighted Wasserstein distance between $Y$ and a linear combination of the $\mathbf{X}$ variables, contaminated by a Gaussian noise, with variance $\sigma^{2}$.
@@
 Lemma 2.2.
-If model (1) holds and if assumption $\mathbf{H}_{1}$ is fulfilled, $\varphi(\boldsymbol{\gamma},\sigma^{2})$ has its unique minimum at $\boldsymbol{\gamma}=\boldsymbol{\beta}$ and $\sigma^{2}=\sigma_{\epsilon}^{2}$.
-2.2 Sampled data and estimators
+If model (2.1) holds and if assumption $\mathbf{H}_{1}$ is fulfilled, $\varphi(\boldsymbol{\gamma},\sigma^{2})$ has its unique minimum at $\boldsymbol{\gamma}=\boldsymbol{\beta}$ and $\sigma^{2}=\sigma_{\epsilon}^{2}$.
+3 Sampled data and estimators
 We suppose that experiments are made for $K\geq 2$ different groups and that for each group $k$, for $k=1,\ldots,K$, we have two independent samples $(Y^{k}_{1},\ldots,Y^{k}_{n_{y}^{k}})$ and $(X^{k}_{j,1},\ldots,X^{k}_{j,n_{x}^{k}})_{j=1,\ldots,d}$, with sizes $n_{y}^{k}$ and $n_{x}^{k}$.
@@
 We denote by $\widehat{\boldsymbol{\mu}}_{1,X}$ the $K\times(d+1)$ matrix, with the first column consisting of ones, and the rest equal to $\widehat{\boldsymbol{\mu}}_{X}$.
-2.2.1 Moment estimators
+3.1 Moment estimators
 If $\widehat{\boldsymbol{\mu}}_{1,X}$ is full rank, moment estimators of $\boldsymbol{\beta}=(\beta_{0},\beta_{1},\ldots,\beta_{d})$ can be built by considering the empirical counterpart of identification equations given in Lemma 2.1,
-$\widehat{\boldsymbol{\beta}}^{M}$ $=\left(\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\widehat{\boldsymbol{\mu}}_{1,X}\right)^{-1}\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\widehat{\boldsymbol{\mu}}_{Y}$
+$\widehat{\boldsymbol{\beta}}^{M}$ $=\left(\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\widehat{\boldsymbol{\mu}}_{1,X}\right)^{-1}\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\widehat{\boldsymbol{\mu}}_{Y}$ (3.1)
 where $\widehat{\boldsymbol{\mu}}_{Y}=(\widehat{\mu}_{Y}^{1},\ldots,\widehat{\mu}_{Y}^{K})$.
@@
 which is unique if $\widehat{\boldsymbol{\mu}}_{1,X}$ is full rank and defined by
-$\widehat{\boldsymbol{\beta}}^{M}$ $=\left(\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\mathbf{w}\widehat{\boldsymbol{\mu}}_{1,X}\right)^{-1}\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\mathbf{w}\widehat{\boldsymbol{\mu}}_{Y}$ (4)
-with $\mathbf{w}=(w_{1},\ldots,w_{K})$. The first moment estimator considered just before corresponds to equal weights $w_{k}=K^{-1}$.
+$\widehat{\boldsymbol{\beta}}^{M}$ $=\left(\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\mathbf{w}\widehat{\boldsymbol{\mu}}_{1,X}\right)^{-1}\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\mathbf{w}\widehat{\boldsymbol{\mu}}_{Y}$ (3.2)
+with $\mathbf{w}=(w_{1},\ldots,w_{K})$. The first moment estimator considered in (3.1) corresponds to equal weights $w_{k}=K^{-1}$.
 We can then define the following estimator of the noise variance,
-$\widehat{\sigma}_{\epsilon}^{2,M}$ $=\widehat{\sigma}_{Y}^{2}-(\widehat{\boldsymbol{\beta}}_{-0}^{M})^{\top}\widehat{\boldsymbol{\Gamma}}_{X}\widehat{\boldsymbol{\beta}}_{-0}^{M}.$ (5)
-2.2.2 Minimum Wasserstein distance estimators
+$\widehat{\sigma}_{\epsilon}^{2,M}$ $=\widehat{\sigma}_{Y}^{2}-(\widehat{\boldsymbol{\beta}}_{-0}^{M})^{\top}\widehat{\boldsymbol{\Gamma}}_{X}\widehat{\boldsymbol{\beta}}_{-0}^{M}.$ (3.3)
+3.2 Optimal transport estimators
 Estimators of $\boldsymbol{\beta}$ and $\sigma^{2}_{\epsilon}$ based on an optimal transport criterion are derived by minimizing the empirical version $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2})$ of functional $\varphi(\boldsymbol{\gamma},\sigma^{2})$ defined by
-$\varphi_{n}(\boldsymbol{\gamma},\sigma^{2})$ $=\sum_{k=1}^{K}\pi_{k}\left[(\widehat{\mu}_{Y}^{k}-\gamma_{0}-\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k})^{2}+\left(\widehat{\sigma}_{Y,k}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma^{2}}\right)^{2}\right].$ (6)
-Note that in absence of an a priori information on the probability $\pi_{k}$ of observing group $k$, we can set $\pi_{k}=K^{-1}$.
+$\varphi_{n}(\boldsymbol{\gamma},\sigma^{2})$ $=\sum_{k=1}^{K}\pi_{k}\left[(\widehat{\mu}_{Y}^{k}-\gamma_{0}-\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k})^{2}+\left(\widehat{\sigma}_{Y,k}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma^{2}}\right)^{2}\right].$ (3.4)
+Note that in absence of a priori information on the probability $\pi_{k}$ of observing group $k$, we can set $\pi_{k}=K^{-1}$.
 We denote by $(\widehat{\boldsymbol{\beta}}^{W},\widehat{\sigma}^{2,W})$ minimizers of $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2})$ which are obtained with iterative optimization algorithms based on gradient descent.
 The algorithm can be initialized with $(\widehat{\boldsymbol{\beta}}^{M},\widehat{\sigma}^{2,M})$
-2.3 Consistency and asymptotic distribution
+4 Consistency and asymptotic distribution
 To study the asymptotic behavior of the estimators of $\boldsymbol{\beta}$ defined in previous section, we suppose that the number $K$ of groups is kept fixed and that for all groups and all variables $\mathbf{X}$ and $Y$, the number of observations tends to infinity.
@@
 Lemma 4.1.
-If ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{2})<+\infty$, and assumption $\mathbf{H}_{1}$ is fulfilled, the sequence of estimators $(\widehat{\boldsymbol{\beta}}^{M},\widehat{\sigma}_{\epsilon}^{2,M})$ defined by (4) and (5) converges in probability to $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$ when $n_{\min}$ tends to infinity.
+If ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{2})<+\infty$, and assumption $\mathbf{H}_{1}$ is fulfilled, the sequence of estimators $(\widehat{\boldsymbol{\beta}}^{M},\widehat{\sigma}_{\epsilon}^{2,M})$ defined by (3.2) and (3.3) converges in probability to $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$ when $n_{\min}$ tends to infinity.
 For the Wasserstein minimum distance estimator, since there is no explicit expression for the estimators, a compactness assumption is also made to obtain the consistency.
 Lemma 4.2.
-If ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{2})<+\infty$, $(\boldsymbol{\beta},\sigma^{2}_{\epsilon})\in\Theta$ and $\Theta$ is a compact set that does not contain 0, suppose that model (1) holds and hypothesis $\mathbf{H}_{1}$ is fulfilled, then the sequence of estimators $(\widehat{\boldsymbol{\beta}}^{W},\widehat{\sigma}^{2,W}_{\epsilon})$ that minimize (6) converges in probability to $(\beta,\sigma^{2}_{\epsilon})$ when $n_{\min}$ tends to infinity.
+If ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{2})<+\infty$, $(\boldsymbol{\beta},\sigma^{2}_{\epsilon})\in\Theta$ and $\Theta$ is a compact set that does not contain 0, suppose that model (2.1) holds and hypothesis $\mathbf{H}_{1}$ is fulfilled, then the sequence of estimators $(\widehat{\boldsymbol{\beta}}^{W},\widehat{\sigma}^{2,W}_{\epsilon})$ that minimize (3.4) converges in probability to $(\beta,\sigma^{2}_{\epsilon})$ when $n_{\min}$ tends to infinity.
 As far as the asymptotic distribution of the estimators is concerned, and for sake of simplicity and lighter notations, we suppose from now on that the number of experiments is the same for all groups and all variables, that is to say $n=n_{y}^{1}=\cdots n_{y}^{K}=n_{x}^{1}\cdots=n_{x}^{K}$.
@@
 The weak convergence toward a Gaussian distribution presented in Proposition 4.1 would remain true, at the expense of heavier notations and a different asymptotic covariance matrix $\boldsymbol{\Gamma}_{\beta_{M}}$, provided that there exist two constants, $0<c\leq C$ such that
-$0<c\leq\frac{\max(n_{y}^{1},\cdots,n_{y}^{K},n_{x}^{1},\cdots,n_{x}^{K})}{\min(n_{y}^{1},\cdots,n_{y}^{K},n_{x}^{1},\cdots,n_{x}^{K})}\leq C<+\infty,$ (7)
+$0<c\leq\frac{\max(n_{y}^{1},\cdots,n_{y}^{K},n_{x}^{1},\cdots,n_{x}^{K})}{\min(n_{y}^{1},\cdots,n_{y}^{K},n_{x}^{1},\cdots,n_{x}^{K})}\leq C<+\infty,$ (4.1)
 and $\min(n_{y}^{1},\cdots,n_{y}^{K},n_{x}^{1},\cdots,n_{x}^{K})\to\infty$.
@@
 Proposition 4.2.
-If model (1) holds and hypothesis $\mathbf{H}_{1}$ is fulfilled, ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{4})<+\infty$, $(\boldsymbol{\beta},\sigma^{2}_{\epsilon})\in\Theta$ and $\Theta$ is a compact set that does not contain $(0,0)$, then, as $n$ tends to infinity,
+If model (2.1) holds and hypothesis $\mathbf{H}_{1}$ is fulfilled, ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{4})<+\infty$, $(\boldsymbol{\beta},\sigma^{2}_{\epsilon})\in\Theta$ and $\Theta$ is a compact set that does not contain $(0,0)$, then, as $n$ tends to infinity,
 $\sqrt{n}\left(\begin{pmatrix}\widehat{\boldsymbol{\beta}}^{W}\\ \widehat{\sigma}^{2,W}_{\epsilon}\end{pmatrix}-\begin{pmatrix}\boldsymbol{\beta}\\ \sigma^{2}_{\epsilon}\end{pmatrix}\right)$ $\rightsquigarrow\mathcal{N}\left(0,\boldsymbol{\Gamma}_{W}\right),$
@@
 Similarly to $\boldsymbol{\Gamma}_{\beta_{M}}$, the expression of the asymptotic covariance matrix $\boldsymbol{\Gamma}_{W}$ is almost impossible to explicitly derive manually, with the exception of some particularly simple cases.
+5 The particular case of simple linear regression
+To show the difficulty, let us examine the case of simple linear regression, that is to say $d=1$. The following linear model
+$Y$ $=\beta_{0}+\beta_{1}X+\epsilon$
+is supposed to hold and if there are two groups $k$ and $j$ such that $\mu_{X}^{k}\neq\mu_{X}^{j}$, the identification assumption $\mathbf{H}_{1}$ is fulfilled and $\beta_{0}$ and $\beta_{1}$ can be uniquely determined.
+The moment estimators of $\beta_{0}$ and $\beta_{1}$ defined in (3.2) have simple expressions,
+$\widehat{\beta}_{0}$ $=\widehat{\mu}_{Y,w}-\widehat{\beta}_{1}\widehat{\mu}_{X,w}$ (5.1)
+$\widehat{\beta}_{1}$ $=\frac{\sum_{k=1}^{K}w_{k}\widehat{\mu}_{X}^{k}\widehat{\mu}_{Y}^{k}-\widehat{\mu}_{X,w}\widehat{\mu}_{Y,w}}{\sum_{k=1}^{K}w_{k}\left(\widehat{\mu}_{X}^{k}\right)^{2}-\widehat{\mu}_{X,w}^{2}}$ (5.2)
+[…]
+We focus on the asymptotic variance of the estimator $\widehat{\beta}_{1}$ of the slope parameter $\beta_{1}$, which is often the parameter of interest. We have
+Lemma 5.1.
+Suppose that the model (2.1) is true, with $d=1$ and $K\geq 2$.
+If there are two groups $k$ and $j$ such that $\mu_{X}^{k}\neq\mu_{X}^{j}$ the vector $(\beta_{0},\beta_{1})$ is identifiable and, as $n$ tends to infinity,
+$\sqrt{n}\left(\widehat{\beta}_{1}-\beta_{1}\right)$ $\rightsquigarrow\mathcal{N}(0,\sigma^{2}_{\beta_{1}})$
+with
+$\sigma^{2}_{\beta_{1}}$ $=\frac{1}{(\mbox{Var}_{w}(X))^{2}}\sum_{k=1}^{K}w_{k}^{2}\left(\beta_{1}^{2}\sigma^{2}_{X,k}+\sigma^{2}_{Y,k}\right)\left(\mu_{X}^{k}-\mu_{X,w}\right)^{2}$
+and $\mu_{X,w}=\sum_{k=1}^{K}w_{k}\mu_{X}^{k}$ and $\mbox{Var}_{w}(X)=\sum_{k=1}^{K}w_{k}\left(\mu_{X}^{k}-\mu_{X,w}\right)^{2}$.
+Lemma 5.1 shows us that even in a very simple framework (only one covariate), the asymptotic variance of the estimator of the slope $\beta_{1}$ is quite complicated.
+It also reveals that minimizing the asymptotic variance $\widehat{\beta}_{1}$ with respect to the weights $w_{k}$ is not a simple task.
 6 Bootstrapping for confidence intervals
@@
 Within each group $k$, we draw, with equal probability and with replacement, $n_{y}^{k}$ observations among $Y_{1}^{k},\ldots,Y_{n_{y}^{k}}^{k}$, We also draw independently, with equal probability and with replacement, $n_{x}^{k}$ observations among $\mathbf{X}_{1}^{k},\ldots,\mathbf{X}_{n_{x}^{k}}^{k}$.
-We denote by $\mu_{Y}^{k*}$ and by $\boldsymbol{\mu}_{X}^{k*}$ the empirical means and by $\sigma_{Y,k}^{2,*}$ and by $\boldsymbol{\Gamma}_{X}^{k*}$ the empirical variances of $Y$ and $\mathbf{X}$ in these these bootstrap samples.
-Bootstrapped estimators $\boldsymbol{\beta}^{M,*}$ and $\boldsymbol{\beta}^{W,*}$ of $\boldsymbol{\beta}$ can now be computed by replacing the empirical moments by the bootstrap moments in (4) and (6).
+We denote by $\mu_{Y}^{k*}$ and by $\boldsymbol{\mu}_{X}^{k*}$ the empirical means and by $\sigma_{Y,k}^{2,*}$ and by $\boldsymbol{\Gamma}_{X}^{k*}$ the empirical variances of $Y$ and $\mathbf{X}$ in these bootstrap samples.
+Bootstrapped estimators $\boldsymbol{\beta}^{M,*}$ and $\boldsymbol{\beta}^{W,*}$ of $\boldsymbol{\beta}$ can now be computed by replacing the empirical moments by the bootstrap moments in (3.2) and (3.4).
 To build confidence sets for the components of $\boldsymbol{\beta}$ based on previous bootstrap procedure, the bootstrap percentile technique, described in Chapter 4 of Shao and Tu, (1995), is simple to use.
@@
 Proposition 6.2.
-Suppose that the assumptions of Proposition 2.2 as fulfilled.
+Suppose that the assumptions of Proposition 4.2 are fulfilled.
 As $n\to+\infty$, the bootstrap estimator $(\boldsymbol{\beta}^{W,*},\sigma_{\epsilon}^{2,W,*})$ is strongly consistent for $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$ in the Kolmogorov metric.
 Consider a risk $\alpha\in(0,1)$, for each component of $\boldsymbol{\beta}$, the bootstrap percentile approach provides, for a given nominal level $1-\alpha$, a consistent confidence set.
-Propositions 3.1 and 3.2 are very important for practical applications since they ensure that even if we are not able to compute the explicit expression of the asymptotic variance of our estimators, we can still construct asymptotic confidence intervals thanks to simple bootstrap procedures.
-4 A simulation study
-Some simulations are performed in order to evaluate the finite sample performances of the proposed approaches on data that resemble in vivo data originating from real experiments on mice.
+Propositions 6.1 and 6.2 are very important for practical applications since they ensure that even if we are not able to compute the explicit expression of the asymptotic variance of our estimators, we can still construct asymptotic confidence intervals thanks to simple bootstrap procedures.
+7 Simulation study
 7.1 Simulation design
-4.1.1 General setting
-Let $K\in\mathbb{N}$ be the number of subpopulations, $g_{k}=\{0,1,\dots,k-1\}$ be the level of the grouping variable for subpopulation $k\in\{1,\dots,K\}$, and the number of animals observed per group was choose for simplicity $n=n_{y}^{1}=\cdots=n_{y}^{K}=n_{x}^{1}\cdots=n_{x}^{K}$.
-For every animal $i\in\{1,\dots,n\}$ and every subpopulation $k\in\{1,\dots,K\}$ we choose the predictor variable $X^{k}_{i}$ in the Gaussian univariate setting:
-$X^{k}_{i}\sim\mathcal{N}(\mu_{X}^{k},\sigma_{X,k}^{2})\text{, where }\mu_{X}^{k}=\Delta_{\mu}g_{k}+C_{\mu},$
-where $C_{\mu}\in\mathbb{R}$ is the global mean scale parameter, and $\Delta_{\mu}\in\mathbb{R}$ is the parameter that determines the difference between groups in terms of the mean.
-We simulate the predicted variable independently as follows, with regression parameters $\beta_{0}$ and $\beta_{1}$:
-$Y^{k}_{i}=\beta_{0}+\beta_{1}X^{\prime k}_{i}+\epsilon^{k}_{i}\text{, where }X^{\prime k}_{i}\sim X^{k}_{i}\text{ and }\epsilon^{k}_{i}\sim\mathcal{N}(0,\sigma_{\epsilon,k}^{2}).$
+We performed a series of simulations in order to evaluate the finite sample performances of the proposed approaches on data that resemble in vivo data originating from real experiments on mice.
+The number of animals observed per group is chosen for simplicity $n=n_{y}^{1}=\cdots=n_{y}^{K}=n_{x}^{1}\cdots=n_{x}^{K}$, and thus the weights for every group are also considered to be equal, i.e.
+$w_{1}=\cdots=w_{K}=\frac{1}{K}$.
+For every animal $i\in\{1,\dots,n\}$ and every subpopulation $k\in\{1,\dots,K\}$, the predictor variable $X^{k}_{i}$ is Gaussian univariate: $X^{k}_{i}\sim\mathcal{N}(\mu_{X}^{k},\sigma_{X}^{2})$, where $\mu_{X}^{k}=9+k$.
+We simulate the predicted variable independently as $Y^{k}_{i}=\beta_{0}+\beta_{1}X^{\prime k}_{i}+\epsilon^{k}_{i}$, where $X^{\prime k}_{i}\sim X^{k}_{i}$ and $\epsilon^{k}_{i}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$, with regression parameters $\beta_{0}=1$ and $\beta_{1}=2$.
 It should be noted that $X^{k}_{i}$ and $X^{\prime k}_{i}$ are independent, we thus recreate the situation of the predictor and predicted variables not being simultaneously observed.
-The variable sets $X=(X^{k}_{i})_{1\leq i\leq n,1\leq k\leq K}$ and $Y=(Y^{k}_{i})_{1\leq i\leq n,1\leq k\leq K}$ are simulated $N_{sim}\in\mathbb{N}$ times.
-For each simulation, $N_{boot}\in\mathbb{N}$ bootstrap samples of the size $n$ are generated from $X^{k}=(X^{k}_{i})_{1\leq k\leq K}$ and $Y^{k}=(Y^{k}_{i})_{1\leq k\leq K}$ for each subpopulation $k\in\{1,\dots,K\}$ independently, then moment estimators $\mu_{X}^{k*}$ and $\mu_{Y}^{k*}$ are calculated.
+The variable sets $X=(X^{k}_{i})_{1\leq i\leq n,1\leq k\leq K}$ and $Y=(Y^{k}_{i})_{1\leq i\leq n,1\leq k\leq K}$ are simulated $N_{sim}$ times.
+For each simulation, $N_{boot}$ bootstrap samples of the size $n$ are generated from $X^{k}=(X^{k}_{i})_{1\leq k\leq K}$ and $Y^{k}=(Y^{k}_{i})_{1\leq k\leq K}$ for each subpopulation $k\in\{1,\dots,K\}$ independently, then moment estimators $\mu_{X}^{k*}$ and $\mu_{Y}^{k*}$ are calculated.
 Finally, the bootstrap sample-based estimators $\boldsymbol{\beta}^{M,*}=(\beta_{0}^{M,*},\beta_{1}^{M,*})$ and $\boldsymbol{\beta}^{W,*}=(\beta_{0}^{W,*},\beta_{1}^{W,*})$ are calculated.
@@
 the proportion of intervals not including 0.
-Along with the bootstrap procedure, we estimate the confidence intervals for the parameter estimators of the regression on the means by group with a naive method, assuming that the deviation of the parameter estimator from the true value divided by the estimator’s standard error follows a Student’s t-distribution:
+Along with the bootstrap estimators, we also considered asymptotic confidence intervals in case of the method of moments, obtained by plugging the estimated values of the moments into the expression for the limit distribution presented in Lemma 5.1.
+Additionally, we estimate the confidence intervals for the parameter estimators of the regression on the means by group with a naive method, assuming that the deviation of the parameter estimator from the true value divided by the estimator’s standard error follows a Student’s t-distribution:
 $\frac{\widehat{\beta}_{j}-\beta_{j}}{SE(\widehat{\beta_{j}})}\rightsquigarrow t_{K-(d+1)}\text{ for }j\in\{0,\dots,d\}.$
-4.1.2 Parameter levels
-We fix the following parameters throughout all simulations: the regression parameters $\beta_{0}=1$ and $\beta_{1}=2$, the number of simulations $N_{sim}=500$, the number of bootstrap samples $N_{boot}=500$, the location parameters of the predictor variable $\Delta_{\mu}=1$ and $C_{\mu}=10$.
-Multiple parameters are varied to study their effect.
-We take the number of animals $n\in\{10,30\}$, in particular to test whether the inference is significantly impaired in case of a small number of animals, which is often the case for real experimental data.
-We consider the number of groups $K\in\{4,10\}$, 4 being the number of groups that is often observed in real data, and 10 being a higher number that may produce sufficiently good results with the naive approach to approximating confidence intervals with Student distribution.
-Additionally, the parameter $\sigma^{2}_{X,k}$ can be adjusted to control the extent to which the observations per group can be easily distinguish one from another.
-We set $\sigma_{X,1}^{2}=\dots=\sigma_{X,K}^{2}$, and $\sigma^{2}_{X,k}\in\{0.75,2\}$, the first value corresponding to lower overlap between groups, and the second to higher overlap.
-Finally, we introduce an additional parameter $\rho\in\mathbb{R}^{+}$ controlling the variance of the response to the variance of noise ratio, i.e.
-$\rho=\cfrac{\sigma_{Y,k}}{\sigma_{\epsilon,k}}$.
-The choice of adjusting the noise to signal ratio instead of the quantity of noise itself through $\sigma_{\epsilon,k}^{2}$ is motivated by the fact that $\sigma_{Y,k}$ depends on $\sigma_{X,k}$, hence the same level of $\sigma_{\epsilon,k}$ cannot be interpreted the same way for different values of $\sigma_{X,k}$.
-The variance of the noise can be expressed as follows: $\sigma_{\epsilon,k}^{2}=\cfrac{\beta_{1}^{2}\sigma_{X,k}^{2}}{\rho^{2}-1}$.
-The values of $\rho$ are chosen to correspond to the realistic situation, namely the very noisy case and a slightly less noisy one: $\rho\in\{1.01,1.1\}$.
-The effect of different values of $\rho$ on the simulated response variable is illustrated on Figure 3.
-Figure 3: The effect of different values of $\rho$ on the data, with $K=4$ and $\sigma_{X,1}^{2}=\dots=\sigma_{X,4}^{2}=0.75$.
+Figure 7.1: The effect of different values of $\rho$ on the data, with $K=4$ and $\sigma_{X}^{2}=0.75$.
 a) Boxplots constructed from the simulated values of $X^{k}_{i}$.
@@
 $\rho=1.01$.
+Throughout all simulations, we fix the number of simulations $N_{sim}=500$ and the number of bootstrap samples $N_{boot}=500$.
+Multiple parameters are varied to study their effect.
+We take the number of animals $n\in\{10,30\}$, in particular to test whether the inference is significantly impaired if the measurements are available only for a small number of animals, which is often the case for real experimental data.
+We consider the number of groups $K\in\{4,10\}$, 4 being the number of groups that is often observed in real data, and 10 being a higher number that may produce sufficiently good results with the naive approach to approximating confidence intervals with Student distribution.
+Additionally, the parameter $\sigma^{2}_{X}$ can be adjusted to control the extent to which the observations per group can be easily distinguished one from another.
+We set $\sigma^{2}_{X}\in\{0.75,2\}$, the first value corresponding to lower overlap between groups, and the second to higher overlap.
+Finally, we introduce an additional parameter $\rho\in\mathbb{R}^{+}$ controlling the variance of the response to the variance of noise ratio, i.e.
+$\rho=\frac{\sigma_{Y}}{\sigma_{\epsilon}}$, where $\sigma^{2}_{Y}=Var(Y^{k}_{i})$ for all $i$ and $k$.
+The choice of adjusting the noise to signal ratio instead of the quantity of noise itself through $\sigma_{\epsilon}^{2}$ is motivated by the fact that $\sigma_{Y}$ depends on $\sigma_{X}$, hence the same level of $\sigma_{\epsilon}$ cannot be interpreted the same way for different values of $\sigma_{X}$.
+The variance of the noise can be expressed as follows: $\sigma_{\epsilon}^{2}=\frac{\beta_{1}^{2}\sigma_{X}^{2}}{\rho^{2}-1}$.
+The values of $\rho$ are chosen to correspond to the realistic situation, namely a very noisy case and a slightly less noisy one: $\rho\in\{1.01,1.1\}$.
+The effect of different values of $\rho$ on the simulated response variable is illustrated on Figure 7.1.
 7.2 Results
-The results of the simulation study are presented in Figure 4, namely the obtained coverage rates for the linear regression slope confidence intervals on Figure 3(a), the amplitudes on Figure 3(b), and the associated test powers on Figure 3(c).
-Every figure contains four tables, each corresponding to a combination of parameters with respect to the number of animals $N_{anim}$ and the number of groups $K$.
-The columns of the tables represent combinations of parameters with respect to the group dispersion parameter $\sigma^{2}_{X,k}$ and the noise to signal ratio $\rho$.
-Finally, the lines of each table indicate the method used to estimate the confidence intervals: the proposed method of moments and optimal transport based bootstrap estimators, and the classical linear regression procedure based on estimated means within each group.
-Figure 4: a) Coverage rates, b) amplitudes, and c) powers of the confidence intervals for the estimators of $\beta_{1}$ obtained from $N_{sim}$ simulations for each parameter combination.
-The method labels on the left: "mm (boot)" stands for the method of moments with bootstrap, "ot (boot)" for the optimal transport method with bootstrap, and "mm (student)" for the naive linear regression on means approach based on Student’s distribution.
-In general, it can be observed that the bootstrap estimators produce confidence intervals with smaller amplitudes and with higher power, but with lower coverage rate.
-Whereas the coverage rates in all cases remain within 91-96%, the extent to which the amplitudes are smaller and the powers are bigger is significantly more important for the bootstrap estimators in almost all cases.
-This implies that the naive approach based on the Student’s distribution is more likely to produce false negatives in terms of significance.This trend of further amplified by the number of groups parameter: whereas the results are overall worsened with the decrease in either the number of animals or groups, it is the case with the small number of groups that demonstrates the biggest difference in the approaches.
+The results of the simulation study are presented in Supplementary Figure C.1.
+Firstly, the results for the moment approach in the asymptotic and bootstrap cases are very similar, thus confirming the effectiveness of the bootstrap approach.
+Concerning the comparison between the bootstrap procedure and the naive approach, it can be observed that in general the bootstrap estimators produce confidence intervals with smaller amplitudes and with higher power, at the expense of a slightly lower coverage rate.
+Whereas the empirical coverage rates in all cases are close to the nominal one (95%), the extent to which the amplitudes are smaller and the powers are bigger for the bootstrap estimators is very important in almost all cases.
+This implies that the naive approach based on the Student’s distribution is more likely to produce false negatives in terms of significance.
+This trend is further amplified by the number of groups parameter: whereas the results are overall worsened with the decrease in either the number of animals or groups, it is the case with the small number of groups that demonstrates the biggest difference in the approaches (the case with few groups and animals is presented in Figure 7.2).
 Indeed, in almost all cases within the tables with $K=4$ we observe the amplitudes approximately twice as important for the naive approach, and a similar trend in terms of lower powers.
@@
 Overall, these results mean that the proposed bootstrap estimators are more effective if the experimental design entails a small number of groups.
-Concerning the remaining two parameters, as expected, in general the best results are obtained with lower $\sigma^{2}_{X,k}$ and higher $\rho$.
-In most case, the results for the naive and the bootstrap estimators are either both good or both bad in terms of power, with the latter being slightly better.
-A particularly complicated case can be distinguished, with high dispersion and high noise level, with few groups and few animals, where all estimators drastically fail: we observe almost equally bad powers (0.1 for both bootstrap estimators and 0.09 for the naive estimator), despite the significant difference in amplitudes.
-On the other hand, we can also distinguish two case where the powers of the bootstrap estimators are more than 90%, whereas those of the naive estimator are under 50%: in both cases there are 4 groups and high noise, in the first case there are only 10 animals but lower dispersion, in the second case high dispersion level is compensated but a higher number of animals.
+Figure 7.2: a) Coverage rates, b) amplitudes, and c) powers of the confidence intervals for the estimators of $\beta_{1}$ obtained from 500 simulations, with number of groups $K=4$ and number of animals per group $n=10$.
+The columns of the tables represent combinations of parameters with respect to the group dispersion parameter $\sigma^{2}_{X}$ and the noise to signal ratio $\rho$.
+The lines indicate the method used to estimate the confidence intervals: "mm (asymp)" stands for the method of moments with asymptotic confidence intervals, "mm (boot)" for the method of moments with bootstrap, "ot (boot)" for the optimal transport method with bootstrap, and "mm (student)" for the naive linear regression on means approach based on Student’s distribution.
+Concerning the remaining two parameters, as expected, in general the best results are obtained with lower $\sigma^{2}_{X}$ and higher $\rho$.
+In most cases, the results for the naive and the bootstrap estimators are either both good or both bad in terms of power, with the latter being slightly better.
+A particularly complicated case can be distinguished, with high dispersion, high noise level, few groups and few animals, where all estimators drastically fail: we observe almost equally bad powers (0.1 for both bootstrap estimators and 0.09 for the naive estimator), despite the significant difference in amplitudes.
+On the other hand, we can also distinguish two cases where the powers of the bootstrap estimators are more than 90%, whereas those of the naive estimator are under 50%: in both cases there are 4 groups and high noise, in the first case there are only 10 animals but lower dispersion, in the second case high dispersion level is compensated but a higher number of animals.
 This implies that if the underlying distributions per group are characterized by a reasonable amount of overlap, or a significant overlap is compensated by having more observations, the bootstrap estimators manage to detect the significant relationship in most cases, unlike the naive estimator.
 Lastly, it can be observed that the estimator based on optimal transport produces confidence intervals with slightly smaller amplitudes compared to the method of moments estimator.
-The difference appears to be relatively more important in the cases with higher group overlap $\sigma^{2}_{X,k}=2$.
+The difference appears to be relatively more important in the cases with higher group overlap $\sigma^{2}_{X}=2$.
 However, the powers are not affected by this difference.
@@
 8 Application to real data
-In order to illustrate the proposed estimators on real data, we studied the data from experiments conducted on mice in order to assess the adverse effects induced in the context of different irradiated volume.
+In order to illustrate the proposed estimators on real data, we studied the data previously mentioned in Section 1, obtained from experiments conducted on mice in order to assess the adverse effects induced in the context of different irradiated volume.
 In the course of these experiments, mice were exposed to either stereotactic body radiation therapy (SBRT) with different beam sizes at 90 Gy on the left lung, or whole-thorax irradiation (WTI) at 19 Gy.
-For one cohort, the expressions of three pro-inflammatory genes (IL1 $\alpha$ and $\beta$, IL6 and TNF) were measured, for the other cohort the measurements of the thickness of the alveolar septas were made due to its role as a macroscopic biomarker of radio-induced pulmonary lesions.
+For one mouse cohort, the expressions of tho pro-inflammatory genes (IL6 and TNF) were measured, for the other cohort the measurements of the thickness of the alveolar septas were made due to its role as a severity measure of radio-induced pulmonary lesions.
 In the case of the SBRT, the measurements were taken in multiple locations: the irradiated patch (inside the irradiation field), the remaining part of the left lung referred to as ipsilateral lung, and the right lung (contralateral lung).
-The goal of this statistical analysis is to determine whether there is a statistical association between the gene expression as predictors and the septal thickening as as outcome.
+A control condition is also considered, where the gene expressions are measured without prior radiation exposure.
+The goal of this statistical analysis is to determine whether there is a statistical association between the gene expression as a predictor and the septal thickening as as outcome.
 Our approach is applied since the variables are measured on different animals, but within each irradiation condition there are shared groups in terms of the measurement time points.
-Table 1: Results of estimation of the linear regression slope predicting septal thickening with the pro-inflammatory genes expression, with three methods, for WTI and SBRT with different beam sizes, with measurements taken in different parts of lungs.
+Table 8.1: Results of estimation of the linear regression slope predicting septal thickening with the pro-inflammatory genes expression, with three methods, for control (no irradiation), WTI and SBRT with different beam sizes, with measurements taken in different parts of lungs.
 Method of moments Optimal transport Lin. Reg. on means
@@
 Loc. Vol. Gene ${\widehat{\beta}_{1}}$ 95% C.I. Signif. ${\widehat{\beta}_{1}}$ 95% C.I. Signif. ${\widehat{\beta}_{1}}$ 95% C.I. Signif.
-IL1b -0.35 (-1.03, 0.4) -0.19 (-0.44, 0.23) -0.35 (-1.99, 1.28)
+Control IL6 2.23 (-1.99, 2.28) 0.83 (-0.83, 0.97) 2.23 (-2.65, 7.12)
+TNF 2 (-2.0, 2.82) 0.88 (-1.0, 1.11) 2 (-1.72, 5.72)
 1 mm IL6 0.43 (-0.23, 1.2) 0.2 (-0.2, 0.84) 0.43 (-1.32, 2.17)
 Ipsilateral TNF 0.2 (-0.23, 0.84) 0.16 (-0.18, 0.84) 0.2 (-1.26, 1.66)
-lung IL1b 0.9 (0.0, 2.11) 1.05 (0.01, 1.44) 0.9 (-0.43, 2.24)
-3 mm IL6 0.05 (-0.34, 0.46) 0.06 (-0.33, 0.49) 0.05 (-1.65, 1.76)
+lung 3 mm IL6 0.05 (-0.34, 0.46) 0.06 (-0.33, 0.49) 0.05 (-1.65, 1.76)
 TNF 0.65 (-0.12, 1.6) 0.63 (-0.12, 1.46) 0.65 (-1.57, 2.87)
-IL1b 2.31 (-1.73, 3.87) 0.82 (-0.48, 0.94) 2.31 (-0.73, 5.35)
+1 mm IL6 1.03 (-0.57, 2.19) 0.46 (-0.3, 1.0) 1.03 (-0.88, 2.94)
+Right TNF 1.05 (-0.47, 2.43) 0.66 (-0.31, 1.26) 1.05 (-1.01, 3.11)
+lung 3 mm IL6 0.3 (-1.45, 1.12) 0.37 (-0.74, 0.86) 0.3 (-4.94, 5.53)
+TNF 2.02 (0.27, 4.1) 1.05 (0.04, 1.44) 2.02 (0.03, 4.02)
 1 mm IL6 0.85 (-0.7, 2.38) 0.61 (-0.56, 1.6) 0.85 (-3.75, 5.45)
 Irradiated TNF 0.85 (-0.6, 2.3) 0.69 (-0.54, 1.53) 0.85 (-3.96, 5.66)
-patch IL1b 2.85 (0.84, 5.14) 2.59 (0.73, 3.58) 2.85 (-1.07, 6.78)
-3 mm IL6 1.35 (0.22, 2.47) 1.3 (0.21, 2.26) 1.35 (-1.8, 4.5)
+patch 3 mm IL6 1.35 (0.22, 2.47) 1.3 (0.21, 2.26) 1.35 (-1.8, 4.5)
 TNF 3.81 (1.01, 6.37) 3.37 (0.99, 5.33) 3.81 (-1.86, 9.48)
-IL1b -0.92 (-2.2, 1.06) -0.42 (-0.89, 0.51) -0.92 (-3.97, 2.14)
-0 mm IL6 2.23 (-1.99, 2.28) 0.83 (-0.83, 0.97) 2.23 (-2.65, 7.12)
-TNF 2 (-2.0, 2.82) 0.88 (-1.0, 1.11) 2 (-1.72, 5.72)
-IL1b 0.63 (-0.82, 2.32) 0.23 (-0.33, 0.56) 0.63 (-2.95, 4.21)
-1 mm IL6 1.03 (-0.57, 2.19) 0.46 (-0.3, 1.0) 1.03 (-0.88, 2.94)
-Right TNF 1.05 (-0.47, 2.43) 0.66 (-0.31, 1.26) 1.05 (-1.01, 3.11)
-lung IL1b 1.07 (0.17, 1.97) 0.92 (0.17, 1.13) 1.07 (-1.96, 4.1)
-3 mm IL6 0.3 (-1.45, 1.12) 0.37 (-0.74, 0.86) 0.3 (-4.94, 5.53)
-TNF 2.02 (0.27, 4.1) 1.05 (0.04, 1.44) 2.02 (0.03, 4.02)
-IL1b 6.41 (-11.6, 15.4) 1.62 (-2.28, 3.1) 6.41 (-16.7, 29.52)
-7 mm IL6 -0.58 (-1.26, 0.09) -0.6 (-1.23, 0.1) -0.58 (-6.23, 5.08)
-TNF 0.51 (-1.7, 3.86) 0.43 (-1.45, 1.89) 0.51 (-23.21, 24.22)
-Whole IL1a 2.93 (-1.18, 6.63) 2.47 (-1.49, 3.02) 2.93 (-11.08, 16.94)
-thorax IL6 3.7 (1.53, 5.99) 2.51 (1.39, 3.43) 3.7 (1.11, 6.29)
+Whole thorax IL6 3.7 (1.53, 5.99) 2.51 (1.39, 3.43) 3.7 (1.11, 6.29)
 irradiation TNF 2.35 (0.57, 4.73) 1.67 (0.53, 1.88) 2.35 (-3.86, 8.57)
-The linear regression parameters were estimated with three estimators the same way it was done in the simulation study in Section 4.2.
-The results are presented in Table 1.
+In order to ensure the comparability of the results for different treatment conditions and genes, the data were centered and reduced with respect to the global mean and standard deviation prior to estimation.
+The linear regression parameters were then estimated with three estimators the same way it was done in the simulation study in Section 7.
 The focus is placed on estimating the slope parameter $\beta_{1}$ in particular.
-The table contains the estimations of $\beta_{1}$ as well as the estimated confidence intervals for the slope estimator, and the corresponding test result on the significance of the estimated relationship.
-Figure 5: Distribution of the data, collected from the irradiated patch under SBRT with 7 mm beam size: the expression of the gene TNF on the left, and septal thickness on the right.
-The measurements were made 30, 45 and 60 days after irradiation.
+The results are presented in Table LABEL:tab:vivo_results, containing the estimations of $\beta_{1}$ as well as the estimated confidence intervals for the slope estimator, and the corresponding test result on the significance of the estimated relationship.
 Major differences between the results for the naive and the bootstrap estimators can be observed in terms of the detected significance.
-On the one hand, the relationship between the pro-inflammatory genes and septal thickening has been identified by all methods in the case of whole-thorax irradiation (two out of three genes for the bootstrap estimators, and only one for the naive estimator), which is an expected result.
-On the other hand, we also expect to identify a strong correlation in the case of the measurements taken directly from the irradiated patch.
+On the one hand, the relationship between the pro-inflammatory genes and septal thickening has been identified by all methods in the case of whole-thorax irradiation (both genes for the bootstrap estimators, and only one for the naive estimator), which is an expected result.
+The results are also consistent for all methods in the case without radiation exposure (control), where no significant relationship has been identified, as it is expected.
+On the other hand, we expect to identify a strong correlation in the case of the measurements taken directly from the irradiated patch.
 This is only the case for the bootstrap estimators, but not for the naive one.
 This results is in accordance with the results we obtain with simulated data: the confidence intervals are often overestimated with the naive approach, which may result in false negatives in terms of significance.
-For the irradiated patch, as well as in all other cases with identified significance, it is only the case for the beam size of 3 mm, the results that is consistent with literature, indicating it as the beam size starting with which the long-term lesions start appearing (Bertho et al.,, 2020).
+Among the SBRT irradiation configurations, only the 3mm beam size has shown a significant correlation between inflammatory genes and septal thickening.
+These results are consistent with literature, indicating it as the beam size starting with which the long-term lesions appear (Bertho et al.,, 2020).
 Multiple significant associations have been identified with the bootstrap estimators in the ipsilateral lung and in the right lung for the beam size of 3 mm.
-However, none of the genes has been identified as significantly linked to lesions in the case of 7 mm.
-This may be caused by the relationship being of non-linear nature, as appears to be the case for the gene TNF, presented in Figure 5.
 Finally, among the cases where a significant relationship has been detected, the estimated values of the slope are always positive, which indicates a general radio-induced up-regulation trend.
-These values are in general bigger in case of the whole-thorax irradiation and within the patch than for the ipsilateral or right lung for SBRT, which is also in line with biological knowledge, since the genes are expected to be more strongly up-regulated in the areas of injury than further away.
-This effect is illustrated in Figure 6 on the example of the linear model prediction made for the gene IL6 .
-Figure 6: Linear model prediction of septal thickness based on IL6 expression, plotted for different locations and beam sizes, with the results from two bootstrap estimators.
+These values are in general bigger in case of whole-thorax irradiaion and within the patch than for the ipsilateral or right lung for SBRT, which suggests a stronger correlation between the inflammatory process and lung injuries under high dose/volume irradiation conditions.
+These results, which are in line with the biological knowledge, could not have been established within the framework of classical statistical regression approaches due to non-simultaneous observations.
+This effect is illustrated in Figure 8.1 on the example of the linear model prediction made for the gene IL6.
+Figure 8.1: Linear model prediction of septal thickness based on IL6 expression, plotted for different locations and beam sizes, with the results from two bootstrap estimators.
 9 Discussion
-This work focuses on a statistical framework designed for extracting dependencies form in vivo experiments, specifically introducing linear regression estimators in the context where the predictor and the predicted variables are never jointly observed.
+This work focuses on a statistical framework designed for extracting dependencies from experiments, specifically introducing linear regression estimators in the context where the predictor and the predicted variables are never jointly observed.
 In this work we chose the basic linear multivariate setting, prioritizing simplicity and computational feasibility.
@@
 To be able to consider such scenarios, our model can be extended to a more general case, namely with generalized linear model.
-The optimal transport estimator appears promising in this context given the fact that Wasserstein distance allows to compare probability distributions of different nature (for example, continuous and discrete).
+The optimal transport estimator appears promising in this context given the fact that the Wasserstein distance allows to compare probability distributions of different nature (for example, continuous and discrete).
 Another further research direction lies in investigating alternative methods based on integrated likelihood and Bayesian approaches, which are likely to produce better results in many cases but require putting priors on distributions.
-Concerning a likelihood-based approach, the joint likelihood to be maximized can be written in a factorized form due to the independence of $X$ and $Y$: $L(x,y;\beta_{0},\beta_{1},\mu_{x}^{k},\sigma_{\epsilon},\sigma_{x})=L_{X}(x;\mu_{x},\sigma_{x})L_{Y}(y;\beta_{0},\beta_{1},\mu_{x},\sigma_{\epsilon},\sigma_{x})$ for a group $k$ in the univariate setting.
-The marginal likelihoods can be expressed explicitly, assuming that the model (1) is true, and that $X|(G=k)\sim\mathcal{N}(\mu_{x}^{k},\sigma_{x,k}^{2})$, and thus $Y|(G=k)\sim\mathcal{N}(\beta_{0}+\beta_{1}\mu_{x}^{k},\sigma_{\epsilon}^{2}+\beta_{1}^{2}\sigma_{x,k}^{2})$, with the noise $\epsilon\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$ independent of $X$.
-Finally, it would be of interest to work on improving theoretical properties of our estimators, namely correcting the negative bias that appears for both estimators.
+Finally, it would be of interest to work on improving theoretical properties of our estimators for finite samples, namely correcting the negative bias that appears for both estimators.
 The latter is particularly important in the case of the optimal transport estimator, which may be arising naturally with the Wasserstein distance (Wasserman,, 2023).
 Correcting this bias would considerably improve the estimator, making it competitive with the approaches mentioned earlier that make numerous assumptions on the data.
-Funding
-This work is supported by the European Union through the PO FEDER-FSE Bourgogne 2014/2020 programs as part of the project ModBioCan2020, and by Institut de Radioprotection et de Sûreté nucléaire as part of the project ROSIRIS.
+Software availability
+Python code and the experimental data are freely available at github.com/parsenteva/vivo_lm.
+Acknowledgements
+The authors gratefully acknowledge the funding from the European Union through the PO FEDER-FSE Bourgogne 2014/2020 programs as part of the project ModBioCan2020, and from the Institut de Radioprotection et de Sûreté nucléaire as part of the project ROSIRIS.
+The authors also warmly thank Dr.
+Agnès François and Dr.
+Annaïg Bertho for the experimental data, as well as Dr.
+Patrick Tardivel for his numerous remarks that helped to improve the presentation of the manuscript.
 Conflict of Interest: none declared.
-Acknowledgements: The authors warmly thank Dr. Patrick Tardivel for his numerous remarks that helped to improve the presentation of the manuscript.
 Appendix A Some classical theorems in asymptotic statistics
@@
 The law of large numbers and the continuous mapping theorem give us that for all $(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\in\Theta$, $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\to\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$ in probability, when $n_{\min}$ tends to infinity.
-Consider now $(\boldsymbol{\alpha},\sigma_{\alpha}^{2})\in\Theta$.
-We have,
+Consider now $(\boldsymbol{\alpha},\sigma_{\alpha}^{2})\in\Theta$. We have,
 $\left|(\widehat{\mu}_{Y}^{k}-\gamma_{0}-\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k})^{2}-(\widehat{\mu}_{Y}^{k}-\alpha_{0}-\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k})^{2}\right|$ $=\left|(\boldsymbol{\alpha}-\boldsymbol{\gamma})^{T}\begin{pmatrix}1\\ \widehat{\boldsymbol{\mu}}_{X}^{k}\end{pmatrix}\left(2\widehat{\mu}_{Y}^{k}-(\boldsymbol{\alpha}+\boldsymbol{\gamma})^{T}\begin{pmatrix}1\\ \widehat{\boldsymbol{\mu}}_{X}^{k}\end{pmatrix}\right)\right|$
@@
 Because $\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\alpha}_{-0}-\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}=\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\left(\boldsymbol{\alpha}_{-0}-\boldsymbol{\gamma}_{-0}\right)+\left(\boldsymbol{\alpha}_{-0}-\boldsymbol{\gamma}_{-0}\right)^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}$ we have, for some constant $C_{2,k}>0$,
-$\left|\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\alpha}_{-0}-\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}\right|$ $\leq C_{2,k}\left\|\widehat{\boldsymbol{\Gamma}}_{X}^{k}\right\|_{sp}\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|.$ (8)
+$\left|\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\alpha}_{-0}-\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}\right|$ $\leq C_{2,k}\left\|\widehat{\boldsymbol{\Gamma}}_{X}^{k}\right\|_{sp}\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|.$ (B.1)
 Using now the fact that function $x\mapsto\sqrt{x}$ is concave and differentiable, we have for $x>0$ and $y>0$ that $\sqrt{y}\leq\sqrt{x}+\frac{y-x}{2\sqrt{x}}$.
@@
 Consequently, we have $|\sqrt{y}-\sqrt{x}|\leq\frac{|x-y|}{2\min(\sqrt{x},\sqrt{y})}$ and we deduce that,
-$\left|\sqrt{\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\alpha}_{-0}+\sigma_{\alpha}^{2}}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma_{\gamma}^{2}}\right|$ $\leq B_{n}^{k}\left(\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|+|\sigma_{\alpha}^{2}-\sigma_{\gamma}^{2}|\right)$ (9)
-where $B_{n}^{k}=O_{p}(1)$. Combining previous inequalities, we get
-$\left|\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi_{n}(\boldsymbol{\alpha},\sigma_{\alpha}^{2})\right|$ $\leq\left(\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|+|\sigma_{\alpha}^{2}-\sigma_{\gamma}^{2}|\right)\sum_{k=1}^{K}\pi_{k}\left(B_{n}^{k}+A_{n}^{k}\right),$ (10)
+$\left|\sqrt{\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\alpha}_{-0}+\sigma_{\alpha}^{2}}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma_{\gamma}^{2}}\right|$ $\leq B_{n}^{k}\left(\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|+|\sigma_{\alpha}^{2}-\sigma_{\gamma}^{2}|\right)$ (B.2)
+where $B_{n}^{k}=O_{p}(1)$.
+Combining previous inequalities, we get
+$\left|\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi_{n}(\boldsymbol{\alpha},\sigma_{\alpha}^{2})\right|$ $\leq\left(\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|+|\sigma_{\alpha}^{2}-\sigma_{\gamma}^{2}|\right)\sum_{k=1}^{K}\pi_{k}\left(B_{n}^{k}+A_{n}^{k}\right),$ (B.3)
 with $\sum_{k=1}^{K}\pi_{k}\left(B_{n}^{k}+A_{n}^{k}\right)=O_{p}(1)$. As a result, it can be deduced from Lemma 2.9 in Newey and McFadden, (1994) that
@@
 The central limit theorem applies directly to the independent sequences of independent random variables $(\mathbf{X}_{1}^{1},\cdots,\mathbf{X}_{n}^{1}),\ldots,(\mathbf{X}_{1}^{K},\cdots,\mathbf{X}_{n}^{K})$ and $(Y_{1}^{1},\cdots,Y_{n}^{1}),,\ldots,(Y_{1}^{K},\cdots,Y_{n}^{K})$ so that, as $n$ tends to infinity
-$\sqrt{n}\begin{pmatrix}\widehat{\boldsymbol{\mu}}_{X}^{1}-\boldsymbol{\mu}_{X}^{1}\\ \vdots\\ \widehat{\boldsymbol{\mu}}_{X}^{K}-\boldsymbol{\mu}_{X}^{K}\\ \widehat{\mu}_{Y}^{1}-\mu_{Y}^{1}\\ \vdots\\ \widehat{\mu}_{Y}^{K}-\mu_{Y}^{K}\end{pmatrix}$ $\rightsquigarrow\mathcal{N}\left(0,\boldsymbol{\Gamma}_{\mu}\right)$ (17)
+$\sqrt{n}\begin{pmatrix}\widehat{\boldsymbol{\mu}}_{X}^{1}-\boldsymbol{\mu}_{X}^{1}\\ \vdots\\ \widehat{\boldsymbol{\mu}}_{X}^{K}-\boldsymbol{\mu}_{X}^{K}\\ \widehat{\mu}_{Y}^{1}-\mu_{Y}^{1}\\ \vdots\\ \widehat{\mu}_{Y}^{K}-\mu_{Y}^{K}\end{pmatrix}$ $\rightsquigarrow\mathcal{N}\left(0,\boldsymbol{\Gamma}_{\mu}\right)$ (B.4)
 where $\boldsymbol{\Gamma}_{\mu}$ is a block diagonal matrix, with diagonal elements $(\boldsymbol{\Gamma}_{X}^{1},\ldots,\boldsymbol{\Gamma}_{X}^{K},\sigma_{Y,1}^{2},\ldots,\sigma_{Y,K}^{2})$, with $\boldsymbol{\Gamma}_{X}^{k}=\mbox{Var}(\mathbf{X}|G=k)={\mathbb{E}}\left(\mathbf{X}^{k}(\mathbf{X}^{k})^{\top}\right)-\boldsymbol{\mu}_{X}^{k}(\boldsymbol{\mu}_{X}^{k})^{\top}$ and $\sigma_{Y,k}^{2}=\mbox{Var}(Y|G=k)$.
@@
 To show that (iii) is fulfilled, we consider the following expansion, based on the empirical version of the gradient of $\varphi$,
-$\nabla\varphi_{n}$ $=\begin{pmatrix}-2\sum_{k=1}^{K}\pi_{k}\left(\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\\ -2\sum_{k=1}^{K}\pi_{k}\left[\left(\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\widehat{\boldsymbol{\mu}}_{X}^{k}+\left(\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-1\right)\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}\right]\\ \sum_{k=1}^{K}\pi_{k}\left(1-\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}\right)\end{pmatrix}$ (21)
-Since model (1) holds, $\nabla\varphi=0$ and $\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}=(\widehat{\mu}_{Y}^{k}-{\mu}_{Y}^{k})-\boldsymbol{\beta}_{-0}^{\top}\left(\widehat{\boldsymbol{\mu}}_{X}^{k}-\boldsymbol{\mu}_{X}^{k}\right)$, we thus deduce with (17) the asymptotic normality of the first component of the gradient $\nabla\varphi_{n}$, that is to say $\sqrt{n}\left(-2\sum_{k=1}^{K}\pi_{k}\left(\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\right)$ converges in distribution to a centered Gaussian distribution.
+$\nabla\varphi_{n}$ $=\begin{pmatrix}-2\sum_{k=1}^{K}\pi_{k}\left(\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\\ -2\sum_{k=1}^{K}\pi_{k}\left[\left(\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\widehat{\boldsymbol{\mu}}_{X}^{k}+\left(\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-1\right)\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}\right]\\ \sum_{k=1}^{K}\pi_{k}\left(1-\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}\right)\end{pmatrix}$ (B.5)
+Since model (2.1) holds, $\nabla\varphi=0$ and $\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}=(\widehat{\mu}_{Y}^{k}-{\mu}_{Y}^{k})-\boldsymbol{\beta}_{-0}^{\top}\left(\widehat{\boldsymbol{\mu}}_{X}^{k}-\boldsymbol{\mu}_{X}^{k}\right)$, we thus deduce with (B.4) the asymptotic normality of the first component of the gradient $\nabla\varphi_{n}$, that is to say $\sqrt{n}\left(-2\sum_{k=1}^{K}\pi_{k}\left(\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\right)$ converges in distribution to a centered Gaussian distribution.
 As far as the second component is concerned, it can be noted that $\widehat{\boldsymbol{\Gamma}}_{X}^{k}$ converges in probability to $\boldsymbol{\Gamma}_{X}^{k}$ and by the continuous mapping theorem, $\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}\to\sigma_{Y,k}$ in probability.
@@
 We first note that $\sigma_{Y,k}=\left(\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}\right)^{1/2}$ so that $\sigma_{Y,k}\left(\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}\right)^{-3/2}=\frac{1}{\sigma_{Y,k}^{2}}$ and $\mathbf{H}(\boldsymbol{\beta}_{-0})$ can be written in a simpler form,
-$\mathbf{H}(\boldsymbol{\beta}_{-0})=$ $2\sum_{k=1}^{K}\pi_{k}\left[\boldsymbol{\mu}_{X}^{k}\left(\boldsymbol{\mu}_{X}^{k}\right)^{\top}+\frac{1}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\left(\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}\right],$ (22)
+$\mathbf{H}(\boldsymbol{\beta}_{-0})=$ $2\sum_{k=1}^{K}\pi_{k}\left[\boldsymbol{\mu}_{X}^{k}\left(\boldsymbol{\mu}_{X}^{k}\right)^{\top}+\frac{1}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\left(\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}\right],$ (B.6)
 which is a definite positive matrix under hypothesis $\mathbf{H}_{1}$. Using a block matrix determinant formula, we have
-$\left|\mathbf{H}(\boldsymbol{\beta},\sigma_{\epsilon}^{2})\right|$ $=\begin{vmatrix}2&0\\ 0&\frac{1}{2}\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\end{vmatrix}\ \begin{vmatrix}\mathbf{H}(\boldsymbol{\beta}_{-0})-\mathbf{C}\begin{pmatrix}\frac{1}{2}&0\\ 0&\frac{2}{\sum_{k}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}}\end{pmatrix}\mathbf{C}^{\top}\end{vmatrix}$ (28)
-where $\mathbf{C}=\begin{pmatrix}2\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}&\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\end{pmatrix}$ and it only has to be verified that the second determinant at the righthand side of (28) is strictly positive.
+$\left|\mathbf{H}(\boldsymbol{\beta},\sigma_{\epsilon}^{2})\right|$ $=\begin{vmatrix}2&0\\ 0&\frac{1}{2}\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\end{vmatrix}\ \begin{vmatrix}\mathbf{H}(\boldsymbol{\beta}_{-0})-\mathbf{C}\begin{pmatrix}\frac{1}{2}&0\\ 0&\frac{2}{\sum_{k}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}}\end{pmatrix}\mathbf{C}^{\top}\end{vmatrix}$ (B.7)
+where $\mathbf{C}=\begin{pmatrix}2\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}&\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\end{pmatrix}$ and it only has to be verified that the second determinant at the righthand side of (B.7) is strictly positive.
 We now have to show that
-$\mathbf{H}(\boldsymbol{\beta}_{-0})-\mathbf{C}\begin{pmatrix}\frac{1}{2}&0\\ 0&\frac{2}{\sum_{k}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}}\end{pmatrix}\mathbf{C}^{\top}=2\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}\left(\boldsymbol{\mu}_{X}^{k}\right)^{\top}-2\left(\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}\right)\left(\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}\right)^{\top}$ (31)
-$+2\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\left(\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}-\frac{2}{\sum_{k}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}}\left(\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)\left(\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}$ (32)
+$\mathbf{H}(\boldsymbol{\beta}_{-0})-\mathbf{C}\begin{pmatrix}\frac{1}{2}&0\\ 0&\frac{2}{\sum_{k}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}}\end{pmatrix}\mathbf{C}^{\top}=2\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}\left(\boldsymbol{\mu}_{X}^{k}\right)^{\top}-2\left(\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}\right)\left(\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}\right)^{\top}$
+$+2\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\left(\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}-\frac{2}{\sum_{k}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}}\left(\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)\left(\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}$ (B.8)
 is a positive matrix. We can remark that by Cauchy Schwarz inequality, for $\mathbf{u}\in\mathbb{R}^{d}$,
@@
 It can be noted that if $\mathbf{u}\neq 0$, previous inequality is strict unless $\mathbf{u}^{\top}\boldsymbol{\mu}_{X}^{1}=\cdots=\mathbf{u}^{\top}\boldsymbol{\mu}_{X}^{K}$, which can not happen under hypothesis $\mathbf{H}_{1}$.
-The second part at the righthand side of (32) is handled the same way.
+The second part at the righthand side of (B.8) is handled the same way.
 We have
@@
 ∎
+Proof.
+of Lemma 5.1
+Note that
+$\widehat{\beta}_{1}=g(\widehat{\mu}_{X}^{1},\ldots,\widehat{\mu}_{X}^{K},\widehat{\mu}_{Y}^{1},\ldots,\widehat{\mu}_{Y}^{K}),$
+with $g:\mathbb{R}^{K+K}\to\mathbb{R}$ defined as follows,
+$g(\mu_{X}^{1},\ldots,\mu_{X}^{K},\mu_{Y}^{1},\ldots,\mu_{Y}^{K})$ $=\frac{\sum_{k=1}^{K}w_{k}\mu_{X}^{k}\mu_{Y}^{k}-\mu_{X,w}\mu_{Y,w}}{\sum_{k=1}^{K}w_{k}(\mu_{X}^{k})^{2}-\left(\sum_{k=1}^{K}w_{k}\mu_{X}^{k}\right)^{2}},$
+$=\frac{\mbox{Cov}_{w}(X,Y)}{\mbox{Var}_{w}(X)},$ (B.9)
+with the notations $\mu_{X,w}=\sum_{k=1}^{K}w_{k}\mu_{X}^{k}$, $\mu_{Y,w}=\sum_{j=1}^{K}w_{j}\mu_{Y}^{j}$, $\mbox{Cov}_{w}(X,Y)=\sum_{k=1}^{K}w_{k}\mu_{X}^{k}\mu_{Y}^{k}-\mu_{X,w}\mu_{Y,w}$ and $\mbox{Var}_{w}(X)=\sum_{k=1}^{K}w_{k}(\mu_{X}^{k})^{2}-(\mu_{X,w})^{2}$.
+The gradient $\nabla g$ of $g$, evaluated at point $(\mu_{X}^{1},\ldots,\mu_{X}^{K},\mu_{Y}^{1},\ldots,\mu_{Y}^{K})$, is equal to
+$\nabla g$ $=\begin{pmatrix}\frac{w_{1}(\mu_{Y}^{1}-\mu_{Y,w})}{\mbox{Var}_{w}(X)}-\frac{2w_{1}\left(\mu_{X}^{1}-\mu_{X,w}\right)\mbox{Cov}_{w}(X,Y)}{(\mbox{Var}_{w}(X))^{2}}\\ \vdots\\ \frac{w_{K}(\mu_{Y}^{K}-\mu_{Y,w})}{\mbox{Var}_{w}(X)}-\frac{2w_{K}\left(\mu_{X}^{K}-\mu_{X,w}\right)\mbox{Cov}_{w}(X,Y)}{(\mbox{Var}_{w}(X))^{2}}\\ \frac{w_{1}\left(\mu_{X}^{1}-\mu_{X,w}\right)}{\mbox{Var}_{w}(X)}\\ \vdots\\ \frac{w_{K}\left(\mu_{X}^{K}-\mu_{X,w}\right)}{\mbox{Var}_{w}(X)}\end{pmatrix}.$
+As in the proof of Proposition 4.1, we get that $\sqrt{n}\left(\widehat{\beta}_{1}-\beta_{1}\right)\rightsquigarrow\mathcal{N}(0,\sigma^{2}_{\beta_{1}})$ with $\sigma^{2}_{\beta_{1}}=\nabla g^{T}\Gamma_{\mu}\nabla g$, so that
+$\sigma^{2}_{\beta_{1}}$ $=\frac{1}{(\mbox{Var}_{w}(X))^{2}}\sum_{k=1}^{K}w_{k}^{2}\left[\sigma^{2}_{X,k}\left(\mu_{Y}^{k}-\mu_{Y,w}-2\beta_{1}\left(\mu_{X}^{k}-\mu_{X,w}\right)\right)^{2}+\sigma^{2}_{Y,k}\left(\mu_{X}^{k}-\mu_{X,w}\right)^{2}\right]$
+$=\frac{1}{(\mbox{Var}_{w}(X))^{2}}\sum_{k=1}^{K}w_{k}^{2}\left[\sigma^{2}_{X,k}\left(-\beta_{1}\left(\mu_{X}^{k}-\mu_{X,w}\right)\right)^{2}+\sigma^{2}_{Y,k}\left(\mu_{X}^{k}-\mu_{X,w}\right)^{2}\right]$
+$=\frac{1}{(\mbox{Var}_{w}(X))^{2}}\sum_{k=1}^{K}w_{k}^{2}\left(\mu_{X}^{k}-\mu_{X,w}\right)^{2}\left(\beta_{1}^{2}\sigma^{2}_{X,k}+\sigma^{2}_{Y,k}\right)$ (B.10)
+remarking that $\beta_{1}=\mbox{Cov}_{w}(X,Y)/\mbox{Var}_{w}(X,Y)$, $\beta_{0}=\mu_{Y,w}-\beta_{1}\mu_{X,w}$ as well as $\beta_{0}=\mu_{Y}^{k}-\beta_{1}\mu_{X}^{k}$.
 Proof.
@@
 The vector of parameters $\boldsymbol{\theta}^{*}$ is the minimizer of functional $\varphi_{n}^{*}$ defined as follows,
-$\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2})$ $=\sum_{k=1}^{K}\pi_{k}\left[(\mu_{Y}^{k,*}-\gamma_{0}-\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\mu}_{X}^{k,*})^{2}+\left(\sigma_{Y,k}^{*}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k,*}\boldsymbol{\gamma}_{-0}+\sigma^{2}}\right)^{2}\right].$ (33)
+$\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2})$ $=\sum_{k=1}^{K}\pi_{k}\left[(\mu_{Y}^{k,*}-\gamma_{0}-\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\mu}_{X}^{k,*})^{2}+\left(\sigma_{Y,k}^{*}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k,*}\boldsymbol{\gamma}_{-0}+\sigma^{2}}\right)^{2}\right].$ (B.11)
 We first show with arguments similar to those employed in the proof of Lemma 4.2, that $\boldsymbol{\theta}^{*}$ is a consistent estimator for $\boldsymbol{\theta}_{0}$, based on the fact that $\varphi_{n}^{*}$ is a smooth function converging to $\varphi$ and the sample mean theorem for bootstrap (see for example Theorem 23.4 in van der Vaart, (1998)).
 Indeed, we first recall that for all $(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\in\Theta$, $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\to\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$ in probability, when $n_{\min}$ tends to infinity and
-$\left|\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\right|$ $\leq\left|\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\right|+\left|\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\right|.$ (34)
+$\left|\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\right|$ $\leq\left|\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\right|+\left|\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\right|.$ (B.12)
 Since the bootstrap means converge to the empirical ones we deduce with the continuous mapping theorem that $\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\to\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$ in probability, when $n_{\min}$ tends to infinity, so that $\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\to\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$.
-We also have, as in (10), where empirical means are replaced by bootstrap means,
-$\left|\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi_{n}^{*}(\boldsymbol{\alpha},\sigma_{\alpha}^{2})\right|$ $\leq\left(\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|+|\sigma_{\alpha}^{2}-\sigma_{\gamma}^{2}|\right)\sum_{k=1}^{K}\pi_{k}\left(B_{n}^{k,*}+A_{n}^{k,*}\right),$ (35)
+We also have, as in (B.3), where empirical means are replaced by bootstrap means,
+$\left|\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi_{n}^{*}(\boldsymbol{\alpha},\sigma_{\alpha}^{2})\right|$ $\leq\left(\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|+|\sigma_{\alpha}^{2}-\sigma_{\gamma}^{2}|\right)\sum_{k=1}^{K}\pi_{k}\left(B_{n}^{k,*}+A_{n}^{k,*}\right),$ (B.13)
 for any $(\boldsymbol{\alpha},\sigma_{\alpha}^{2})\in\Theta$, with $\sum_{k=1}^{K}\pi_{k}\left(B_{n}^{k,*}+A_{n}^{k,*}\right)=O_{p}(1)$.
-As a result, we deduce from Lemma 2.4, inequality (34) and Lemma 2.9 in Newey and McFadden, (1994) that
+As a result, we deduce from Lemma 4.2, inequality (B.12) and Lemma 2.9 in Newey and McFadden, (1994) that
 $\sup_{(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\in\Theta}\left|\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\right|\to 0\quad\mbox{ in probability.}$.
@@
 By definition of $\widehat{\boldsymbol{\theta}}$ and Taylor expansion we have
-$\nabla\varphi_{n}(\widehat{\boldsymbol{\theta}})=\nabla\varphi_{n}(\boldsymbol{\theta}_{0})+\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\left(\widehat{\boldsymbol{\theta}}-\boldsymbol{\theta}_{0}\right)=0$ (36)
+$\nabla\varphi_{n}(\widehat{\boldsymbol{\theta}})=\nabla\varphi_{n}(\boldsymbol{\theta}_{0})+\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\left(\widehat{\boldsymbol{\theta}}-\boldsymbol{\theta}_{0}\right)=0$ (B.14)
 where $\overline{\boldsymbol{\theta}}$ belongs, componentwise, to the segment between $\boldsymbol{\theta}_{0}$ and $\widehat{\boldsymbol{\theta}}$.
 We have a similar expansion for boostrap estimators, as well as
-$\nabla\varphi_{n}^{*}(\boldsymbol{\theta}^{*})=\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})+\nabla_{00}^{*}\varphi_{n}(\overline{\boldsymbol{\theta}}^{*})\left(\boldsymbol{\theta}^{*}-\boldsymbol{\theta}_{0}\right)=0$ (37)
-where $\overline{\boldsymbol{\theta}}^{*}$ belongs, componentwise, to the segment between $\boldsymbol{\theta}_{0}$ and $\boldsymbol{\theta}^{*}$. Combining (36) and (37), we deduce
+$\nabla\varphi_{n}^{*}(\boldsymbol{\theta}^{*})=\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})+\nabla_{00}^{*}\varphi_{n}(\overline{\boldsymbol{\theta}}^{*})\left(\boldsymbol{\theta}^{*}-\boldsymbol{\theta}_{0}\right)=0$ (B.15)
+where $\overline{\boldsymbol{\theta}}^{*}$ belongs, componentwise, to the segment between $\boldsymbol{\theta}_{0}$ and $\boldsymbol{\theta}^{*}$. Combining (B.14) and (B.15), we deduce
 $\boldsymbol{\theta}^{*}-\widehat{\boldsymbol{\theta}}$ $=\left(\nabla_{00}^{*}\varphi_{n}(\overline{\boldsymbol{\theta}}^{*})\right)^{-1}\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})-\left(\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\right)^{-1}\nabla\varphi_{n}(\boldsymbol{\theta}_{0})$
-$=\left(\left(\nabla_{00}^{*}\varphi_{n}(\overline{\boldsymbol{\theta}}^{*})\right)^{-1}-\left(\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\right)^{-1}\right)\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})+\left(\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\right)^{-1}\left(\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})-\nabla\varphi_{n}(\boldsymbol{\theta}_{0})\right)$ (38)
+$=\left(\left(\nabla_{00}^{*}\varphi_{n}(\overline{\boldsymbol{\theta}}^{*})\right)^{-1}-\left(\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\right)^{-1}\right)\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})+\left(\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\right)^{-1}\left(\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})-\nabla\varphi_{n}(\boldsymbol{\theta}_{0})\right)$ (B.16)
 Noticing that $\nabla_{00}^{*}\varphi_{n}(\overline{\boldsymbol{\theta}}^{*})$ and $\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})$ both tend in probability to the same limit $\mathbf{H}(\boldsymbol{\beta},\sigma^{2}_{\epsilon})$ and we have, with similar arguments as those used in the proof of Proposition 4.2, that $\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})$ is $O_{p}(n^{-1/2})$, it can be deduce that
-$\boldsymbol{\theta}^{*}-\widehat{\boldsymbol{\theta}}$ $=\left(\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\right)^{-1}\left(\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})-\nabla\varphi_{n}(\boldsymbol{\theta}_{0})\right)+o_{P}(n^{-1/2}).$ (39)
+$\boldsymbol{\theta}^{*}-\widehat{\boldsymbol{\theta}}$ $=\left(\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\right)^{-1}\left(\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})-\nabla\varphi_{n}(\boldsymbol{\theta}_{0})\right)+o_{P}(n^{-1/2}).$ (B.17)
 Using arguments similar to those employed in the expansion of $\nabla\varphi_{n}$ in the proof of Proposition 4.2, we make appear the difference between bootstrap means and empirical means or a differentiable functional of these quantities:
-$\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})-\nabla\varphi_{n}(\boldsymbol{\theta}_{0})=\footnotesize\begin{pmatrix}2\sum_{k=1}^{K}\pi_{k}\left((\widehat{\mu}_{Y}^{k}-{\mu}_{Y}^{k,*})-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\left(\widehat{\boldsymbol{\mu}}_{X}-\boldsymbol{\mu}_{X}^{k,*}\right)\right)\\ \begin{pmatrix}2\sum_{k=1}^{K}\pi_{k}\left[\left(\widehat{\mu}_{Y}^{k}-\mu_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\widehat{\boldsymbol{\mu}}_{X}^{k}+\left(\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-1\right)\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}\right]\\ -2\sum_{k=1}^{K}\pi_{k}\left[\left(\mu_{Y}^{k,*}-\mu_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\mu}_{X}^{k,*}\right)\boldsymbol{\mu}_{X}^{k,*}+\left(\frac{\sigma_{Y,k}^{*}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k,*}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-1\right)\boldsymbol{\Gamma}_{X}^{k,*}\boldsymbol{\beta}_{-0}\right]\end{pmatrix}\\ \sum_{k=1}^{K}\left(\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-\frac{\sigma_{Y,k}^{*}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k,*}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}\right)\end{pmatrix},$ (45)
+$\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})-\nabla\varphi_{n}(\boldsymbol{\theta}_{0})=\footnotesize\begin{pmatrix}2\sum_{k=1}^{K}\pi_{k}\left((\widehat{\mu}_{Y}^{k}-{\mu}_{Y}^{k,*})-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\left(\widehat{\boldsymbol{\mu}}_{X}-\boldsymbol{\mu}_{X}^{k,*}\right)\right)\\ \begin{pmatrix}2\sum_{k=1}^{K}\pi_{k}\left[\left(\widehat{\mu}_{Y}^{k}-\mu_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\widehat{\boldsymbol{\mu}}_{X}^{k}+\left(\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-1\right)\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}\right]\\ -2\sum_{k=1}^{K}\pi_{k}\left[\left(\mu_{Y}^{k,*}-\mu_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\mu}_{X}^{k,*}\right)\boldsymbol{\mu}_{X}^{k,*}+\left(\frac{\sigma_{Y,k}^{*}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k,*}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-1\right)\boldsymbol{\Gamma}_{X}^{k,*}\boldsymbol{\beta}_{-0}\right]\end{pmatrix}\\ \sum_{k=1}^{K}\left(\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-\frac{\sigma_{Y,k}^{*}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k,*}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}\right)\end{pmatrix},$ (B.18)
 which satisfies the central limit theorem for bootstrap means, or the Delta method for bootstrap estimators (see the Appendix as well as Theorem 23.4 and Theorem 23.5 in van der Vaart, (1998)).
@@
 ∎
+Appendix C Supplementary figures
+Figure C.1: a) Coverage rates, b) amplitudes, and c) powers of the confidence intervals for the estimators of $\beta_{1}$ obtained from $N_{sim}$ simulations for each parameter combination.
+The method labels on the left: "mm (asymp)" stands for the method of moments with asymptotic confidence intervals, "mm (boot)" for the method of moments with bootstrap, "ot (boot)" for the optimal transport method with bootstrap, and "mm (student)" for the naive linear regression on means approach based on Student’s distribution.

### v4

@@
-Estimating the linear relation between variables that are never jointly observed
+Bootstrap inference for linear regression between variables that are never jointly observed: application in in vivo experiments
 Abstract
-In modern experimental science there is a commonly encountered problem of estimating the coefficients of a linear regression in the context where the variables of interest can never be observed simultaneously.
-Assuming that the global experiment can be decomposed into sub-experiments with distinct first moments, we propose two estimators of the linear regression that take this additional information into account.
-We consider an estimator based on moments, and an estimator based on optimal transport theory.
-These estimators are proven to be consistent as well as asymptotically Gaussian under weak hypotheses.
-The asymptotic variance has no explicit expression, except in some particular cases, for which reason a stratified bootstrap approach is developed to build confidence intervals for the estimated parameters, whose consistency is also shown.
-A simulation study, assessing and comparing the finite sample performances of these estimators, demonstrated the advantages of the bootstrap approach in multiple realistic scenarios.
-An application to in vivo experiments, conducted in the context of studying radio-induced adverse effects on mice, revealed important relationships between the biomarkers of interest that could not be identified with the considered naive approach.
+In modern experimental science, there is a common problem of estimating the coefficients of a linear regression in a context where the variables of interest cannot be observed simultaneously.
+When there is a categorical variable that is observed on all statistical units, we consider two estimators of linear regression that take this additional information into account: an estimator based on moments and an estimator based on optimal transport theory.
+These estimators are shown to be consistent and asymptotically Gaussian under weak hypotheses.
+The asymptotic variance has no explicit expression, except in some special cases, for which reason a stratified bootstrap approach is developed to construct confidence intervals for the estimated parameters, whose consistency is also shown.
+A simulation study evaluating and comparing the finite sample performance of these estimators demonstrates the advantages of the bootstrap approach in several realistic scenarios.
+An application to in vivo experiments, conducted in the context of studying radio-induced adverse effects in mice, revealed important relationships between the biomarkers of interest that could not be identified with the considered naive approach.
 1 Introduction
-In order to study the effects of a certain treatment on a living organism, in vivo experiments are often conducted.
-In the context of a complex organism response, scientists may be interested in studying multiple variables describing the effect on different scales.
-In particular, such variables of interest often include a macroscopic biomarker, only available through in vivo data, and a microscopic biomarker that can also be observed on a cellular level (i.e.
+In vivo experiments are often used to study the effects of a treatment on a living organism.
+In the context of a complex organism response, scientists may be interested in studying multiple variables that describe the effect at different scales.
+In particular, such variables of interest often include a macroscopic biomarker that is only available through in vivo data, and a microscopic biomarker that can also be observed at the cellular level (i.e.
 in vitro).
-The interest in this case lies in predicting the former with the latter.
-For instance, in the context of studying the adverse effects induced by radiotherapy on healthy tissues, the potential outcomes of interest manifest in the form of the severity of macroscopic lesions, and predictors such as gene expression.
-These measures often require sacrificing the animals that they are taken from.
-As a result, the quantities of interest cannot be observed on the same animals.
-Since the goal is to establish relationships between these variables, a problem of statistical data fusion arises.
-Figure 1.1: Schematic representation of the design of an in vivo experiment studying the effect of irradiated volume.
-An illustrative example of an in vivo experiment where the variables of interest are never observed simultaneously is presented in Figure 1.1.
-In this experiment, mice are irradiated on the lungs with different volume, with a goal of studying the role of irradiated volume in the appearance of radio-induced adverse effects.
-The latter are assessed by measuring septal thickening, a histological marker of lung injuries.
-The other variable that is measured with the purpose of predicting the adverse effect-related variable is the expression of multiple pro-inflammatory genes.
-As shown in Figure 1.1, there are two independent cohorts in the study, one is used to measure gene expression, whereas the other for measuring septal thickness.
-This is a results of both measures being of destructive nature, which does not allow them to be taken on the same animals.
-Comparing distributions of the measurements arising from the two cohorts, such as those presented in Figure 1.2, one may suspect a correlation or even a linear relationship between the variables.
-In order to assert whether such a relationship exists, one has to connect two variables that are never observed simultaneously, which translates into solving a data fusion problem.
-In this example, there are four groups indicating time points (1, 3, 6 and 12 months after irradiation) when the corresponding animals are sacrificed and the measurements are taken.
-Thus, the categorical variable indicating the time point, which is known for every observation, can be used as an additional variable in order to link the predictor and the predicted variables between each other.
-Figure 1.2: Distribution of the data, collected from the irradiated patch under SBRT with 3 mm beam size: the expression of the gene IL6 on the left, and septal thickness on the right.
+The interest in this case is to use the latter to predict the former.
+For example, in the context of studying the adverse effects of radiotherapy on healthy tissue, the potential outcomes of interest are the severity of macroscopic lesions and predictors such as gene expression.
+In preclinical research, these quantities of interest are often observed in different animals from independent cohorts.
+Since the goal is to establish relationships between these variables, the problem of statistical data fusion arises.
+Figure 1: Schematic representation of the design of an in vivo experiment studying the effect of irradiated volume.
+An illustrative example of an in vivo experiment where the variables of interest are not observed simultaneously is shown in Figure 1.
+In this experiment, presented in Bertho et al., (2020), mice are irradiated in the lungs with different volumes to study the role of irradiated volume in the occurrence of radiation-induced adverse effects.
+The latter are assessed by measuring septal thickening, a histological marker of lung injury.
+The other variable measured to predict the adverse effect variable is the expression of several pro-inflammatory genes.
+As shown in Figure 1, there are two independent cohorts in the study, one used to measure gene expression and the other to measure septal thickness.
+Comparing the distributions of measurements from the two cohorts, as shown in Figure 2, may suggest a correlation or even a linear relationship between the variables.
+To establish whether such a relationship exists, it is necessary to link two variables that are not observed on the same statistical units, which is equivalent to solving a data fusion or a file matching problem according to the terminology employed in Little and Rubin, (2002).
+In this example, there are four groups indicating the time points (1, 3, 6 and 12 months after irradiation) when the measurements were taken on the corresponding animals.
+Thus, the categorical variable indicating the time point, which is known for each observation, can be used as an additional variable to link the predictor and the predicted variables.
+Figure 2: Distribution of the data, collected from the irradiated patch under SBRT with 3 mm beam size: the expression of the gene IL6 on the left, and septal thickness on the right.
 The measurements were made 1, 3, 6 and 12 months after irradiation.
-The task of linking variables that are never jointly observed cannot be approached as a typical missing values problem, since most methods for inference on incomplete data require a sufficient overlap, which is completely absent in the case treated here.
-As a result, any approaches applying frameworks such as multiple imputation in the context of data fusion are unsuitable for our application.
-For example, Carrig et al., (2015) use multiple imputation to integrate disparate datasets allowing for the absence of the overlap, but requiring a calibration dataset, where all the variables of interest must be jointly observed.
-Other approaches to data fusion available in literature include factor analysis (Cudeck,, 2000), statistical matching (Mitsuhiro and Hoshino,, 2020), Bayesian network inference (Triantafillou et al.,, 2010; Tsamardinos et al.,, 2012) and Gaussian Markov combinations (Massa and Riccomagno,, 2017).
-These methods are designed for linking variables that are not observed simultaneously through covariates, present for both variables of interest.
-This corresponds to the properties of the in vivo data described above.
-However, the covariates in these approaches are random variables, typically continuous, and often assumed to be Gaussian, which is the case in Cudeck, (2000) and Massa and Riccomagno, (2017).
-The grouping variable available through in vivo experiments cannot be represented in a continuous form since categories such as control and sham make it impossible to assume continuity and normality.
-The Bayesian network approaches introduced by Triantafillou et al., (2010) and Tsamardinos et al., (2012), aimed at inferring binary causal relationships between variables, are more suitable for large datasets with a high number of covariates.
-Finally, currently available research in statistical matching addresses such aspects as not-at-random missingness (Mitsuhiro and Hoshino,, 2021) and high dimensionality (Mitsuhiro and Hoshino,, 2020).
-This approach is based on the idea of comparing distances between the covariates from the datasets of interest, which cannot be done by taking the group variable as the covariate.
-It can be noted that the goal of the aforementioned examples in statistical matching is to group individuals before imputation, which is not necessary in our case since the groups are already known.
-In this work we propose an approach to estimate the relationship between the variables that are not simultaneously observed under the experimental setting described above, based on a conditional model assuming linear relationship within every component, defined by the grouping variable that is assumed to be present.
-An estimator derived with the method of moments as well as optimal transport solution using Wasserstein distance are considered for the real data problem in question.
-Both approaches do not require any overlap between the two cohorts of the experiment, and are based on weak assumptions ensuring model identifiability and moments existence.
-To construct confidence intervals, we propose both asymptotic and bootstrap-based results, the former being often unfeasible in practice, with the latter demonstrating practical advantages in numerous realistic settings.
+The task of linking variables that are not observed together cannot be approached as a typical missing value problem, since most methods of inference on incomplete data require sufficient overlap, which is completely absent in the case we are dealing with.
+As a result, all approaches that use frameworks such as multiple imputation in the context of data fusion are inappropriate for our application.
+For example, Carrig et al., (2015) use multiple imputation to integrate different datasets, which allows for the absence of overlap, but requires a calibration dataset in which all variables of interest must be jointly observed.
+Other approaches to data fusion available in the literature include factor analysis (Cudeck,, 2000), statistical matching (Mitsuhiro and Hoshino,, 2020), Bayesian network inference (Triantafillou et al.,, 2010; Tsamardinos et al.,, 2012) and Gaussian Markov combinations (Massa and Riccomagno,, 2017).
+These methods are designed to link variables that are not simultaneously observed through covariates that exist for both variables of interest.
+This corresponds to the characteristics of in vivo data described above.
+However, the covariates in these approaches are continuous random variables, often assumed to be Gaussian, as is the case in Cudeck, (2000) and Massa and Riccomagno, (2017).
+The grouping variable available from in vivo experiments cannot be represented in continuous form, since categories such as control and sham make it impossible to assume continuity and normality.
+The Bayesian network approaches introduced by Triantafillou et al., (2010) and Tsamardinos et al., (2012), which aim to infer binary causal relationships between variables, are more suitable for large datasets with a high number of covariates.
+Current research in statistical matching addresses aspects such as not-at-random missingness (Mitsuhiro and Hoshino,, 2021) and high dimensionality (Mitsuhiro and Hoshino,, 2020).
+This approach is based on the idea of comparing distances between covariates from the datasets of interest, which cannot be done by taking the group variable as a covariate.
+It can be noted that the goal of the aforementioned examples in statistical matching is to group individuals prior to imputation, which is not necessary in our case since the groups are already known.
+Finally, ecological regression-based approaches are also employed, where the correlations of interest between two covariates relate to their means or percentages within groups.
+The primary difference with our study design is that we have access to the individual marginal observations of each variable.
+This allows us to conduct statistical inferences that ecological regression does not permit (Gelman et al.,, 2001).
+In this work, we assume that there is a linear relationship between the continuous variables that are not simultaneously observed, and that the linear regression coefficients are the same for all sub-populations defined by the categorical variable.
+A similar context is treated in Evans et al., (2021), where a more general model is considered.
+The authors propose an approach that requires correctly specifying various relationships between variables (e.g., the distribution of the predictor variable conditional on what corresponds to the grouping variable in our case) in order to perform successful inference in the general case.
+Their approach is illustrated using survey data, which is the case when the latter can be expected to be successful due to large samples and/or prior knowledge.
+However, this is not the case for the experimental data considered in this work, which are characterized by small sample sizes and lack of prior knowledge about the underlying distributions.
+To address the specificities of the considered context, our approach is based on the assumption that there is a linear relationship conditional on the group, and that this relationship is the same for all groups, without making any assumptions about the distributions of the variables or requiring their specification.
+We propose two estimators derived with the method of moments as well as an optimal transport solution using Wasserstein distance.
+Both approaches do not require any overlap between the two cohorts of the experiment and are based on weak assumptions ensuring model identifiability and the existence of moments.
+These estimators are shown to be consistent and asymptotically Gaussian under weak hypotheses.
+The asymptotic variance has no explicit expression, except in some special simple cases.
+For that reason a consistent stratified bootstrap approach is developed to construct confidence intervals for the estimated parameters.
+Not previously considered in the missing data literature, the bootstrap-based approach can be seen as the most important contribution of this paper, since the asymptotic solution is often infeasible in practice, whereas the bootstrap shows practical advantages in many realistic settings, especially in the case of a small sample size.
 2 Identification approaches
-We consider a real random variable $Y$ and a vector of $d$ real valued random regressors $\mathbf{X}=(X_{1},\ldots,X_{d})$ and suppose that the following linear regression hold
-$Y$ $=\beta_{0}+\sum_{j=1}^{d}\beta_{j}X_{j}+\epsilon.$ (2.1)
+We consider a real random variable $Y$ and a vector of $d$ real valued random regressors $\mathbf{X}=(X_{1},\ldots,X_{d})$, and suppose that the following linear regression holds:
+$Y$ $=\beta_{0}+\sum_{j=1}^{d}\beta_{j}X_{j}+\epsilon.$ (1)
 […]
-In the absence of additional information and without any strong additional hypothesis, the parameters $(\beta_{0},\beta_{1},\ldots,\beta_{d})$ and the variance of the noise $\sigma^{2}_{\epsilon}$ cannot be identified.
-Indeed, if for example $X_{1}$ is centered with symmetric distribution, the coefficient $\beta_{1}$ can only be determined up to sign change since $\beta_{1}X_{1}$ and $\beta_{1}(-X_{1})$ have the same distribution.
-To deal with this identification issue, we consider that we can perform different experiments in which the mean value of $X$ is allowed to vary.
-For that, we suppose that there are $K$ groups (corresponding to $K$ different experiments), defined by a discrete variable $G$ taking values in $\{1,\cdots,K\}$ observed simultaneously with $Y$ and with $X$.
-This means that we have now access to $(\mathbf{X},G)$ and $(Y,G)$ but not to $(\mathbf{X},Y,G)$.
-We also suppose that $\epsilon$ is independent of $G$.
+In the absence of additional information and without a strong additional hypothesis, the parameters $(\beta_{0},\beta_{1},\ldots,\beta_{d})$ and the variance of the noise $\sigma^{2}_{\epsilon}$ cannot be identified.
+For example, if $X_{1}$ is centered and has a symmetric distribution, the coefficient $\beta_{1}$ can only be determined up to the sign change, since $\beta_{1}X_{1}$ and $\beta_{1}(-X_{1})$ have the same distribution.
+To deal with this identification problem, we consider that we can perform different experiments in which the mean of $X$ is allowed to vary.
+To do this, we assume that there are $K$ groups (i.e.
+$K$ different experiments), defined by a categorical variable $G$ taking values in $\{1,\ldots,K\}$ observed simultaneously with $Y$ and $X$.
+This means that we now have access to $(\mathbf{X},G)$ and $(Y,G)$, but not to $(\mathbf{X},Y,G)$.
+We also assume that $\epsilon$ is independent of $G$.
 Given $G=k$, for $k=1,\ldots,K$, we denote by $\mu_{Y}^{k}={\mathbb{E}}(Y|G=k)$ and $\mu_{X_{j}}^{k}={\mathbb{E}}(X_{j}|G=k)$, $j=1,\ldots,d$, the expected values within each group.
-We now present two different approaches developed to identify the vector $\boldsymbol{\beta}=(\beta_{0},\ldots,\beta_{d})$ of unknown regression coefficients and the noise variance $\sigma_{\epsilon}^{2}$, taking account of the additional information given by the discrete variable $G$.
+We present two different approaches to identify the vector $\boldsymbol{\beta}=(\beta_{0},\ldots,\beta_{d})$ of unknown regression coefficients and the noise variance $\sigma_{\epsilon}^{2}$, taking into account the additional information provided by the discrete variable $G$.
 2.1 Moment approach
-A first and simple approach is based on first moments identification. Taking the conditional expectation, given $G=k$, in (2.1), we have for $k=1,\ldots,K$,
-$\mu_{Y}^{k}$ $=\beta_{0}+\sum_{j=1}^{d}\beta_{j}\mu_{X_{j}}^{k},$ (2.2)
-since the residual term $\epsilon$ is supposed to satisfy $\mathbb{E}(\epsilon|G=k)=0$ for $k=1,\ldots,K$.
-We denote by $\boldsymbol{\mu}_{1,X}$ the $K\times(d+1)$ design matrix, whose $k$th row is equal to $(1,\boldsymbol{\mu}_{X}^{k\top})$ with $\boldsymbol{\mu}_{X}^{k}=(\mu_{X_{1}}^{k},\cdots,\mu_{X_{d}}^{k})^{\top}$, by $\boldsymbol{\mu}_{Y}$ the $K$ dimensional vector with elements $(\mu_{Y}^{1},\ldots,\mu_{Y}^{K})$.
-The $K$ linear equations in (2.2) can we written in a matrix way, $\boldsymbol{\mu}_{Y}=\boldsymbol{\mu}_{1,X}\boldsymbol{\beta}$.
-We introduce the following assumption, guaranteeing the identifiability of the model parameters:
+The first simple approach is based on the first moments identification. Taking the conditional expectation, given $G=k$, in (1), we have for $k=1,\ldots,K$,
+$\mu_{Y}^{k}$ $=\beta_{0}+\sum_{j=1}^{d}\beta_{j}\mu_{X_{j}}^{k},$ (2)
+since the residual term $\epsilon$ is assumed to satisfy $\mathbb{E}(\epsilon|G=k)=0$ for $k=1,\ldots,K$.
+We denote by $\boldsymbol{\mu}_{1,X}$ the $K\times(d+1)$ design matrix, with the $k$th row equal to $(1,\boldsymbol{\mu}_{X}^{k\top})$ with $\boldsymbol{\mu}_{X}^{k}=(\mu_{X_{1}}^{k},\cdots,\mu_{X_{d}}^{k})^{\top}$, and by $\boldsymbol{\mu}_{Y}$ the $K$ dimensional vector with elements $(\mu_{Y}^{1},\ldots,\mu_{Y}^{K})$.
+The $K$ linear equations in (2) can be written in a matrix form: $\boldsymbol{\mu}_{Y}=\boldsymbol{\mu}_{1,X}\boldsymbol{\beta}$.
+The following assumption guarantees the identifiability of the model parameters:
 $\mathbf{H}_{1}$ $\quad\mbox{rank}(\boldsymbol{\mu}_{1,X})=d+1,$
 meaning that there are at least $K\geq d+1$ groups and that the $d+1$ column vectors of $\boldsymbol{\mu}_{1,X}$ span a vector space of dimension $d+1$ in $\mathbb{R}^{K}$.
-Lemma 2.1.
-If model (2.1) holds and assumption $\mathbf{H}_{1}$ is fulfilled, $\boldsymbol{\beta}$ is uniquely identified in terms of the conditional first order moments of $\mathbf{X}$ and $Y$ given $G$,
+Lemma 2.1
+If the model (1) holds and the assumption $\mathbf{H}_{1}$ is fulfilled, $\boldsymbol{\beta}$ is uniquely identified in terms of the conditional first order moments of $\mathbf{X}$ and $Y$ given $G$,
 $\boldsymbol{\beta}$ $=\left(\boldsymbol{\mu}_{1,X}^{\top}\boldsymbol{\mu}_{1,X}\right)^{-1}\boldsymbol{\mu}_{1,X}^{\top}\boldsymbol{\mu}_{Y}.$
@@
 […]
-If we assume that, given $G=k$, $\mathbf{X}$ is a Gaussian random vector and $\epsilon$ is Gaussian, we have, if the linear model (2.1) holds, that $Y$ is also Gaussian given $G=k$, with expectation $\mu_{Y}^{k}=\beta_{0}+\sum_{j=1}^{d}\beta_{j}\mu_{X_{j}}^{k}$ and variance $\sigma_{\epsilon}^{2}+\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}$, where $\boldsymbol{\Gamma}_{X}^{k}$ is the variance matrix of $\mathbf{X}$ given $G=k$.
+If the linear model (1) holds, and if we assume that, given $G=k$, $\mathbf{X}$ is a Gaussian random vector and $\epsilon$ is Gaussian, we have that $Y$ is also Gaussian given $G=k$, with expectation $\mu_{Y}^{k}=\beta_{0}+\sum_{j=1}^{d}\beta_{j}\mu_{X_{j}}^{k}$ and variance $\sigma_{\epsilon}^{2}+\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}$, where $\boldsymbol{\Gamma}_{X}^{k}$ is the variance matrix of $\mathbf{X}$ given $G=k$.
 Thus, the Wasserstein distance between $D_{\gamma}$, the distribution of $\gamma_{0}+\boldsymbol{\gamma}_{-0}^{\top}\mathbf{X}+\epsilon$, and $D_{Y}$, the distribution of $Y$, is equal to
@@
 We introduce the following loss criterion
-$\varphi(\boldsymbol{\gamma},\sigma^{2})$ $=\sum_{k=1}^{K}\pi_{k}\left[(\mu_{Y}^{k}-\gamma_{0}-\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\mu}_{X}^{k})^{2}+\left(\sigma_{Y,k}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma^{2}}\right)^{2}\right].$ (2.3)
-which evaluates the weighted Wasserstein distance between $Y$ and a linear combination of the $\mathbf{X}$ variables, contaminated by a Gaussian noise, with variance $\sigma^{2}$.
-We state, without proof, the following Lemma ensuring that the parameters $\boldsymbol{\beta}$ and $\sigma_{\epsilon}^{2}$ can be identified under general conditions.
-Lemma 2.2.
-If model (2.1) holds and if assumption $\mathbf{H}_{1}$ is fulfilled, $\varphi(\boldsymbol{\gamma},\sigma^{2})$ has its unique minimum at $\boldsymbol{\gamma}=\boldsymbol{\beta}$ and $\sigma^{2}=\sigma_{\epsilon}^{2}$.
+$\varphi(\boldsymbol{\gamma},\sigma^{2})$ $=\sum_{k=1}^{K}\pi_{k}\left[(\mu_{Y}^{k}-\gamma_{0}-\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\mu}_{X}^{k})^{2}+\left(\sigma_{Y,k}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma^{2}}\right)^{2}\right].$ (3)
+which evaluates the weighted Wasserstein distance between $Y$ and a linear combination of the $\mathbf{X}$ variables, contaminated by Gaussian noise, with variance $\sigma^{2}$.
+We state, without proof, the following lemma which ensures that the parameters $\boldsymbol{\beta}$ and $\sigma_{\epsilon}^{2}$ can be identified under general conditions:
+Lemma 2.2
+If the model (1) holds and the assumption $\mathbf{H}_{1}$ is fulfilled, $\varphi(\boldsymbol{\gamma},\sigma^{2})$ has its unique minimum at $\boldsymbol{\gamma}=\boldsymbol{\beta}$ and $\sigma^{2}=\sigma_{\epsilon}^{2}$.
 3 Sampled data and estimators
-We suppose that experiments are made for $K\geq 2$ different groups and that for each group $k$, for $k=1,\ldots,K$, we have two independent samples $(Y^{k}_{1},\ldots,Y^{k}_{n_{y}^{k}})$ and $(X^{k}_{j,1},\ldots,X^{k}_{j,n_{x}^{k}})_{j=1,\ldots,d}$, with sizes $n_{y}^{k}$ and $n_{x}^{k}$.
+We assume that the experiments are performed for $K\geq 2$ different groups, and that for each group $k$, for $k=1,\ldots,K$ we have two independent samples $(Y^{k}_{1},\ldots,Y^{k}_{n_{y}^{k}})$ and $(X^{k}_{j,1},\ldots,X^{k}_{j,n_{x}^{k}})_{j=1,\ldots,d}$, with sizes $n_{y}^{k}$ and $n_{x}^{k}$.
 For each unit $i=1,\ldots,n_{x}^{k}$ from group $k$, the vector of covariates is denoted by $\mathbf{X}_{i}^{k}=(X_{1,i},\ldots,X_{d,i})$.
 We also define $N_{x}=\sum_{k=1}^{K}n_{x}^{k}$ and $N_{y}=\sum_{k=1}^{K}n_{y}^{k}$, the total number of observations of the response $Y$ and the covariates $X_{1},\ldots,X_{d}$.
-As shown in Lemma 2.1 and Lemma 2.2, identification of parameter $\boldsymbol{\beta}$ relies on the knowledge of the first two conditional moments, given $G$, of $Y$ and $\mathbf{X}$.
-For $k=1,\ldots,K$, we denote by $\widehat{\mu}_{Y}^{k}=\frac{1}{n_{y}^{k}}\sum_{i=1}^{n_{y}^{k}}Y^{k}_{i}$ and $\widehat{\mu}_{X_{j}}^{k}=\frac{1}{n_{x}^{k}}\sum_{i=1}^{n_{x}^{k}}X^{k}_{j,i}$ the empirical means within each group $k$ and by $\widehat{\sigma}^{2}_{Y,k}=\frac{1}{n_{y}^{k}}\sum_{i=1}^{n_{y}^{k}}\left(Y_{i}^{k}\right)^{2}-\left(\widehat{\mu}_{Y}^{k}\right)^{2}$, and $\widehat{\boldsymbol{\Gamma}}^{k}_{X}=\frac{1}{n_{x}^{k}}\sum_{i=1}^{n_{x}^{k}}\mathbf{X}_{i}^{k}(\mathbf{X}_{i}^{k})^{\top}-\widehat{\boldsymbol{\mu}}_{X}^{k}(\widehat{\boldsymbol{\mu}}_{X}^{k})^{\top}$, the empirical variances, with $\widehat{\boldsymbol{\mu}}_{X}^{k}=(\widehat{\mu}_{X_{1}}^{k},\ldots,\widehat{\mu}_{X_{d}}^{k})$.
-We also define the overall empirical mean and variance $\widehat{\mu}_{Y}=\frac{1}{N_{y}}\sum_{k=1}^{K}{n_{y}^{k}}\widehat{\mu}_{Y}^{k}$, $\widehat{\boldsymbol{\mu}}_{X}=\frac{1}{N_{x}}\sum_{k=1}^{K}{n_{x}^{k}}\widehat{\boldsymbol{\mu}}_{X}^{k}$, $\widehat{\sigma}^{2}_{Y}=\frac{1}{N_{y}}\sum_{k=1}^{K}\sum_{i=1}^{n_{y}^{k}}\left(Y_{i}^{k}\right)^{2}-\left(\widehat{\mu}_{Y}\right)^{2}$, $\widehat{\boldsymbol{\Gamma}}_{X}=\frac{1}{N_{x}}\sum_{k=1}^{K}\sum_{i=1}^{n_{x}^{k}}\mathbf{X}_{i}^{k}(\mathbf{X}_{i}^{k})^{\top}-\widehat{\boldsymbol{\mu}}_{X}\widehat{\boldsymbol{\mu}}_{X}^{\top}$.
+As shown in Lemma 2.1 and Lemma 2.2, the identification of the parameter $\boldsymbol{\beta}$ depends on the knowledge of the first two conditional moments, given $G$, of $Y$ and $\mathbf{X}$.
+For $k=1,\ldots,K$, we denote by $\widehat{\mu}_{Y}^{k}=\frac{1}{n_{y}^{k}}\sum_{i=1}^{n_{y}^{k}}Y^{k}_{i}$ and $\widehat{\mu}_{X_{j}}^{k}=\frac{1}{n_{x}^{k}}\sum_{i=1}^{n_{x}^{k}}X^{k}_{j,i}$ the empirical mean within each group $k$ and by $\widehat{\sigma}^{2}_{Y,k}=\frac{1}{n_{y}^{k}}\sum_{i=1}^{n_{y}^{k}}\left(Y_{i}^{k}\right)^{2}-\left(\widehat{\mu}_{Y}^{k}\right)^{2}$, and $\widehat{\boldsymbol{\Gamma}}^{k}_{X}=\frac{1}{n_{x}^{k}}\sum_{i=1}^{n_{x}^{k}}\mathbf{X}_{i}^{k}(\mathbf{X}_{i}^{k})^{\top}-\widehat{\boldsymbol{\mu}}_{X}^{k}(\widehat{\boldsymbol{\mu}}_{X}^{k})^{\top}$, the empirical variances, with $\widehat{\boldsymbol{\mu}}_{X}^{k}=(\widehat{\mu}_{X_{1}}^{k},\ldots,\widehat{\mu}_{X_{d}}^{k})$.
+We also define the total empirical mean and variance $\widehat{\mu}_{Y}=\frac{1}{N_{y}}\sum_{k=1}^{K}{n_{y}^{k}}\widehat{\mu}_{Y}^{k}$, $\widehat{\boldsymbol{\mu}}_{X}=\frac{1}{N_{x}}\sum_{k=1}^{K}{n_{x}^{k}}\widehat{\boldsymbol{\mu}}_{X}^{k}$, $\widehat{\sigma}^{2}_{Y}=\frac{1}{N_{y}}\sum_{k=1}^{K}\sum_{i=1}^{n_{y}^{k}}\left(Y_{i}^{k}\right)^{2}-\left(\widehat{\mu}_{Y}\right)^{2}$, $\widehat{\boldsymbol{\Gamma}}_{X}=\frac{1}{N_{x}}\sum_{k=1}^{K}\sum_{i=1}^{n_{x}^{k}}\mathbf{X}_{i}^{k}(\mathbf{X}_{i}^{k})^{\top}-\widehat{\boldsymbol{\mu}}_{X}\widehat{\boldsymbol{\mu}}_{X}^{\top}$.
 We denote by $\widehat{\boldsymbol{\mu}}_{1,X}$ the $K\times(d+1)$ matrix, with the first column consisting of ones, and the rest equal to $\widehat{\boldsymbol{\mu}}_{X}$.
 3.1 Moment estimators
-If $\widehat{\boldsymbol{\mu}}_{1,X}$ is full rank, moment estimators of $\boldsymbol{\beta}=(\beta_{0},\beta_{1},\ldots,\beta_{d})$ can be built by considering the empirical counterpart of identification equations given in Lemma 2.1,
-$\widehat{\boldsymbol{\beta}}^{M}$ $=\left(\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\widehat{\boldsymbol{\mu}}_{1,X}\right)^{-1}\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\widehat{\boldsymbol{\mu}}_{Y}$ (3.1)
+If $\widehat{\boldsymbol{\mu}}_{1,X}$ is full rank, moment estimators of $\boldsymbol{\beta}=(\beta_{0},\beta_{1},\ldots,\beta_{d})$ can be built by considering the empirical counterpart of the identification equations given in Lemma 2.1,
+$\widehat{\boldsymbol{\beta}}^{M}$ $=\left(\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\widehat{\boldsymbol{\mu}}_{1,X}\right)^{-1}\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\widehat{\boldsymbol{\mu}}_{Y}$ (4)
 where $\widehat{\boldsymbol{\mu}}_{Y}=(\widehat{\mu}_{Y}^{1},\ldots,\widehat{\mu}_{Y}^{K})$.
-We consider in the following a slightly more general moment estimator of $\boldsymbol{\beta}$, by introducing a weight $w_{k}$ given to each group $k$ of observations, with $w_{k}>0$ and $\sum_{k=1}^{K}w_{k}=1$.
+In the following, we consider a slightly more general moment estimator of $\boldsymbol{\beta}$, by introducing a weight $w_{k}$ given to each group $k$ of observations, with $w_{k}>0$ and $\sum_{k=1}^{K}w_{k}=1$.
 The weighted moment estimator $\boldsymbol{\beta}$ is defined as the minimizer of
@@
 which is unique if $\widehat{\boldsymbol{\mu}}_{1,X}$ is full rank and defined by
-$\widehat{\boldsymbol{\beta}}^{M}$ $=\left(\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\mathbf{w}\widehat{\boldsymbol{\mu}}_{1,X}\right)^{-1}\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\mathbf{w}\widehat{\boldsymbol{\mu}}_{Y}$ (3.2)
-with $\mathbf{w}=(w_{1},\ldots,w_{K})$. The first moment estimator considered in (3.1) corresponds to equal weights $w_{k}=K^{-1}$.
+$\widehat{\boldsymbol{\beta}}^{M}$ $=\left(\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\mathbf{w}\widehat{\boldsymbol{\mu}}_{1,X}\right)^{-1}\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\mathbf{w}\widehat{\boldsymbol{\mu}}_{Y}$ (5)
+where $\mathbf{w}$ is a diagonal matrix with diagonal elements $w_{1},\ldots,w_{K}$. The first moment estimator considered in (4) corresponds to the case with equal weights $w_{k}=K^{-1}$.
 We can then define the following estimator of the noise variance,
-$\widehat{\sigma}_{\epsilon}^{2,M}$ $=\widehat{\sigma}_{Y}^{2}-(\widehat{\boldsymbol{\beta}}_{-0}^{M})^{\top}\widehat{\boldsymbol{\Gamma}}_{X}\widehat{\boldsymbol{\beta}}_{-0}^{M}.$ (3.3)
+$\widehat{\sigma}_{\epsilon}^{2,M}$ $=\widehat{\sigma}_{Y}^{2}-(\widehat{\boldsymbol{\beta}}_{-0}^{M})^{\top}\widehat{\boldsymbol{\Gamma}}_{X}\widehat{\boldsymbol{\beta}}_{-0}^{M}.$ (6)
 3.2 Optimal transport estimators
-Estimators of $\boldsymbol{\beta}$ and $\sigma^{2}_{\epsilon}$ based on an optimal transport criterion are derived by minimizing the empirical version $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2})$ of functional $\varphi(\boldsymbol{\gamma},\sigma^{2})$ defined by
-$\varphi_{n}(\boldsymbol{\gamma},\sigma^{2})$ $=\sum_{k=1}^{K}\pi_{k}\left[(\widehat{\mu}_{Y}^{k}-\gamma_{0}-\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k})^{2}+\left(\widehat{\sigma}_{Y,k}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma^{2}}\right)^{2}\right].$ (3.4)
+Estimators of $\boldsymbol{\beta}$ and $\sigma^{2}_{\epsilon}$ based on an optimal transport criterion are derived by minimizing the empirical version $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2})$ of a functional $\varphi(\boldsymbol{\gamma},\sigma^{2})$ defined by
+$\varphi_{n}(\boldsymbol{\gamma},\sigma^{2})$ $=\sum_{k=1}^{K}\pi_{k}\left[(\widehat{\mu}_{Y}^{k}-\gamma_{0}-\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k})^{2}+\left(\widehat{\sigma}_{Y,k}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma^{2}}\right)^{2}\right].$ (7)
 Note that in absence of a priori information on the probability $\pi_{k}$ of observing group $k$, we can set $\pi_{k}=K^{-1}$.
-We denote by $(\widehat{\boldsymbol{\beta}}^{W},\widehat{\sigma}^{2,W})$ minimizers of $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2})$ which are obtained with iterative optimization algorithms based on gradient descent.
+We denote by $(\widehat{\boldsymbol{\beta}}^{W},\widehat{\sigma}^{2,W})$ the minimizers of $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2})$, which are obtained with iterative optimization algorithms based on gradient descent.
 The algorithm can be initialized with $(\widehat{\boldsymbol{\beta}}^{M},\widehat{\sigma}^{2,M})$
 4 Consistency and asymptotic distribution
-To study the asymptotic behavior of the estimators of $\boldsymbol{\beta}$ defined in previous section, we suppose that the number $K$ of groups is kept fixed and that for all groups and all variables $\mathbf{X}$ and $Y$, the number of observations tends to infinity.
-This means that $n_{\min}=\min(n_{y}^{1},\cdots,n_{y}^{K},n_{x}^{1},\cdots,n_{x}^{K})$ the smallest sample size among all experiments is also supposed to tend to infinity.
-Under general moment conditions, without any particular assumption on the probability law of $Y$ and $\mathbf{X}$ we get in the following two Lemmas that the moment estimators and the Wasserstein distance based estimators are consistent.
-The proofs are postponed to the Appendix.
-Lemma 4.1.
-If ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{2})<+\infty$, and assumption $\mathbf{H}_{1}$ is fulfilled, the sequence of estimators $(\widehat{\boldsymbol{\beta}}^{M},\widehat{\sigma}_{\epsilon}^{2,M})$ defined by (3.2) and (3.3) converges in probability to $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$ when $n_{\min}$ tends to infinity.
+To study the asymptotic behavior of the estimators of $\boldsymbol{\beta}$ defined in the previous section, we assume that the number $K$ of groups is kept fixed, and that for all groups and all variables $\mathbf{X}$ and $Y$, the number of observations tends to infinity.
+This means that $n_{\min}=\min(n_{y}^{1},\cdots,n_{y}^{K},n_{x}^{1},\cdots,n_{x}^{K})$, the smallest sample size among all experiments, should also tend to infinity.
+Lemma 4.1
+If ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{2})<+\infty$, and the assumption $\mathbf{H}_{1}$ is fulfilled, the sequence of estimators $(\widehat{\boldsymbol{\beta}}^{M},\widehat{\sigma}_{\epsilon}^{2,M})$ defined by (5) and (6) converges in probability to $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$ when $n_{\min}$ tends to infinity.
 For the Wasserstein minimum distance estimator, since there is no explicit expression for the estimators, a compactness assumption is also made to obtain the consistency.
-Lemma 4.2.
-If ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{2})<+\infty$, $(\boldsymbol{\beta},\sigma^{2}_{\epsilon})\in\Theta$ and $\Theta$ is a compact set that does not contain 0, suppose that model (2.1) holds and hypothesis $\mathbf{H}_{1}$ is fulfilled, then the sequence of estimators $(\widehat{\boldsymbol{\beta}}^{W},\widehat{\sigma}^{2,W}_{\epsilon})$ that minimize (3.4) converges in probability to $(\beta,\sigma^{2}_{\epsilon})$ when $n_{\min}$ tends to infinity.
-As far as the asymptotic distribution of the estimators is concerned, and for sake of simplicity and lighter notations, we suppose from now on that the number of experiments is the same for all groups and all variables, that is to say $n=n_{y}^{1}=\cdots n_{y}^{K}=n_{x}^{1}\cdots=n_{x}^{K}$.
-Proposition 4.1.
-Assume that the assumptions of Lemma 4.1 are fulfilled. Then as $n$ tends to infinity,
+Lemma 4.2
+If ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{2})<+\infty$, $(\boldsymbol{\beta},\sigma^{2}_{\epsilon})\in\Theta$ and $\Theta$ is a compact set that does not contain 0, if the model (1) holds and the hypothesis $\mathbf{H}_{1}$ is fulfilled, then the sequence of estimators $(\widehat{\boldsymbol{\beta}}^{W},\widehat{\sigma}^{2,W}_{\epsilon})$ that minimize (7) converges in probability to $(\beta,\sigma^{2}_{\epsilon})$ when $n_{\min}$ tends to infinity.
+As far as the asymptotic distribution of the estimators is concerned, and for the sake of simplicity and simpler notation, from now on we will assume that the number of experiments is the same for all groups and all variables, i.e.
+$n=n_{y}^{1}=\ldots=n_{y}^{K}=n_{x}^{1}=\ldots=n_{x}^{K}$.
+Proposition 4.1
+If the assumptions of Lemma 4.1 are fulfilled, as $n$ tends to infinity,
 $\sqrt{n}\left(\widehat{\boldsymbol{\beta}}^{M}-\boldsymbol{\beta}\right)\rightsquigarrow\mathcal{N}\left(0,\boldsymbol{\Gamma}_{\beta_{M}}\right)$
 where the expression of the asymptotic covariance matrix $\boldsymbol{\Gamma}_{\beta_{M}}$ is given in the proof.
-This result is based on the CLT for empirical means and the application of the delta method, which involves computing the Jacobian of the inverse of matrices, making obtaining the explicit expression for $\boldsymbol{\Gamma}_{\beta_{M}}$ a difficult task when $d>1$.
-Remark 4.1.
-The weak convergence toward a Gaussian distribution presented in Proposition 4.1 would remain true, at the expense of heavier notations and a different asymptotic covariance matrix $\boldsymbol{\Gamma}_{\beta_{M}}$, provided that there exist two constants, $0<c\leq C$ such that
-$0<c\leq\frac{\max(n_{y}^{1},\cdots,n_{y}^{K},n_{x}^{1},\cdots,n_{x}^{K})}{\min(n_{y}^{1},\cdots,n_{y}^{K},n_{x}^{1},\cdots,n_{x}^{K})}\leq C<+\infty,$ (4.1)
+This result is based on the central limit theorem for empirical means and the application of the delta method, which involves computing the Jacobian of the inverse of matrices, making it difficult to obtain the explicit expression for $\boldsymbol{\Gamma}_{\beta_{M}}$ when $d>1$.
+Remark 4.1
+The weak convergence toward a Gaussian distribution presented in Proposition 4.1 remains true, at the expense of heavier notation and a different asymptotic covariance matrix $\boldsymbol{\Gamma}_{\beta_{M}}$, provided that there exist two constants, $0<c\leq C$ such that
+$0<c\leq\frac{\max(n_{y}^{1},\cdots,n_{y}^{K},n_{x}^{1},\cdots,n_{x}^{K})}{\min(n_{y}^{1},\cdots,n_{y}^{K},n_{x}^{1},\cdots,n_{x}^{K})}\leq C<+\infty,$ (8)
 and $\min(n_{y}^{1},\cdots,n_{y}^{K},n_{x}^{1},\cdots,n_{x}^{K})\to\infty$.
-The asymptotic normality of $\widehat{\boldsymbol{\beta}}^{W}$ relies on classical results for M-estimators recalled in the Appendix (see Theorem A.4).
-Note that since we estimate simultaneously $\beta$ and $\sigma_{\epsilon}^{2}$, an additional condition on the existence of the moments of order four for the covariates is required.
-Proposition 4.2.
-If model (2.1) holds and hypothesis $\mathbf{H}_{1}$ is fulfilled, ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{4})<+\infty$, $(\boldsymbol{\beta},\sigma^{2}_{\epsilon})\in\Theta$ and $\Theta$ is a compact set that does not contain $(0,0)$, then, as $n$ tends to infinity,
+The asymptotic normality of $\widehat{\boldsymbol{\beta}}^{W}$ relies on classical results for M-estimators recalled in Supplementary Material (see Theorem A.4).
+Note that since we estimate $\beta$ and $\sigma_{\epsilon}^{2}$ simultaneously, an additional condition on the existence of the moments of order four is required for the covariates.
+Proposition 4.2
+If the model (1) holds, the hypothesis $\mathbf{H}_{1}$ is fulfilled, ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{4})<+\infty$, $(\boldsymbol{\beta},\sigma^{2}_{\epsilon})\in\Theta$ and $\Theta$ is a compact set that does not contain $(0,0)$, then, as $n$ tends to infinity,
 $\sqrt{n}\left(\begin{pmatrix}\widehat{\boldsymbol{\beta}}^{W}\\ \widehat{\sigma}^{2,W}_{\epsilon}\end{pmatrix}-\begin{pmatrix}\boldsymbol{\beta}\\ \sigma^{2}_{\epsilon}\end{pmatrix}\right)$ $\rightsquigarrow\mathcal{N}\left(0,\boldsymbol{\Gamma}_{W}\right),$
@@
 5 The particular case of simple linear regression
-To show the difficulty, let us examine the case of simple linear regression, that is to say $d=1$. The following linear model
+To illustrate the difficulty, consider the case of a simple linear regression, i.e. $d=1$. The following linear model
 $Y$ $=\beta_{0}+\beta_{1}X+\epsilon$
-is supposed to hold and if there are two groups $k$ and $j$ such that $\mu_{X}^{k}\neq\mu_{X}^{j}$, the identification assumption $\mathbf{H}_{1}$ is fulfilled and $\beta_{0}$ and $\beta_{1}$ can be uniquely determined.
-The moment estimators of $\beta_{0}$ and $\beta_{1}$ defined in (3.2) have simple expressions,
-$\widehat{\beta}_{0}$ $=\widehat{\mu}_{Y,w}-\widehat{\beta}_{1}\widehat{\mu}_{X,w}$ (5.1)
-$\widehat{\beta}_{1}$ $=\frac{\sum_{k=1}^{K}w_{k}\widehat{\mu}_{X}^{k}\widehat{\mu}_{Y}^{k}-\widehat{\mu}_{X,w}\widehat{\mu}_{Y,w}}{\sum_{k=1}^{K}w_{k}\left(\widehat{\mu}_{X}^{k}\right)^{2}-\widehat{\mu}_{X,w}^{2}}$ (5.2)
+is assumed to hold, and if there are two groups $k$ and $j$ such that $\mu_{X}^{k}\neq\mu_{X}^{j}$, the identification assumption $\mathbf{H}_{1}$ is fulfilled and $\beta_{0}$ and $\beta_{1}$ can be uniquely determined.
+The moment estimators of $\beta_{0}$ and $\beta_{1}$ defined in (5) have simple expressions:
+$\widehat{\beta}_{0}$ $=\widehat{\mu}_{Y,w}-\widehat{\beta}_{1}\widehat{\mu}_{X,w}$ (9)
+$\widehat{\beta}_{1}$ $=\frac{\sum_{k=1}^{K}w_{k}\widehat{\mu}_{X}^{k}\widehat{\mu}_{Y}^{k}-\widehat{\mu}_{X,w}\widehat{\mu}_{Y,w}}{\sum_{k=1}^{K}w_{k}\left(\widehat{\mu}_{X}^{k}\right)^{2}-\widehat{\mu}_{X,w}^{2}}$ (10)
 […]
-We focus on the asymptotic variance of the estimator $\widehat{\beta}_{1}$ of the slope parameter $\beta_{1}$, which is often the parameter of interest. We have
-Lemma 5.1.
-Suppose that the model (2.1) is true, with $d=1$ and $K\geq 2$.
+We focus on the asymptotic variance of the estimator $\widehat{\beta}_{1}$ of the slope parameter $\beta_{1}$, which is often the parameter of interest.
+Lemma 5.1
+Suppose that the model (1) is true, with $d=1$ and $K\geq 2$.
 If there are two groups $k$ and $j$ such that $\mu_{X}^{k}\neq\mu_{X}^{j}$ the vector $(\beta_{0},\beta_{1})$ is identifiable and, as $n$ tends to infinity,
@@
 and $\mu_{X,w}=\sum_{k=1}^{K}w_{k}\mu_{X}^{k}$ and $\mbox{Var}_{w}(X)=\sum_{k=1}^{K}w_{k}\left(\mu_{X}^{k}-\mu_{X,w}\right)^{2}$.
-Lemma 5.1 shows us that even in a very simple framework (only one covariate), the asymptotic variance of the estimator of the slope $\beta_{1}$ is quite complicated.
+Lemma 5.1 shows that even in a very simple framework (only one covariate) the asymptotic variance of the estimator of the slope $\beta_{1}$ is quite complicated.
 It also reveals that minimizing the asymptotic variance $\widehat{\beta}_{1}$ with respect to the weights $w_{k}$ is not a simple task.
 6 Bootstrapping for confidence intervals
-Since, as noted in the previous section, it is complicated to compute explicitly the asymptotic variance matrix of $\widehat{\boldsymbol{\beta}}^{M}$ and $\widehat{\boldsymbol{\beta}}^{W}$, we consider stratified bootstrap approaches in order to build confidence sets for $\boldsymbol{\beta}$.
-Our bootstrap procedure takes account of the independence between the different groups $k=1,\ldots,K$, as well as the independence of inputs $(X_{1}^{k},\ldots,X_{d}^{k})$ and output $Y^{k}$ within each group, meaning more formally that, given $G=k$, the joint probability measure $\mathbb{P}^{k}$ of $Y$ and $\mathbf{X}$ is a product measure of the marginal measures $\mathbb{P}^{k}=\mathbb{P}_{Y}^{k}\otimes\mathbb{P}_{\mathbf{X}}^{k}$.
-Within each group $k$, we draw, with equal probability and with replacement, $n_{y}^{k}$ observations among $Y_{1}^{k},\ldots,Y_{n_{y}^{k}}^{k}$, We also draw independently, with equal probability and with replacement, $n_{x}^{k}$ observations among $\mathbf{X}_{1}^{k},\ldots,\mathbf{X}_{n_{x}^{k}}^{k}$.
+Since, as noted in the previous section, it is complicated to explicitly compute the asymptotic variance matrix of $\widehat{\boldsymbol{\beta}}^{M}$ and $\widehat{\boldsymbol{\beta}}^{W}$, we consider stratified bootstrap approaches in order to build confidence sets for $\boldsymbol{\beta}$.
+Our bootstrap procedure takes into account the independence between the different groups $k=1,\ldots,K$ as well as the independence of the inputs $(X_{1}^{k},\ldots,X_{d}^{k})$ and the output $Y^{k}$ within each group.
+More formally, given $G=k$, the joint probability measure $\mathbb{P}^{k}$ of $Y$ and $\mathbf{X}$ is a product measure of the marginal measures $\mathbb{P}^{k}=\mathbb{P}_{Y}^{k}\otimes\mathbb{P}_{\mathbf{X}}^{k}$.
+Within each group $k$, we draw, with equal probability and with replacement, $n_{y}^{k}$ observations among $Y_{1}^{k},\ldots,Y_{n_{y}^{k}}^{k}$.
+We also draw independently, with equal probability and with replacement, $n_{x}^{k}$ observations among $\mathbf{X}_{1}^{k},\ldots,\mathbf{X}_{n_{x}^{k}}^{k}$.
 We denote by $\mu_{Y}^{k*}$ and by $\boldsymbol{\mu}_{X}^{k*}$ the empirical means and by $\sigma_{Y,k}^{2,*}$ and by $\boldsymbol{\Gamma}_{X}^{k*}$ the empirical variances of $Y$ and $\mathbf{X}$ in these bootstrap samples.
-Bootstrapped estimators $\boldsymbol{\beta}^{M,*}$ and $\boldsymbol{\beta}^{W,*}$ of $\boldsymbol{\beta}$ can now be computed by replacing the empirical moments by the bootstrap moments in (3.2) and (3.4).
-To build confidence sets for the components of $\boldsymbol{\beta}$ based on previous bootstrap procedure, the bootstrap percentile technique, described in Chapter 4 of Shao and Tu, (1995), is simple to use.
-It can be noted that our estimators are smooth functions of sample means so that classical bootstrap theory applies (see for example Shao and Tu, (1995), Chapter 3).
-For simplicity, as in Proposition 4.1, we suppose that $n=n_{y}^{1}=\ldots=n_{y}^{K}=n_{x}^{1}=\ldots=n_{x}^{K}$.
-Because of the considered experimental design, our global "empirical distribution" is made of products of marginal empirical distributions, the bootstrap for means is almost surely consistent for the Kolmogorov metric, and with Theorem 3.1 in Shao and Tu, (1995) the same result holds for the estimators of $\boldsymbol{\beta}$ considered in this work.
-Then, the application of Theorem 4.1 in Shao and Tu, (1995) allows to conclude that bootstrap percentile method gives consistent confidence sets for each component of $\boldsymbol{\beta}$.
-Proposition 6.1.
-Suppose that $\mathbb{E}(Y^{2})<\infty$ and $\mathbb{E}\|\mathbf{X}\|^{2}<\infty$ and hypothesis $\mathbf{H}_{1}$ is fulfilled.
+Bootstrapped estimators $\boldsymbol{\beta}^{M,*}$ and $\boldsymbol{\beta}^{W,*}$ of $\boldsymbol{\beta}$ can now be computed by replacing the empirical moments by the bootstrap moments in (5) and (7).
+To build confidence sets for the components of $\boldsymbol{\beta}$ based on this bootstrap procedure, the bootstrap percentile technique described in Chapter 4 of Shao and Tu, (1995) can be applied.
+It can be noted that our estimators are smooth functions of the sample means, so classical bootstrap theory applies (see for example Shao and Tu, (1995), Chapter 3).
+For simplicity, we assume, as in Proposition 4.1, that $n=n_{y}^{1}=\ldots=n_{y}^{K}=n_{x}^{1}=\ldots=n_{x}^{K}$.
+Because of the experimental design considered, our global ”empirical distribution” consists of products of marginal empirical distributions, the bootstrap for the means is almost surely consistent for the Kolmogorov metric, and with Theorem 3.1 in Shao and Tu, (1995) the same result holds for the estimators of $\boldsymbol{\beta}$ considered in this work.
+The application of Theorem 4.1 in Shao and Tu, (1995) allows to conclude that the bootstrap percentile method gives consistent confidence bounds for each component of $\boldsymbol{\beta}$.
+Proposition 6.1
+Suppose that $\mathbb{E}(Y^{2})<\infty$ and $\mathbb{E}\|\mathbf{X}\|^{2}<\infty$ and the hypothesis $\mathbf{H}_{1}$ is fulfilled.
 Then as $n\to+\infty$, the bootstrap estimator $\boldsymbol{\beta}^{M,*}$ is strongly consistent for $\boldsymbol{\beta}$ in the Kolmogorov metric.
-Consider a risk $\alpha\in(0,1)$, for each component of $\boldsymbol{\beta}$, the bootstrap percentile approach provides, for a given nominal level $1-\alpha$, a consistent confidence set.
-We can also state a similar result for the estimators minimizing the Wasserstein distance, under slightly more restrictive moment conditions and compactness assumption.
-Proposition 6.2.
+Consider a risk $\alpha\in(0,1)$, then for a given nominal level $1-\alpha$, for each component of $\boldsymbol{\beta}$, the bootstrap percentile approach provides a consistent confidence set.
+We can also state a similar result for the estimators minimizing the Wasserstein distance, under slightly more restrictive moment conditions and a compactness assumption.
+Proposition 6.2
 Suppose that the assumptions of Proposition 4.2 are fulfilled.
 As $n\to+\infty$, the bootstrap estimator $(\boldsymbol{\beta}^{W,*},\sigma_{\epsilon}^{2,W,*})$ is strongly consistent for $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$ in the Kolmogorov metric.
-Consider a risk $\alpha\in(0,1)$, for each component of $\boldsymbol{\beta}$, the bootstrap percentile approach provides, for a given nominal level $1-\alpha$, a consistent confidence set.
-Propositions 6.1 and 6.2 are very important for practical applications since they ensure that even if we are not able to compute the explicit expression of the asymptotic variance of our estimators, we can still construct asymptotic confidence intervals thanks to simple bootstrap procedures.
+Consider a risk $\alpha\in(0,1)$, then for a given nominal level $1-\alpha$, for each component of $\boldsymbol{\beta}$, the bootstrap percentile approach provides a consistent confidence set.
+Propositions 6.1 and 6.2 are very important for practical applications since they ensure that even if we are unable to compute the explicit expression of the asymptotic variance of our estimators, the asymptotic confidence intervals can still be constructed with a simple bootstrap procedure.
 7 Simulation study
 7.1 Simulation design
-We performed a series of simulations in order to evaluate the finite sample performances of the proposed approaches on data that resemble in vivo data originating from real experiments on mice.
-The number of animals observed per group is chosen for simplicity $n=n_{y}^{1}=\cdots=n_{y}^{K}=n_{x}^{1}\cdots=n_{x}^{K}$, and thus the weights for every group are also considered to be equal, i.e.
+We performed a series of simulations to evaluate the finite sample performance of the proposed approaches on data resembling in vivo data from real experiments on mice.
+The number of animals observed per group is chosen for simplicity to be $n=n_{y}^{1}=\cdots=n_{y}^{K}=n_{x}^{1}\cdots=n_{x}^{K}$, and thus the weights for each group are also assumed to be equal, i.e.
 $w_{1}=\cdots=w_{K}=\frac{1}{K}$.
-For every animal $i\in\{1,\dots,n\}$ and every subpopulation $k\in\{1,\dots,K\}$, the predictor variable $X^{k}_{i}$ is Gaussian univariate: $X^{k}_{i}\sim\mathcal{N}(\mu_{X}^{k},\sigma_{X}^{2})$, where $\mu_{X}^{k}=9+k$.
+For each animal $i\in\{1,\dots,n\}$ and each subpopulation $k\in\{1,\dots,K\}$, the predictor variable $X^{k}_{i}$ is Gaussian univariate: $X^{k}_{i}\sim\mathcal{N}(\mu_{X}^{k},\sigma_{X}^{2})$, where $\mu_{X}^{k}=9+k$.
 We simulate the predicted variable independently as $Y^{k}_{i}=\beta_{0}+\beta_{1}X^{\prime k}_{i}+\epsilon^{k}_{i}$, where $X^{\prime k}_{i}\sim X^{k}_{i}$ and $\epsilon^{k}_{i}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$, with regression parameters $\beta_{0}=1$ and $\beta_{1}=2$.
-It should be noted that $X^{k}_{i}$ and $X^{\prime k}_{i}$ are independent, we thus recreate the situation of the predictor and predicted variables not being simultaneously observed.
+By making $X^{k}_{i}$ and $X^{\prime k}_{i}$ independent, we recreate the situation where the predictor and predicted variables are not observed simultaneously.
 The variable sets $X=(X^{k}_{i})_{1\leq i\leq n,1\leq k\leq K}$ and $Y=(Y^{k}_{i})_{1\leq i\leq n,1\leq k\leq K}$ are simulated $N_{sim}$ times.
-For each simulation, $N_{boot}$ bootstrap samples of the size $n$ are generated from $X^{k}=(X^{k}_{i})_{1\leq k\leq K}$ and $Y^{k}=(Y^{k}_{i})_{1\leq k\leq K}$ for each subpopulation $k\in\{1,\dots,K\}$ independently, then moment estimators $\mu_{X}^{k*}$ and $\mu_{Y}^{k*}$ are calculated.
-Finally, the bootstrap sample-based estimators $\boldsymbol{\beta}^{M,*}=(\beta_{0}^{M,*},\beta_{1}^{M,*})$ and $\boldsymbol{\beta}^{W,*}=(\beta_{0}^{W,*},\beta_{1}^{W,*})$ are calculated.
-Based on $N_{boot}$ estimates, we calculate 95% confidence intervals using the function quantile of the Python library NumPy.
-As a result, we obtain $N_{sim}$ confidence intervals for each regression parameter, that we use to calculate the following quantities of interest: coverage rate of the intervals, their average amplitude, and power, i.e.
-the proportion of intervals not including 0.
-Along with the bootstrap estimators, we also considered asymptotic confidence intervals in case of the method of moments, obtained by plugging the estimated values of the moments into the expression for the limit distribution presented in Lemma 5.1.
-Additionally, we estimate the confidence intervals for the parameter estimators of the regression on the means by group with a naive method, assuming that the deviation of the parameter estimator from the true value divided by the estimator’s standard error follows a Student’s t-distribution:
+For each simulation, $N_{boot}$ bootstrap samples of size $n$ are generated from $X^{k}=(X^{k}_{i})_{1\leq k\leq K}$ and $Y^{k}=(Y^{k}_{i})_{1\leq k\leq K}$ for each subpopulation $k\in\{1,\dots,K\}$ independently, then the moment estimators $\mu_{X}^{k*}$ and $\mu_{Y}^{k*}$ are computed.
+Finally, the bootstrap sample-based estimators $\boldsymbol{\beta}^{M,*}=(\beta_{0}^{M,*},\beta_{1}^{M,*})$ and $\boldsymbol{\beta}^{W,*}=(\beta_{0}^{W,*},\beta_{1}^{W,*})$ are computed.
+Based on the $N_{boot}$ estimates, we compute 95% confidence intervals using the quantile function from the Python library NumPy.
+As a result, we obtain $N_{sim}$ confidence intervals for each regression parameter, which we use to calculate the following quantities of interest: the coverage rate of the intervals, their average amplitude, and the power, i.e.
+the proportion of intervals that do not contain 0.
+In addition to the bootstrap estimators, we also considered asymptotic confidence intervals in the case of the method of moments, obtained by plugging the estimated values of the moments into the expression for the limit distribution presented in Lemma 5.1.
+In addition, we estimate the confidence intervals for the parameter estimators of the regression on the means per group with a naive method, assuming that the deviation of the parameter estimator from the true value divided by the standard error of the estimator follows a Student’s t-distribution:
 $\frac{\widehat{\beta}_{j}-\beta_{j}}{SE(\widehat{\beta_{j}})}\rightsquigarrow t_{K-(d+1)}\text{ for }j\in\{0,\dots,d\}.$
-Figure 7.1: The effect of different values of $\rho$ on the data, with $K=4$ and $\sigma_{X}^{2}=0.75$.
+Finally, we considered the case where the predictor and the predicted variable are observed simultaneously, i.e.
+$X^{k}_{i}$ and $Y^{k}_{i}$ are such that $Y^{k}_{i}=\beta_{0}+\beta_{1}X^{k}_{i}+\epsilon^{k}_{i}$.
+In this case the parameters are estimated with the classical linear regression approach, and the confidence intervals are obtained with a Student’s t-distribution.
+Throughout all simulations we fix the number of simulations $N_{sim}=500$ and the number of bootstrap samples $N_{boot}=500$.
+Multiple parameters are varied to study their effect.
+We take the number of animals $n\in\{10,30\}$, in particular to test whether inference is significantly affected when measurements are only available for a small number of animals, which is often the case in real experimental data.
+We consider the number of groups $K\in\{4,10\}$, where 4 is the number of groups often observed in real data, and 10 is a higher number that may produce sufficiently good results with the naive approach of approximating confidence intervals with the Student distribution.
+Additionally, the parameter $\sigma^{2}_{X}$ can be adjusted to control the extent to which the observations per group can be easily distinguished from each other.
+We set $\sigma^{2}_{X}\in\{0.75,2\}$, the first value corresponding to less overlap between groups and the second to more overlap.
+Finally, we introduce an additional parameter $\rho\in\mathbb{R}^{+}$ which controls the variance of the response to the variance of the noise ratio, i.e.
+$\rho=\frac{\sigma_{Y}}{\sigma_{\epsilon}}$, where $\sigma^{2}_{Y}=Var(Y^{k}_{i})$ for all $i$ and $k$.
+The choice of adjusting the signal-to-noise ratio rather than the amount of noise itself through $\sigma_{\epsilon}^{2}$ is motivated by the fact that $\sigma_{Y}$ depends on $\sigma_{X}$, so the same level of $\sigma_{\epsilon}$ cannot be interpreted in the same way for different values of $\sigma_{X}$.
+The variance of the noise can be expressed as follows $\sigma_{\epsilon}^{2}=\frac{\beta_{1}^{2}\sigma_{X}^{2}}{\rho^{2}-1}$.
+The values of $\rho$ are chosen to correspond to the realistic situation, namely a very noisy case and a slightly less noisy one: $\rho\in\{1.01,1.1\}$.
+The effect of different values of $\rho$ on the simulated response variable is illustrated in Figure 3.
+Figure 3: The effect of different values of $\rho$ on the data, with $K=4$ and $\sigma_{X}^{2}=0.75$.
 a) Boxplots constructed from the simulated values of $X^{k}_{i}$.
@@
 $\rho=1.01$.
-Throughout all simulations, we fix the number of simulations $N_{sim}=500$ and the number of bootstrap samples $N_{boot}=500$.
-Multiple parameters are varied to study their effect.
-We take the number of animals $n\in\{10,30\}$, in particular to test whether the inference is significantly impaired if the measurements are available only for a small number of animals, which is often the case for real experimental data.
-We consider the number of groups $K\in\{4,10\}$, 4 being the number of groups that is often observed in real data, and 10 being a higher number that may produce sufficiently good results with the naive approach to approximating confidence intervals with Student distribution.
-Additionally, the parameter $\sigma^{2}_{X}$ can be adjusted to control the extent to which the observations per group can be easily distinguished one from another.
-We set $\sigma^{2}_{X}\in\{0.75,2\}$, the first value corresponding to lower overlap between groups, and the second to higher overlap.
-Finally, we introduce an additional parameter $\rho\in\mathbb{R}^{+}$ controlling the variance of the response to the variance of noise ratio, i.e.
-$\rho=\frac{\sigma_{Y}}{\sigma_{\epsilon}}$, where $\sigma^{2}_{Y}=Var(Y^{k}_{i})$ for all $i$ and $k$.
-The choice of adjusting the noise to signal ratio instead of the quantity of noise itself through $\sigma_{\epsilon}^{2}$ is motivated by the fact that $\sigma_{Y}$ depends on $\sigma_{X}$, hence the same level of $\sigma_{\epsilon}$ cannot be interpreted the same way for different values of $\sigma_{X}$.
-The variance of the noise can be expressed as follows: $\sigma_{\epsilon}^{2}=\frac{\beta_{1}^{2}\sigma_{X}^{2}}{\rho^{2}-1}$.
-The values of $\rho$ are chosen to correspond to the realistic situation, namely a very noisy case and a slightly less noisy one: $\rho\in\{1.01,1.1\}$.
-The effect of different values of $\rho$ on the simulated response variable is illustrated on Figure 7.1.
 7.2 Results
-The results of the simulation study are presented in Supplementary Figure C.1.
-Firstly, the results for the moment approach in the asymptotic and bootstrap cases are very similar, thus confirming the effectiveness of the bootstrap approach.
-Concerning the comparison between the bootstrap procedure and the naive approach, it can be observed that in general the bootstrap estimators produce confidence intervals with smaller amplitudes and with higher power, at the expense of a slightly lower coverage rate.
-Whereas the empirical coverage rates in all cases are close to the nominal one (95%), the extent to which the amplitudes are smaller and the powers are bigger for the bootstrap estimators is very important in almost all cases.
+The results of the simulation study are shown in Supplementary Figure 7.
+Firstly, the results for the moment approach are very similar in the asymptotic and bootstrap cases, confirming the effectiveness of the bootstrap approach.
+When comparing the bootstrap procedure with the naive approach, it can be observed that the bootstrap estimators generally produce confidence intervals with smaller average amplitudes and higher power at the expense of a slightly lower coverage rate.
+Whereas the empirical coverage rates are close to the nominal one (95%) in all cases, the extent to which the average amplitudes are smaller and the powers are greater for the bootstrap estimators is very important in almost all cases.
 This implies that the naive approach based on the Student’s distribution is more likely to produce false negatives in terms of significance.
-This trend is further amplified by the number of groups parameter: whereas the results are overall worsened with the decrease in either the number of animals or groups, it is the case with the small number of groups that demonstrates the biggest difference in the approaches (the case with few groups and animals is presented in Figure 7.2).
-Indeed, in almost all cases within the tables with $K=4$ we observe the amplitudes approximately twice as important for the naive approach, and a similar trend in terms of lower powers.
-The latter result is important since lower power implies bigger probability of not detecting a significant relationship between the predictor and the predicted variables, when it is actually present.
-Overall, these results mean that the proposed bootstrap estimators are more effective if the experimental design entails a small number of groups.
-Figure 7.2: a) Coverage rates, b) amplitudes, and c) powers of the confidence intervals for the estimators of $\beta_{1}$ obtained from 500 simulations, with number of groups $K=4$ and number of animals per group $n=10$.
-The columns of the tables represent combinations of parameters with respect to the group dispersion parameter $\sigma^{2}_{X}$ and the noise to signal ratio $\rho$.
-The lines indicate the method used to estimate the confidence intervals: "mm (asymp)" stands for the method of moments with asymptotic confidence intervals, "mm (boot)" for the method of moments with bootstrap, "ot (boot)" for the optimal transport method with bootstrap, and "mm (student)" for the naive linear regression on means approach based on Student’s distribution.
-Concerning the remaining two parameters, as expected, in general the best results are obtained with lower $\sigma^{2}_{X}$ and higher $\rho$.
-In most cases, the results for the naive and the bootstrap estimators are either both good or both bad in terms of power, with the latter being slightly better.
-A particularly complicated case can be distinguished, with high dispersion, high noise level, few groups and few animals, where all estimators drastically fail: we observe almost equally bad powers (0.1 for both bootstrap estimators and 0.09 for the naive estimator), despite the significant difference in amplitudes.
-On the other hand, we can also distinguish two cases where the powers of the bootstrap estimators are more than 90%, whereas those of the naive estimator are under 50%: in both cases there are 4 groups and high noise, in the first case there are only 10 animals but lower dispersion, in the second case high dispersion level is compensated but a higher number of animals.
-This implies that if the underlying distributions per group are characterized by a reasonable amount of overlap, or a significant overlap is compensated by having more observations, the bootstrap estimators manage to detect the significant relationship in most cases, unlike the naive estimator.
-Lastly, it can be observed that the estimator based on optimal transport produces confidence intervals with slightly smaller amplitudes compared to the method of moments estimator.
-The difference appears to be relatively more important in the cases with higher group overlap $\sigma^{2}_{X}=2$.
+This trend is further strengthened by the parameter encoding the number of groups: while the overall results worsen with a decrease in either the number of animals or the number of groups, it is the case with a small number of groups that shows the greatest difference between the approaches (the case with few groups and animals is shown in Figure 4).
+Indeed, in almost all cases within the tables with $K=4$ we observe that the average amplitudes are approximately twice as important for the naive approach, and a similar trend in terms of lower powers.
+The latter result is important because lower power implies a higher probability of not detecting a significant relationship between the predictor and the predicted variables, if it is indeed present.
+Overall, these results imply that the proposed bootstrap estimators are more effective when the experimental design involves a small number of groups.
+Figure 4: a) Coverage rates, b) average amplitudes, and c) powers of the confidence intervals for the estimators of $\beta_{1}$ obtained from 500 simulations, with number of groups $K=4$ and number of animals per group $n=10$.
+The columns of the tables indicate simulation scenarios with different combinations of parameters: scenario S1 with lower group overlap ($\sigma^{2}_{X}=0.75$) and higher signal-to-noise ratio ($\rho=1.1$), S2 with higher group overlap ($\sigma^{2}_{X}=2$) and higher signal-to-noise ratio ($\rho=1.1$), S3 with lower group overlap ($\sigma^{2}_{X}=0.75$) and lower signal-to-noise ratio ($\rho=1.01$), and S4 with higher group overlap ($\sigma^{2}_{X}=2$) and lower signal-to-noise ratio ($\rho=1.01$).
+The lines indicate the method used to estimate the confidence intervals: ”mm (asymp)” stands for the method of moments with asymptotic confidence intervals, ”mm (boot)” for the method of moments with bootstrap, ”ot (boot)” for the optimal transport method with bootstrap, ”mm (student)” for the naive linear regression on means approach based on Student’s distribution, and ”simultaneous” for the classical linear regression estimation in the case where the predictor and the predicted variable are observed simultaneously.
+Regarding the remaining two parameters, as expected, the best results are generally obtained with lower $\sigma^{2}_{X}$ and higher $\rho$.
+In most cases, the results for the naive and bootstrap estimators are either both good or both bad in terms of power, with the latter being slightly better.
+A particularly complicated case can be distinguished, with high overlap, high noise, few groups and few animals, where all estimators fail drastically: we observe almost equally bad powers (0.1 for both bootstrap estimators and 0.09 for the naive estimator), despite the significant difference in average amplitudes.
+On the other hand, we can also distinguish two cases where the powers of the bootstrap estimators are above 90% while those of the naive estimator are below 50%: in both cases there are 4 groups and high noise, in the first case there are only 10 animals but less overlap, in the second case a high level of overlap is compensated by a higher number of animals.
+This means that if the underlying distributions per group are characterized by a reasonable amount of overlap, or if a significant overlap is compensated by having more observations, the bootstrap estimators manage to detect the significant relationship in most cases, unlike the naive estimator.
+Figure 5: Distributions of amplitudes of confidence intervals obtained with different methods based on 500 simulations under scenarios S2 (a) and S3 (b), with $K=4$ and $n=10$.
+Concerning the comparison with the simultaneous case, as expected, the latter globally produces better results than any method in the context of non-simultaneous design.
+In particular, average amplitudes are systematically 10%-40% smaller compared to our methods in the non-simultaneous case.
+However, the difference seems to be less important than that between our approaches and the naive approach.
+Boxplots of the amplitude distributions for the most challenging combination of $K$ and $n$ under the most interesting scenarios (S2 and S3) are shown in Figure 5.
+The figure confirms the fact that the amplitude distributions obtained with our methods are more similar to those obtained in the simultaneous case than to those obtained with the naive approach in the non-simultaneous case.
+Moreover, the powers are almost as high as those obtained in the simultaneous case in 3 out of 4 scenarios.
+The only scenario where this difference is striking is the one with little noise and high group overlap.
+It can be concluded that if the groups are separable, in the case where the predictor and the predicted variable are not observed simultaneously, with our approach we can detect a significant relationship with a success comparable to the case when the variables are observed simultaneously.
+It should be noted that if the data are too noisy, all methods fail to detect the significance in all cases.
+Finally, it can be observed that the estimator based on optimal transport produces confidence intervals with slightly smaller average amplitudes compared to the method of moments estimator.
+The difference appears to be relatively more important in cases with higher group overlap $\sigma^{2}_{X}=2$.
 However, the powers are not affected by this difference.
 This may be explained by a more important bias associated with the optimal transport estimator.
-The estimator will likely produce better results in terms of the power than the method of moments estimator if the bias is corrected.
+The estimator is likely to produce better results in terms of power than the method of moments estimator when the bias is corrected.
 8 Application to real data
-In order to illustrate the proposed estimators on real data, we studied the data previously mentioned in Section 1, obtained from experiments conducted on mice in order to assess the adverse effects induced in the context of different irradiated volume.
-In the course of these experiments, mice were exposed to either stereotactic body radiation therapy (SBRT) with different beam sizes at 90 Gy on the left lung, or whole-thorax irradiation (WTI) at 19 Gy.
-For one mouse cohort, the expressions of tho pro-inflammatory genes (IL6 and TNF) were measured, for the other cohort the measurements of the thickness of the alveolar septas were made due to its role as a severity measure of radio-induced pulmonary lesions.
-In the case of the SBRT, the measurements were taken in multiple locations: the irradiated patch (inside the irradiation field), the remaining part of the left lung referred to as ipsilateral lung, and the right lung (contralateral lung).
-A control condition is also considered, where the gene expressions are measured without prior radiation exposure.
-The goal of this statistical analysis is to determine whether there is a statistical association between the gene expression as a predictor and the septal thickening as as outcome.
-Our approach is applied since the variables are measured on different animals, but within each irradiation condition there are shared groups in terms of the measurement time points.
-Table 8.1: Results of estimation of the linear regression slope predicting septal thickening with the pro-inflammatory genes expression, with three methods, for control (no irradiation), WTI and SBRT with different beam sizes, with measurements taken in different parts of lungs.
+To illustrate the proposed estimators on real data, we examined the data mentioned in Section 1, obtained from experiments conducted in mice to assess the adverse effects induced in the context of different irradiated volumes.
+In these experiments (see Bertho et al., (2020) for a more detailed presentation), mice were exposed to either stereotactic body radiation therapy (SBRT) with different beam sizes at 90 Gy to the left lung, or whole thorax irradiation (WTI) at 19 Gy.
+For one cohort of mice, the expression of pro-inflammatory genes (IL6 and TNF) was measured, and for the other cohort, the thickness of the alveolar septum was measured as a measure of the severity of radiation-induced lung lesions.
+In the case of SBRT, measurements were taken at multiple sites: the irradiated patch (within the irradiation field), the remaining part of the left lung, referred to as the ipsilateral lung, and the right lung (contralateral lung).
+A control condition is also included where gene expression is measured without prior exposure to radiation.
+The goal of this statistical analysis is to determine whether there is a statistical association between gene expression as a predictor and septal thickening as an outcome.
+Our approach is applied because the variables are measured on different animals, but within each irradiation condition there are common groups in terms of measurement time points.
+To ensure comparability of the results for different treatment conditions and genes, the data were centered and reduced with respect to the global mean and standard deviation prior to estimation.
+The linear regression parameters were then estimated with three estimators in the same way as in the simulation study in Section 7.
+The focus is on the estimation of the slope parameter $\beta_{1}$.
+The results are presented in Table 1, which contains the estimates of $\beta_{1}$ as well as the estimated confidence intervals for the slope estimator and the corresponding test result for the significance of the estimated relationship.
+Table 1: Results of estimation of the linear regression slope predicting septal thickening with the pro-inflammatory genes expression, with three methods, for control (no irradiation), WTI and SBRT with different beam sizes, with measurements taken in different parts of lungs.
 Method of moments Optimal transport Lin. Reg. on means
@@
 irradiation TNF 2.35 (0.57, 4.73) 1.67 (0.53, 1.88) 2.35 (-3.86, 8.57)
-In order to ensure the comparability of the results for different treatment conditions and genes, the data were centered and reduced with respect to the global mean and standard deviation prior to estimation.
-The linear regression parameters were then estimated with three estimators the same way it was done in the simulation study in Section 7.
-The focus is placed on estimating the slope parameter $\beta_{1}$ in particular.
-The results are presented in Table LABEL:tab:vivo_results, containing the estimations of $\beta_{1}$ as well as the estimated confidence intervals for the slope estimator, and the corresponding test result on the significance of the estimated relationship.
-Major differences between the results for the naive and the bootstrap estimators can be observed in terms of the detected significance.
-On the one hand, the relationship between the pro-inflammatory genes and septal thickening has been identified by all methods in the case of whole-thorax irradiation (both genes for the bootstrap estimators, and only one for the naive estimator), which is an expected result.
-The results are also consistent for all methods in the case without radiation exposure (control), where no significant relationship has been identified, as it is expected.
-On the other hand, we expect to identify a strong correlation in the case of the measurements taken directly from the irradiated patch.
-This is only the case for the bootstrap estimators, but not for the naive one.
-This results is in accordance with the results we obtain with simulated data: the confidence intervals are often overestimated with the naive approach, which may result in false negatives in terms of significance.
-Among the SBRT irradiation configurations, only the 3mm beam size has shown a significant correlation between inflammatory genes and septal thickening.
-These results are consistent with literature, indicating it as the beam size starting with which the long-term lesions appear (Bertho et al.,, 2020).
-Multiple significant associations have been identified with the bootstrap estimators in the ipsilateral lung and in the right lung for the beam size of 3 mm.
-Finally, among the cases where a significant relationship has been detected, the estimated values of the slope are always positive, which indicates a general radio-induced up-regulation trend.
-These values are in general bigger in case of whole-thorax irradiaion and within the patch than for the ipsilateral or right lung for SBRT, which suggests a stronger correlation between the inflammatory process and lung injuries under high dose/volume irradiation conditions.
-These results, which are in line with the biological knowledge, could not have been established within the framework of classical statistical regression approaches due to non-simultaneous observations.
-This effect is illustrated in Figure 8.1 on the example of the linear model prediction made for the gene IL6.
-Figure 8.1: Linear model prediction of septal thickness based on IL6 expression, plotted for different locations and beam sizes, with the results from two bootstrap estimators.
+Considerable differences can be observed between the results for the naive and the bootstrap estimators with respect to the significance found.
+On the one hand, the relationship between the pro-inflammatory genes and septal thickening was identified by all methods in the case of whole thorax irradiation (both genes for the bootstrap estimators and only one gene for the naive estimator), which is an expected result.
+The results are also consistent for all methods in the case of no radiation exposure (control), where no significant association was identified, as expected.
+On the other hand, we expect to identify a strong correlation in the case of measurements taken directly from the irradiated patch.
+This is the case only for the bootstrap estimators, but not for the naive one.
+This result is in accordance with the results obtained with simulated data: the confidence intervals are often overestimated with the naive approach, which can lead to false negatives in terms of significance.
+Among the SBRT irradiation configurations, only the 3 mm beam size showed a significant correlation between inflammatory genes and septal thickening.
+These results are consistent with the literature, indicating that this is the beam size from which the long-term lesions appear (Bertho et al.,, 2020).
+Several significant associations were identified with the bootstrap estimators in the ipsilateral lung and in the right lung for the beam size of 3 mm.
+Finally, in the cases where a significant relationship was detected, the estimated values of the slope are always positive, indicating a general radio-induced upregulation trend.
+These values are generally greater for whole thorax irradiation and within the patch than for the ipsilateral or right lung for SBRT, suggesting a stronger correlation between the inflammatory process and lung injury under high dose/volume irradiation conditions.
+These results, which are in line with biological knowledge, could not have been obtained using classical statistical regression approaches due to non-simultaneous observations.
+This effect is illustrated in Figure 6 using the example of linear model prediction for the gene IL6.
+Figure 6: Linear model prediction of septal thickness based on IL6 expression, plotted for different locations and beam sizes, with the results from two bootstrap estimators.
 9 Discussion
-This work focuses on a statistical framework designed for extracting dependencies from experiments, specifically introducing linear regression estimators in the context where the predictor and the predicted variables are never jointly observed.
-In this work we chose the basic linear multivariate setting, prioritizing simplicity and computational feasibility.
-Particularly, the estimator based on the method of moments makes no hypotheses on data distribution and can be calculated explicitly.
-The estimator based on optimal transport includes a simple optimization problem, and is based on the Gaussian form of the Wasserstein distance but does not technically require the data to be Gaussian, seeking to approximate them with Gaussian variables in whatever case.
-However, these approaches are inapplicable in the cases where linear relationship hypothesis cannot be satisfied.
-For instance, it is the case with predicting survival data with some continuous biomarker, which is of particular interest in research into radio-induced adverse effects.
-To be able to consider such scenarios, our model can be extended to a more general case, namely with generalized linear model.
-The optimal transport estimator appears promising in this context given the fact that the Wasserstein distance allows to compare probability distributions of different nature (for example, continuous and discrete).
-Another further research direction lies in investigating alternative methods based on integrated likelihood and Bayesian approaches, which are likely to produce better results in many cases but require putting priors on distributions.
-Finally, it would be of interest to work on improving theoretical properties of our estimators for finite samples, namely correcting the negative bias that appears for both estimators.
-The latter is particularly important in the case of the optimal transport estimator, which may be arising naturally with the Wasserstein distance (Wasserman,, 2023).
-Correcting this bias would considerably improve the estimator, making it competitive with the approaches mentioned earlier that make numerous assumptions on the data.
-Software availability
-Python code and the experimental data are freely available at github.com/parsenteva/vivo_lm.
+This work focuses on a statistical framework designed to extract dependencies from experiments, specifically introducing linear regression estimators in the context where the predictor and predicted variables are not jointly observed but share a common observed categorical variable.
+In this work we have chosen the basic linear multivariate setting, prioritizing simplicity and computational feasibility.
+In particular, the estimator based on the method of moments makes no hypotheses about the data distribution and can be computed explicitly.
+The optimal transport estimator involves a simple optimization problem and is based on the Gaussian form of the Wasserstein distance, but does not technically require the data to be Gaussian, seeking to approximate them with Gaussian variables in any case.
+The proposed bootstrap procedure produces confidence intervals for the regression parameters that are smaller than those obtained with the naive approach, while preserving a high coverage rate.
+In practice, this allows better detection of significant effects in cases where the sample size is small, which is often the case in in vivo experiments.
+However, these approaches are not applicable in cases where the linear relationship hypothesis cannot be satisfied.
+For example, this is the case when predicting survival data with some continuous biomarkers, which is of particular interest in the research on the adverse effects of radiation.
+To be able to consider such scenarios, our model can be extended to a more general case, namely with a generalized linear model.
+The optimal transport estimator seems promising in this context, given the fact that the Wasserstein distance allows to compare probability distributions of different nature (e.g.
+continuous and discrete).
+Another potential direction of research is to investigate alternative methods based on integrated likelihood and Bayesian approaches, which are likely to produce better results in many cases, but require the imposition of priors on distributions.
+Finally, it would be of interest to work on improving the theoretical properties of our finite sample estimators, namely the correction of the negative bias that appears for both estimators.
+The latter is particularly important in the case of the optimal transport estimator, which can arise naturally with the Wasserstein distance (e.g.
+Manole et al., (2024)).
+Correcting this bias would considerably improve the estimator, making it competitive with the aforementioned approaches, which make numerous assumptions about the data.
 Acknowledgements
-The authors gratefully acknowledge the funding from the European Union through the PO FEDER-FSE Bourgogne 2014/2020 programs as part of the project ModBioCan2020, and from the Institut de Radioprotection et de Sûreté nucléaire as part of the project ROSIRIS.
-The authors also warmly thank Dr.
-Agnès François and Dr.
-Annaïg Bertho for the experimental data, as well as Dr.
+The authors gratefully acknowledge the funding from the European Union through the PO FEDER-FSE Bourgogne 2014/2020 programs as part of the ModBioCan2020 project, and from the Institut de Radioprotection et de Sûreté nucléaire as part of the ROSIRIS project.
+The authors would also like to express their gratitude to Dr.
+Agnès François, Dr.
+Annaïg Bertho, Dr.
+Morgane Dos Santos and Dr.
+Fabien Milliat for the experimental data.
+Finally, the authors warmly thank Dr.
+Agnès François for her help in the biological interpretation of the data, as well as Dr.
 Patrick Tardivel for his numerous remarks that helped to improve the presentation of the manuscript.
-Conflict of Interest: none declared.
+Supporting Information
+The Python code is freely available at github.com/parsenteva/vivo_lm. The data that support the findings in this paper are available upon request from Dr. Agnès François (agnes.françois@irsn.fr).
 Appendix A Some classical theorems in asymptotic statistics
 A proof of the classical continuous mapping theorem can be found in van der Vaart, (1998) (Theorem 2.3).
-Theorem A.1.
+Theorem A.1
 (Continuous mapping theorem).
@@
 We also recall some well known results that are useful to show the consistency of estimators $\widehat{\theta}_{n}$ defined as the minimizers of functionals $Q_{n}(\theta)$ which have some regularity properties at the limit.
-Theorem A.2.
+Theorem A.2
 (Lemma 2.9 in Newey and McFadden, (1994))
@@
 $\sup_{\theta\in\Theta}\left|Q_{n}(\theta)-Q_{0}(\theta)\right|\to 0\ \mbox{ inprobability.}$.
-Theorem A.3.
+Theorem A.3
 (Theorem 2.1 in Newey and McFadden, (1994))
@@
 Under additional hypotheses, we also get the asymptotic normality of the sequence of estimators $\widehat{\theta}_{n}$ of $\theta_{0}$.
-we denote by $\nabla_{00}Q_{n}(\theta)$ the Hessian matrix of functional $Q_{n}$ evaluated at $\theta$.
-Theorem A.4.
+We denote by $\nabla_{00}Q_{n}(\theta)$ the Hessian matrix of the functional $Q_{n}$ evaluated at $\theta$.
+Theorem A.4
 (Theorem 3.1 in Newey and McFadden, (1994))
@@
 We also recall the central limit theorem for bootstrap means (see Theorem 23.4 in van der Vaart, (1998) for a proof).
-Theorem A.5.
-(CLT for bootstrap means)
+Theorem A.5
+(Central limit theorem for the bootstrap means)
 Let $X_{1},X_{2},\ldots$ be i.i.d.
@@
 Appendix B Proofs
-Proof.
-of Lemma 4.1
+Proofof Lemma 4.1
 First note that the assumptions ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{2})<+\infty$ ensure the existence of $\sigma_{Y}^{2}$ and $\boldsymbol{\Gamma}_{X}$.
@@
 The law of large numbers gives that $\widehat{\boldsymbol{\Gamma}}_{X}^{2}\to\boldsymbol{\Gamma}_{X}^{2}$ and $\widehat{\sigma}_{Y}^{2}\to\sigma_{Y}^{2}$ in probability and we deduce, with another application of the continuous mapping theorem, that $\widehat{\sigma}_{Y}^{2}-\widehat{\boldsymbol{\beta}}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{2}\widehat{\boldsymbol{\beta}}\to\sigma_{Y}^{2}-\boldsymbol{\beta}^{\top}\boldsymbol{\Gamma}_{X}^{2}\boldsymbol{\beta}=\sigma_{\epsilon}^{2}$ in probability as $n_{\min}$ tends to infinity.
-∎
-Proof.
-of Lemma 4.2
+$\Box$
+Proofof Lemma 4.2
 The proof is based on Lemma 2.9 and Theorem 2.1 in Newey and McFadden, (1994), which are recalled in Appendix A.
@@
 Because $\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\alpha}_{-0}-\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}=\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\left(\boldsymbol{\alpha}_{-0}-\boldsymbol{\gamma}_{-0}\right)+\left(\boldsymbol{\alpha}_{-0}-\boldsymbol{\gamma}_{-0}\right)^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}$ we have, for some constant $C_{2,k}>0$,
-$\left|\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\alpha}_{-0}-\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}\right|$ $\leq C_{2,k}\left\|\widehat{\boldsymbol{\Gamma}}_{X}^{k}\right\|_{sp}\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|.$ (B.1)
+$\left|\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\alpha}_{-0}-\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}\right|$ $\leq C_{2,k}\left\|\widehat{\boldsymbol{\Gamma}}_{X}^{k}\right\|_{sp}\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|.$ (11)
 Using now the fact that function $x\mapsto\sqrt{x}$ is concave and differentiable, we have for $x>0$ and $y>0$ that $\sqrt{y}\leq\sqrt{x}+\frac{y-x}{2\sqrt{x}}$.
@@
 Consequently, we have $|\sqrt{y}-\sqrt{x}|\leq\frac{|x-y|}{2\min(\sqrt{x},\sqrt{y})}$ and we deduce that,
-$\left|\sqrt{\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\alpha}_{-0}+\sigma_{\alpha}^{2}}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma_{\gamma}^{2}}\right|$ $\leq B_{n}^{k}\left(\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|+|\sigma_{\alpha}^{2}-\sigma_{\gamma}^{2}|\right)$ (B.2)
+$\left|\sqrt{\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\alpha}_{-0}+\sigma_{\alpha}^{2}}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma_{\gamma}^{2}}\right|$ $\leq B_{n}^{k}\left(\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|+|\sigma_{\alpha}^{2}-\sigma_{\gamma}^{2}|\right)$ (12)
 where $B_{n}^{k}=O_{p}(1)$.
 Combining previous inequalities, we get
-$\left|\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi_{n}(\boldsymbol{\alpha},\sigma_{\alpha}^{2})\right|$ $\leq\left(\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|+|\sigma_{\alpha}^{2}-\sigma_{\gamma}^{2}|\right)\sum_{k=1}^{K}\pi_{k}\left(B_{n}^{k}+A_{n}^{k}\right),$ (B.3)
+$\left|\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi_{n}(\boldsymbol{\alpha},\sigma_{\alpha}^{2})\right|$ $\leq\left(\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|+|\sigma_{\alpha}^{2}-\sigma_{\gamma}^{2}|\right)\sum_{k=1}^{K}\pi_{k}\left(B_{n}^{k}+A_{n}^{k}\right),$ (13)
 with $\sum_{k=1}^{K}\pi_{k}\left(B_{n}^{k}+A_{n}^{k}\right)=O_{p}(1)$. As a result, it can be deduced from Lemma 2.9 in Newey and McFadden, (1994) that
@@
 We conclude the proof by recalling that $\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$ attains its unique minimum at $(\boldsymbol{\beta},\sigma^{2}_{\epsilon})\in\Theta$ if assumption $\mathbf{H}_{1}$ is fulfilled, so that $(\widehat{\boldsymbol{\beta}}^{W},\widehat{\sigma}^{2,W})\to(\boldsymbol{\beta},\sigma^{2}_{\epsilon})$ in probability in view of Theorem 2.1 in Newey and McFadden, (1994).
-∎
-Proof.
-of Proposition 4.1
+$\Box$
+Proofof Proposition 4.1
 The central limit theorem applies directly to the independent sequences of independent random variables $(\mathbf{X}_{1}^{1},\cdots,\mathbf{X}_{n}^{1}),\ldots,(\mathbf{X}_{1}^{K},\cdots,\mathbf{X}_{n}^{K})$ and $(Y_{1}^{1},\cdots,Y_{n}^{1}),,\ldots,(Y_{1}^{K},\cdots,Y_{n}^{K})$ so that, as $n$ tends to infinity
-$\sqrt{n}\begin{pmatrix}\widehat{\boldsymbol{\mu}}_{X}^{1}-\boldsymbol{\mu}_{X}^{1}\\ \vdots\\ \widehat{\boldsymbol{\mu}}_{X}^{K}-\boldsymbol{\mu}_{X}^{K}\\ \widehat{\mu}_{Y}^{1}-\mu_{Y}^{1}\\ \vdots\\ \widehat{\mu}_{Y}^{K}-\mu_{Y}^{K}\end{pmatrix}$ $\rightsquigarrow\mathcal{N}\left(0,\boldsymbol{\Gamma}_{\mu}\right)$ (B.4)
+$\sqrt{n}\begin{pmatrix}\widehat{\boldsymbol{\mu}}_{X}^{1}-\boldsymbol{\mu}_{X}^{1}\\ \vdots\\ \widehat{\boldsymbol{\mu}}_{X}^{K}-\boldsymbol{\mu}_{X}^{K}\\ \widehat{\mu}_{Y}^{1}-\mu_{Y}^{1}\\ \vdots\\ \widehat{\mu}_{Y}^{K}-\mu_{Y}^{K}\end{pmatrix}$ $\rightsquigarrow\mathcal{N}\left(0,\boldsymbol{\Gamma}_{\mu}\right)$ (14)
 where $\boldsymbol{\Gamma}_{\mu}$ is a block diagonal matrix, with diagonal elements $(\boldsymbol{\Gamma}_{X}^{1},\ldots,\boldsymbol{\Gamma}_{X}^{K},\sigma_{Y,1}^{2},\ldots,\sigma_{Y,K}^{2})$, with $\boldsymbol{\Gamma}_{X}^{k}=\mbox{Var}(\mathbf{X}|G=k)={\mathbb{E}}\left(\mathbf{X}^{k}(\mathbf{X}^{k})^{\top}\right)-\boldsymbol{\mu}_{X}^{k}(\boldsymbol{\mu}_{X}^{k})^{\top}$ and $\sigma_{Y,k}^{2}=\mbox{Var}(Y|G=k)$.
@@
 Application $g$ is differentiable at $\boldsymbol{\theta}=(\boldsymbol{\mu}_{X}^{1},\ldots,\boldsymbol{\mu}_{X}^{K},\mu_{Y}^{1},\ldots,\mu_{Y}^{K})$, with non null Jacobian matrix denoted by $\mathbf{J}_{\theta}$ (see Chapter 8 and more particularly Theorem 8.3 in Magnus and Neudecker, (2019)).
-The application of the Delta method (see Theorem 3.1 in van der Vaart, (1998)) permits to get the asymptotic normality convergence result,
+The application of the Delta method (see Theorem 3.1 in van der Vaart, (1998)) allows to get the asymptotic normality convergence result,
 $\sqrt{n}\left(\widehat{\boldsymbol{\beta}}^{M}-\boldsymbol{\beta}\right)$ $\rightsquigarrow\mathcal{N}\left(0,\boldsymbol{\Gamma}_{\beta_{M}}\right),$
-where $\boldsymbol{\Gamma}_{\beta_{M}}=\mathbf{J}_{\theta}\boldsymbol{\Gamma}_{\mu}\mathbf{J}_{\theta}^{\top}$. ∎
-Proof.
-of Proposition 4.2
+where $\boldsymbol{\Gamma}_{\beta_{M}}=\mathbf{J}_{\theta}\boldsymbol{\Gamma}_{\mu}\mathbf{J}_{\theta}^{\top}$. $\Box$
+Proofof Proposition 4.2
 The proof consists in checking the different points of Theorem A.4.
-Point (i) is satisfied by the hypotheses, and the point (ii) follows directly from the fact $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2})$ is twice-differentiable in a neighborhood of $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$.
-To show that (iii) is fulfilled, we consider the following expansion, based on the empirical version of the gradient of $\varphi$,
-$\nabla\varphi_{n}$ $=\begin{pmatrix}-2\sum_{k=1}^{K}\pi_{k}\left(\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\\ -2\sum_{k=1}^{K}\pi_{k}\left[\left(\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\widehat{\boldsymbol{\mu}}_{X}^{k}+\left(\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-1\right)\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}\right]\\ \sum_{k=1}^{K}\pi_{k}\left(1-\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}\right)\end{pmatrix}$ (B.5)
-Since model (2.1) holds, $\nabla\varphi=0$ and $\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}=(\widehat{\mu}_{Y}^{k}-{\mu}_{Y}^{k})-\boldsymbol{\beta}_{-0}^{\top}\left(\widehat{\boldsymbol{\mu}}_{X}^{k}-\boldsymbol{\mu}_{X}^{k}\right)$, we thus deduce with (B.4) the asymptotic normality of the first component of the gradient $\nabla\varphi_{n}$, that is to say $\sqrt{n}\left(-2\sum_{k=1}^{K}\pi_{k}\left(\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\right)$ converges in distribution to a centered Gaussian distribution.
-As far as the second component is concerned, it can be noted that $\widehat{\boldsymbol{\Gamma}}_{X}^{k}$ converges in probability to $\boldsymbol{\Gamma}_{X}^{k}$ and by the continuous mapping theorem, $\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}\to\sigma_{Y,k}$ in probability.
+Point (i) is satisfied by the assumptions, and the point (ii) follows directly from the fact that $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2})$ is twice-differentiable in a neighborhood of $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$.
+To show that (iii) is fulfilled, we consider the following expansion, based on the empirical version of the gradient of $\varphi$:
+$\nabla\varphi_{n}$ $=\begin{pmatrix}-2\sum_{k=1}^{K}\pi_{k}\left(\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\\ -2\sum_{k=1}^{K}\pi_{k}\left[\left(\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\widehat{\boldsymbol{\mu}}_{X}^{k}+\left(\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-1\right)\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}\right]\\ \sum_{k=1}^{K}\pi_{k}\left(1-\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}\right)\end{pmatrix}$ (15)
+Since model (1) holds, $\nabla\varphi=0$ and $\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}=(\widehat{\mu}_{Y}^{k}-{\mu}_{Y}^{k})-\boldsymbol{\beta}_{-0}^{\top}\left(\widehat{\boldsymbol{\mu}}_{X}^{k}-\boldsymbol{\mu}_{X}^{k}\right)$, we thus deduce with (14) the asymptotic normality of the first component of the gradient $\nabla\varphi_{n}$, that is to say $\sqrt{n}\left(-2\sum_{k=1}^{K}\pi_{k}\left(\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\right)$ converges in distribution to a centered Gaussian distribution.
+As far as the second component is concerned, it can be noted that $\widehat{\boldsymbol{\Gamma}}_{X}^{k}$ converges in probability to $\boldsymbol{\Gamma}_{X}^{k}$, and by the continuous mapping theorem $\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}\to\sigma_{Y,k}$ in probability.
 It can also be noted that, under the moment condition ${\mathbb{E}}\left[\|\mathbf{X}\|^{4}|G=k\right]<\infty$, the central limit theorem gives that $\sqrt{n}\left(\widehat{\boldsymbol{\Gamma}}_{X}^{k}-\boldsymbol{\Gamma}_{X}^{k}\right)$ converges in distribution to a centered Gaussian multivariate distribution, and we deduce with the Cramer-Wold device, the continuous mapping theorem and Slutsky’s theorem that the second component of $\nabla\varphi_{n}$ multiplied by $\sqrt{n}$ also in distribution to a centered Gaussian random vector.
@@
 We first note that $\sigma_{Y,k}=\left(\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}\right)^{1/2}$ so that $\sigma_{Y,k}\left(\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}\right)^{-3/2}=\frac{1}{\sigma_{Y,k}^{2}}$ and $\mathbf{H}(\boldsymbol{\beta}_{-0})$ can be written in a simpler form,
-$\mathbf{H}(\boldsymbol{\beta}_{-0})=$ $2\sum_{k=1}^{K}\pi_{k}\left[\boldsymbol{\mu}_{X}^{k}\left(\boldsymbol{\mu}_{X}^{k}\right)^{\top}+\frac{1}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\left(\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}\right],$ (B.6)
-which is a definite positive matrix under hypothesis $\mathbf{H}_{1}$. Using a block matrix determinant formula, we have
-$\left|\mathbf{H}(\boldsymbol{\beta},\sigma_{\epsilon}^{2})\right|$ $=\begin{vmatrix}2&0\\ 0&\frac{1}{2}\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\end{vmatrix}\ \begin{vmatrix}\mathbf{H}(\boldsymbol{\beta}_{-0})-\mathbf{C}\begin{pmatrix}\frac{1}{2}&0\\ 0&\frac{2}{\sum_{k}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}}\end{pmatrix}\mathbf{C}^{\top}\end{vmatrix}$ (B.7)
-where $\mathbf{C}=\begin{pmatrix}2\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}&\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\end{pmatrix}$ and it only has to be verified that the second determinant at the righthand side of (B.7) is strictly positive.
+$\mathbf{H}(\boldsymbol{\beta}_{-0})=$ $2\sum_{k=1}^{K}\pi_{k}\left[\boldsymbol{\mu}_{X}^{k}\left(\boldsymbol{\mu}_{X}^{k}\right)^{\top}+\frac{1}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\left(\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}\right],$ (16)
+which is a positive definite matrix under the hypothesis $\mathbf{H}_{1}$. Using a block matrix determinant formula, we have
+$\left|\mathbf{H}(\boldsymbol{\beta},\sigma_{\epsilon}^{2})\right|$ $=\begin{vmatrix}2&0\\ 0&\frac{1}{2}\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\end{vmatrix}\ \begin{vmatrix}\mathbf{H}(\boldsymbol{\beta}_{-0})-\mathbf{C}\begin{pmatrix}\frac{1}{2}&0\\ 0&\frac{2}{\sum_{k}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}}\end{pmatrix}\mathbf{C}^{\top}\end{vmatrix}$ (17)
+where $\mathbf{C}=\begin{pmatrix}2\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}&\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\end{pmatrix}$, and it only has to be verified that the second determinant at the righthand side of (17) is strictly positive.
 We now have to show that
 $\mathbf{H}(\boldsymbol{\beta}_{-0})-\mathbf{C}\begin{pmatrix}\frac{1}{2}&0\\ 0&\frac{2}{\sum_{k}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}}\end{pmatrix}\mathbf{C}^{\top}=2\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}\left(\boldsymbol{\mu}_{X}^{k}\right)^{\top}-2\left(\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}\right)\left(\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}\right)^{\top}$
-$+2\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\left(\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}-\frac{2}{\sum_{k}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}}\left(\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)\left(\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}$ (B.8)
+$+2\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\left(\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}-\frac{2}{\sum_{k}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}}\left(\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)\left(\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}$ (18)
 is a positive matrix. We can remark that by Cauchy Schwarz inequality, for $\mathbf{u}\in\mathbb{R}^{d}$,
@@
 using the fact that $\sum_{k}(\sqrt{\pi_{k}})^{2}=1$.
-It can be noted that if $\mathbf{u}\neq 0$, previous inequality is strict unless $\mathbf{u}^{\top}\boldsymbol{\mu}_{X}^{1}=\cdots=\mathbf{u}^{\top}\boldsymbol{\mu}_{X}^{K}$, which can not happen under hypothesis $\mathbf{H}_{1}$.
-The second part at the righthand side of (B.8) is handled the same way.
+It can be noted that if $\mathbf{u}\neq 0$, previous inequality is strict unless $\mathbf{u}^{\top}\boldsymbol{\mu}_{X}^{1}=\cdots=\mathbf{u}^{\top}\boldsymbol{\mu}_{X}^{K}$, which cannot happen under the hypothesis $\mathbf{H}_{1}$.
+The second part at the righthand side of (18) is handled in the same way.
 We have
@@
 This is a direct consequence of the continuous mapping theorem, which gives us that for all $(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\in\mathcal{N}$, $\|\nabla_{00}\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\mathbf{H}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\|\to 0$ in probability, and the fact that third order partial derivatives of $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$ are bounded in probability for $(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$ so that Theorem A.2 can apply.
-∎
-Proof.
-of Lemma 5.1
+$\Box$
+Proofof Lemma 5.1
 Note that
@@
 $g(\mu_{X}^{1},\ldots,\mu_{X}^{K},\mu_{Y}^{1},\ldots,\mu_{Y}^{K})$ $=\frac{\sum_{k=1}^{K}w_{k}\mu_{X}^{k}\mu_{Y}^{k}-\mu_{X,w}\mu_{Y,w}}{\sum_{k=1}^{K}w_{k}(\mu_{X}^{k})^{2}-\left(\sum_{k=1}^{K}w_{k}\mu_{X}^{k}\right)^{2}},$
-$=\frac{\mbox{Cov}_{w}(X,Y)}{\mbox{Var}_{w}(X)},$ (B.9)
+$=\frac{\mbox{Cov}_{w}(X,Y)}{\mbox{Var}_{w}(X)},$ (19)
 with the notations $\mu_{X,w}=\sum_{k=1}^{K}w_{k}\mu_{X}^{k}$, $\mu_{Y,w}=\sum_{j=1}^{K}w_{j}\mu_{Y}^{j}$, $\mbox{Cov}_{w}(X,Y)=\sum_{k=1}^{K}w_{k}\mu_{X}^{k}\mu_{Y}^{k}-\mu_{X,w}\mu_{Y,w}$ and $\mbox{Var}_{w}(X)=\sum_{k=1}^{K}w_{k}(\mu_{X}^{k})^{2}-(\mu_{X,w})^{2}$.
-The gradient $\nabla g$ of $g$, evaluated at point $(\mu_{X}^{1},\ldots,\mu_{X}^{K},\mu_{Y}^{1},\ldots,\mu_{Y}^{K})$, is equal to
+The gradient $\nabla g$ of $g$, evaluated at the point $(\mu_{X}^{1},\ldots,\mu_{X}^{K},\mu_{Y}^{1},\ldots,\mu_{Y}^{K})$, is equal to
 $\nabla g$ $=\begin{pmatrix}\frac{w_{1}(\mu_{Y}^{1}-\mu_{Y,w})}{\mbox{Var}_{w}(X)}-\frac{2w_{1}\left(\mu_{X}^{1}-\mu_{X,w}\right)\mbox{Cov}_{w}(X,Y)}{(\mbox{Var}_{w}(X))^{2}}\\ \vdots\\ \frac{w_{K}(\mu_{Y}^{K}-\mu_{Y,w})}{\mbox{Var}_{w}(X)}-\frac{2w_{K}\left(\mu_{X}^{K}-\mu_{X,w}\right)\mbox{Cov}_{w}(X,Y)}{(\mbox{Var}_{w}(X))^{2}}\\ \frac{w_{1}\left(\mu_{X}^{1}-\mu_{X,w}\right)}{\mbox{Var}_{w}(X)}\\ \vdots\\ \frac{w_{K}\left(\mu_{X}^{K}-\mu_{X,w}\right)}{\mbox{Var}_{w}(X)}\end{pmatrix}.$
@@
 $=\frac{1}{(\mbox{Var}_{w}(X))^{2}}\sum_{k=1}^{K}w_{k}^{2}\left[\sigma^{2}_{X,k}\left(-\beta_{1}\left(\mu_{X}^{k}-\mu_{X,w}\right)\right)^{2}+\sigma^{2}_{Y,k}\left(\mu_{X}^{k}-\mu_{X,w}\right)^{2}\right]$
-$=\frac{1}{(\mbox{Var}_{w}(X))^{2}}\sum_{k=1}^{K}w_{k}^{2}\left(\mu_{X}^{k}-\mu_{X,w}\right)^{2}\left(\beta_{1}^{2}\sigma^{2}_{X,k}+\sigma^{2}_{Y,k}\right)$ (B.10)
+$=\frac{1}{(\mbox{Var}_{w}(X))^{2}}\sum_{k=1}^{K}w_{k}^{2}\left(\mu_{X}^{k}-\mu_{X,w}\right)^{2}\left(\beta_{1}^{2}\sigma^{2}_{X,k}+\sigma^{2}_{Y,k}\right)$ (20)
 remarking that $\beta_{1}=\mbox{Cov}_{w}(X,Y)/\mbox{Var}_{w}(X,Y)$, $\beta_{0}=\mu_{Y,w}-\beta_{1}\mu_{X,w}$ as well as $\beta_{0}=\mu_{Y}^{k}-\beta_{1}\mu_{X}^{k}$.
-Proof.
-of Proposition 6.1
+$\Box$
+Proofof Proposition 6.1
 The fact that the bootstrap estimator $\boldsymbol{\beta}^{M,*}$ is strongly consistent for $\boldsymbol{\beta}$ is a direct consequence of Theorem 3.1 in Shao and Tu, (1995), noting that
@@
 The fact that confidence sets based on the percentile approach are consistent is proved by checking the assumptions in Theorem 4.1 (iii) Shao and Tu, (1995), namely the bootstrap estimator $\boldsymbol{\beta}^{M,*}$ is consistent, $\widehat{\boldsymbol{\beta}}^{M}$ is consistent (Lemma 4.1), with asymptotic Gaussian distribution (Proposition 4.1).
-∎
-Proof.
-of Proposition 6.2
+$\Box$
+Proofof Proposition 6.2
 We denote by $\boldsymbol{\theta}_{0}=(\boldsymbol{\beta},\sigma_{\epsilon}^{2,W})$ the vector of true parameters, by $\widehat{\boldsymbol{\theta}}=(\boldsymbol{\beta}^{W},\sigma_{\epsilon}^{2})$ the sequence of minimum Wasserstein distance estimators and by $\boldsymbol{\theta}^{*}=(\boldsymbol{\beta}^{W,*},\sigma_{\epsilon}^{2,W,*})$ bootstrap estimators of $\boldsymbol{\theta}_{0}$.
 The vector of parameters $\boldsymbol{\theta}^{*}$ is the minimizer of functional $\varphi_{n}^{*}$ defined as follows,
-$\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2})$ $=\sum_{k=1}^{K}\pi_{k}\left[(\mu_{Y}^{k,*}-\gamma_{0}-\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\mu}_{X}^{k,*})^{2}+\left(\sigma_{Y,k}^{*}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k,*}\boldsymbol{\gamma}_{-0}+\sigma^{2}}\right)^{2}\right].$ (B.11)
+$\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2})$ $=\sum_{k=1}^{K}\pi_{k}\left[(\mu_{Y}^{k,*}-\gamma_{0}-\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\mu}_{X}^{k,*})^{2}+\left(\sigma_{Y,k}^{*}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k,*}\boldsymbol{\gamma}_{-0}+\sigma^{2}}\right)^{2}\right].$ (21)
 We first show with arguments similar to those employed in the proof of Lemma 4.2, that $\boldsymbol{\theta}^{*}$ is a consistent estimator for $\boldsymbol{\theta}_{0}$, based on the fact that $\varphi_{n}^{*}$ is a smooth function converging to $\varphi$ and the sample mean theorem for bootstrap (see for example Theorem 23.4 in van der Vaart, (1998)).
 Indeed, we first recall that for all $(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\in\Theta$, $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\to\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$ in probability, when $n_{\min}$ tends to infinity and
-$\left|\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\right|$ $\leq\left|\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\right|+\left|\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\right|.$ (B.12)
-Since the bootstrap means converge to the empirical ones we deduce with the continuous mapping theorem that $\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\to\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$ in probability, when $n_{\min}$ tends to infinity, so that $\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\to\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$.
-We also have, as in (B.3), where empirical means are replaced by bootstrap means,
-$\left|\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi_{n}^{*}(\boldsymbol{\alpha},\sigma_{\alpha}^{2})\right|$ $\leq\left(\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|+|\sigma_{\alpha}^{2}-\sigma_{\gamma}^{2}|\right)\sum_{k=1}^{K}\pi_{k}\left(B_{n}^{k,*}+A_{n}^{k,*}\right),$ (B.13)
+$\left|\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\right|$ $\leq\left|\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\right|+\left|\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\right|.$ (22)
+Since the bootstrap means converge to the empirical ones, we deduce with the continuous mapping theorem that $\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\to\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$ in probability, when $n_{\min}$ tends to infinity, so that $\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\to\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$.
+We also have, as in (13), where empirical means are replaced by the bootstrap means,
+$\left|\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi_{n}^{*}(\boldsymbol{\alpha},\sigma_{\alpha}^{2})\right|$ $\leq\left(\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|+|\sigma_{\alpha}^{2}-\sigma_{\gamma}^{2}|\right)\sum_{k=1}^{K}\pi_{k}\left(B_{n}^{k,*}+A_{n}^{k,*}\right),$ (23)
 for any $(\boldsymbol{\alpha},\sigma_{\alpha}^{2})\in\Theta$, with $\sum_{k=1}^{K}\pi_{k}\left(B_{n}^{k,*}+A_{n}^{k,*}\right)=O_{p}(1)$.
-As a result, we deduce from Lemma 4.2, inequality (B.12) and Lemma 2.9 in Newey and McFadden, (1994) that
+As a result, we deduce from Lemma 4.2, inequality (22) and Lemma 2.9 in Newey and McFadden, (1994) that
 $\sup_{(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\in\Theta}\left|\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\right|\to 0\quad\mbox{ in probability.}$.
@@
 By definition of $\widehat{\boldsymbol{\theta}}$ and Taylor expansion we have
-$\nabla\varphi_{n}(\widehat{\boldsymbol{\theta}})=\nabla\varphi_{n}(\boldsymbol{\theta}_{0})+\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\left(\widehat{\boldsymbol{\theta}}-\boldsymbol{\theta}_{0}\right)=0$ (B.14)
+$\nabla\varphi_{n}(\widehat{\boldsymbol{\theta}})=\nabla\varphi_{n}(\boldsymbol{\theta}_{0})+\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\left(\widehat{\boldsymbol{\theta}}-\boldsymbol{\theta}_{0}\right)=0,$ (24)
 where $\overline{\boldsymbol{\theta}}$ belongs, componentwise, to the segment between $\boldsymbol{\theta}_{0}$ and $\widehat{\boldsymbol{\theta}}$.
 We have a similar expansion for boostrap estimators, as well as
-$\nabla\varphi_{n}^{*}(\boldsymbol{\theta}^{*})=\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})+\nabla_{00}^{*}\varphi_{n}(\overline{\boldsymbol{\theta}}^{*})\left(\boldsymbol{\theta}^{*}-\boldsymbol{\theta}_{0}\right)=0$ (B.15)
-where $\overline{\boldsymbol{\theta}}^{*}$ belongs, componentwise, to the segment between $\boldsymbol{\theta}_{0}$ and $\boldsymbol{\theta}^{*}$. Combining (B.14) and (B.15), we deduce
+$\nabla\varphi_{n}^{*}(\boldsymbol{\theta}^{*})=\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})+\nabla_{00}^{*}\varphi_{n}(\overline{\boldsymbol{\theta}}^{*})\left(\boldsymbol{\theta}^{*}-\boldsymbol{\theta}_{0}\right)=0,$ (25)
+where $\overline{\boldsymbol{\theta}}^{*}$ belongs, componentwise, to the segment between $\boldsymbol{\theta}_{0}$ and $\boldsymbol{\theta}^{*}$. Combining (24) and (25), we deduce
 $\boldsymbol{\theta}^{*}-\widehat{\boldsymbol{\theta}}$ $=\left(\nabla_{00}^{*}\varphi_{n}(\overline{\boldsymbol{\theta}}^{*})\right)^{-1}\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})-\left(\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\right)^{-1}\nabla\varphi_{n}(\boldsymbol{\theta}_{0})$
-$=\left(\left(\nabla_{00}^{*}\varphi_{n}(\overline{\boldsymbol{\theta}}^{*})\right)^{-1}-\left(\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\right)^{-1}\right)\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})+\left(\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\right)^{-1}\left(\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})-\nabla\varphi_{n}(\boldsymbol{\theta}_{0})\right)$ (B.16)
-Noticing that $\nabla_{00}^{*}\varphi_{n}(\overline{\boldsymbol{\theta}}^{*})$ and $\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})$ both tend in probability to the same limit $\mathbf{H}(\boldsymbol{\beta},\sigma^{2}_{\epsilon})$ and we have, with similar arguments as those used in the proof of Proposition 4.2, that $\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})$ is $O_{p}(n^{-1/2})$, it can be deduce that
-$\boldsymbol{\theta}^{*}-\widehat{\boldsymbol{\theta}}$ $=\left(\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\right)^{-1}\left(\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})-\nabla\varphi_{n}(\boldsymbol{\theta}_{0})\right)+o_{P}(n^{-1/2}).$ (B.17)
-Using arguments similar to those employed in the expansion of $\nabla\varphi_{n}$ in the proof of Proposition 4.2, we make appear the difference between bootstrap means and empirical means or a differentiable functional of these quantities:
-$\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})-\nabla\varphi_{n}(\boldsymbol{\theta}_{0})=\footnotesize\begin{pmatrix}2\sum_{k=1}^{K}\pi_{k}\left((\widehat{\mu}_{Y}^{k}-{\mu}_{Y}^{k,*})-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\left(\widehat{\boldsymbol{\mu}}_{X}-\boldsymbol{\mu}_{X}^{k,*}\right)\right)\\ \begin{pmatrix}2\sum_{k=1}^{K}\pi_{k}\left[\left(\widehat{\mu}_{Y}^{k}-\mu_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\widehat{\boldsymbol{\mu}}_{X}^{k}+\left(\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-1\right)\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}\right]\\ -2\sum_{k=1}^{K}\pi_{k}\left[\left(\mu_{Y}^{k,*}-\mu_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\mu}_{X}^{k,*}\right)\boldsymbol{\mu}_{X}^{k,*}+\left(\frac{\sigma_{Y,k}^{*}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k,*}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-1\right)\boldsymbol{\Gamma}_{X}^{k,*}\boldsymbol{\beta}_{-0}\right]\end{pmatrix}\\ \sum_{k=1}^{K}\left(\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-\frac{\sigma_{Y,k}^{*}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k,*}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}\right)\end{pmatrix},$ (B.18)
-which satisfies the central limit theorem for bootstrap means, or the Delta method for bootstrap estimators (see the Appendix as well as Theorem 23.4 and Theorem 23.5 in van der Vaart, (1998)).
+$=\left(\left(\nabla_{00}^{*}\varphi_{n}(\overline{\boldsymbol{\theta}}^{*})\right)^{-1}-\left(\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\right)^{-1}\right)\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})+\left(\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\right)^{-1}\left(\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})-\nabla\varphi_{n}(\boldsymbol{\theta}_{0})\right).$ (26)
+Noticing that $\nabla_{00}^{*}\varphi_{n}(\overline{\boldsymbol{\theta}}^{*})$ and $\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})$ both tend in probability to the same limit $\mathbf{H}(\boldsymbol{\beta},\sigma^{2}_{\epsilon})$ and we have, with similar arguments as those used in the proof of Proposition 4.2, that $\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})$ is $O_{p}(n^{-1/2})$.
+It can be deduced that
+$\boldsymbol{\theta}^{*}-\widehat{\boldsymbol{\theta}}$ $=\left(\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\right)^{-1}\left(\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})-\nabla\varphi_{n}(\boldsymbol{\theta}_{0})\right)+o_{P}(n^{-1/2}).$ (27)
+Using arguments similar to those employed in the expansion of $\nabla\varphi_{n}$ in the proof of Proposition 4.2, we make appear the difference between the bootstrap means and the empirical means or a differentiable functional of these quantities:
+$\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})-\nabla\varphi_{n}(\boldsymbol{\theta}_{0})=\footnotesize\begin{pmatrix}2\sum_{k=1}^{K}\pi_{k}\left((\widehat{\mu}_{Y}^{k}-{\mu}_{Y}^{k,*})-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\left(\widehat{\boldsymbol{\mu}}_{X}-\boldsymbol{\mu}_{X}^{k,*}\right)\right)\\ \begin{pmatrix}2\sum_{k=1}^{K}\pi_{k}\left[\left(\widehat{\mu}_{Y}^{k}-\mu_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\widehat{\boldsymbol{\mu}}_{X}^{k}+\left(\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-1\right)\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}\right]\\ -2\sum_{k=1}^{K}\pi_{k}\left[\left(\mu_{Y}^{k,*}-\mu_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\mu}_{X}^{k,*}\right)\boldsymbol{\mu}_{X}^{k,*}+\left(\frac{\sigma_{Y,k}^{*}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k,*}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-1\right)\boldsymbol{\Gamma}_{X}^{k,*}\boldsymbol{\beta}_{-0}\right]\end{pmatrix}\\ \sum_{k=1}^{K}\left(\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-\frac{\sigma_{Y,k}^{*}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k,*}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}\right)\end{pmatrix},$ (28)
+which satisfies the central limit theorem for the bootstrap means, or the Delta method for the bootstrap estimators (see Theorem A.5 in Section A, as well as Theorem 23.4 and Theorem 23.5 in van der Vaart, (1998)).
 Consequently, $\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})-\nabla\varphi_{n}(\boldsymbol{\theta}_{0})$ and $\nabla\varphi_{n}(\boldsymbol{\theta}_{0})-\nabla\varphi(\boldsymbol{\theta}_{0})$ have the same asymptotic distribution.
-By Slustky’s theorem, the asymptotic distribution of $\sqrt{n}\left(\boldsymbol{\theta}^{*}-\widehat{\boldsymbol{\theta}}\right)$ is thus the same as the asymptotic distribution of $\mathbf{H}(\boldsymbol{\beta},\sigma^{2}_{\epsilon})\sqrt{n}\nabla\varphi_{n}(\boldsymbol{\theta}_{0})$, and we can conclude that $\sqrt{n}\left(\boldsymbol{\theta}^{*}-\widehat{\boldsymbol{\theta}}\right)$ and $\sqrt{n}\left(\widehat{\boldsymbol{\theta}}-\boldsymbol{\theta}_{0}\right)$ have also the same asymptotic Gaussian distribution.
-∎
+By Slustky’s theorem, the asymptotic distribution of $\sqrt{n}\left(\boldsymbol{\theta}^{*}-\widehat{\boldsymbol{\theta}}\right)$ is the same as the asymptotic distribution of $\mathbf{H}(\boldsymbol{\beta},\sigma^{2}_{\epsilon})\sqrt{n}\nabla\varphi_{n}(\boldsymbol{\theta}_{0})$, and we can conclude that $\sqrt{n}\left(\boldsymbol{\theta}^{*}-\widehat{\boldsymbol{\theta}}\right)$ and $\sqrt{n}\left(\widehat{\boldsymbol{\theta}}-\boldsymbol{\theta}_{0}\right)$ also have the same asymptotic Gaussian distributions.
+$\Box$
 Appendix C Supplementary figures
-Figure C.1: a) Coverage rates, b) amplitudes, and c) powers of the confidence intervals for the estimators of $\beta_{1}$ obtained from $N_{sim}$ simulations for each parameter combination.
-The method labels on the left: "mm (asymp)" stands for the method of moments with asymptotic confidence intervals, "mm (boot)" for the method of moments with bootstrap, "ot (boot)" for the optimal transport method with bootstrap, and "mm (student)" for the naive linear regression on means approach based on Student’s distribution.
+Figure 7: a) Coverage rates, b) average amplitudes, and c) powers of the confidence intervals for the estimators of $\beta_{1}$ obtained from $N_{sim}$ simulations for each parameter combination.
+The method labels on the left: ”mm (asymp)” stands for the method of moments with asymptotic confidence intervals, ”mm (boot)” for the method of moments with bootstrap, ”ot (boot)” for the optimal transport method with bootstrap, and ”mm (student)” for the naive linear regression on means approach based on Student’s distribution, and ”simultaneous” for the classical linear regression estimation in the case where the predictor and the predicted variable are observed simultaneously.
+The columns of the tables indicate simulation scenarios with different combinations of parameters: scenario S1 with lower group overlap ($\sigma^{2}_{X}=0.75$) and higher signal-to-noise ratio ($\rho=1.1$), S2 with higher group overlap ($\sigma^{2}_{X}=2$) and higher signal-to-noise ratio ($\rho=1.1$), S3 with lower group overlap ($\sigma^{2}_{X}=0.75$) and lower signal-to-noise ratio ($\rho=1.01$), and S4 with higher group overlap ($\sigma^{2}_{X}=2$) and lower signal-to-noise ratio ($\rho=1.01$).